### v1

# Inference-time Scaling of Diffusion Models through Classical Search

Abstract
Classical search algorithms have long underpinned modern artificial intelligence. In this work, we tackle the challenge of inference-time control in diffusion models—adapting generated outputs to meet diverse test-time objectives—using principles from classical search. We propose a general framework that orchestrates local and global search to efficiently navigate the generative space. It employs a theoretically grounded local search via annealed Langevin MCMC and performs compute-efficient global exploration using breadth-first and depth-first tree search. We evaluate our approach on a range of challenging domains, including planning, offline reinforcement learning, and image generation. Across all tasks, we observe significant gains in both performance and efficiency. These results show that classical search provides a principled and practical foundation for inference-time scaling in diffusion models. Project page at diffusion-inference-scaling.github.io.
1 Introduction
Classical search algorithms have laid the foundation for modern artificial intelligence 10.5555/1671238 . In discrete settings, graph search algorithms are widely used to explore the state space. Breadth-first search (BFS) moore1959shortest and depth-first search (DFS) tarjan1972depth traverse the search tree in a fixed order. To better leverage problem-specific information, best-first search methods pearl1984heuristics , such as A* hart1968formal , use a heuristic to evaluate and prioritize states. Alternatively, local search methods, such as hill-climbing (10.5555/1671238, , Sec. 4.1), explore neighboring states. More recent techniques like gradient descent and Markov Chain Monte Carlo (MCMC) have become widely adopted in optimization and probabilistic inference, underpinning many modern AI models.
Diffusion models ho2020denoising have shown impressive performance in generative modeling for continuous domains such as images dhariwal2021diffusion , videos ho2022video , and world modeling yanglearning . They are also increasingly used in robotics and decision-making liu2024rdt ; black2024pi_0 ; team2024octo to generate diverse actions chi2023diffusion . However, generated samples may not always align with physical laws song2023loss or human intent wallace2024diffusion , and the vast generative space often necessitates multiple trials to produce satisfactory outputs xie2025sana . To address this, we scale up inference-time compute using strategic search methods that navigate the generative manifold for high-quality samples. We formalize sample evaluation using a verifier function $f({\bm{x}}_{0})$ defined on ground truth samples, which measures the quality of the sample. Such verifiers could be reward functions xu2023imagereward , Q-functions lu2023contrastive , classifier conditions $p(c|{\bm{x}}_{0})$ ye2024tfg ; dhariwal2021diffusion , and multi-modal LLMs huang2023t2icompbench .
To efficiently search the generative space induced by diffusion, we revisit classical search principles. Inspired by local search, when the verifier $f$ is differentiable, we use Langevin MCMC to explore neighborhoods of current samples, guided by the verifier gradient and “score function” vincent2011score_matching from the diffusion model. To explore diverse modes in the complex distribution generated by diffusion models, we search the Markov chain of the “denoising process” with global search methods. Viewing sampling as a search tree, we use BFS and DFS to gradually explore the states while denoising. Like best-first search methods, we can evaluate intermediate states ${\bm{x}}_{t}$ using the verifier $f({\bm{x}}_{0|t})$, expanding high-quality paths more effectively. An illustration of local and global search is shown in Fig. 1.
Recently, Ma et al. ma2025inference also explored inference-time scaling of diffusion models using search-based techniques, focusing on image generation tasks. Their zero-order search approach can be interpreted as a instance of local search, while their path-based strategies a combination of both local and global search. However, their method does not effectively leverage verifier gradients, and we introduce a local search strategy with annealed Langevin MCMC, efficiently steering sampling toward high-score regions. Moreover, we systematically evaluate the compute efficiency of search methods. Finally, we extend inference-time scaling beyond image generation to a broader set of domains.
To this end, we summarize our key contributions as follows:
  • i)
    We propose a general and unified framework for efficient inference-time search in diffusion models, drawing inspiration from classical search algorithms. Specifically, we introduce a local search method based on Langevin MCMC that theoretically unifies prior gradient-based approaches, and we develop efficient global search strategies using tree-based algorithms such as BFS and DFS.
  • ii)
    We implement inference-time scaling across a range of tasks—including planning, offline reinforcement learning, and image generation—and introduce task-specific verifiers. Our results demonstrate that inference-time scaling with search against verifiers can significantly improve the performance on many downstream tasks. Throughout the experiments, we show that our proposed approaches consistently surpass baselines in terms of compute efficiency.
Figure 1: Illustration of our search framework. Bottom left: direct sampling results in samples with low verifier scores. Bottom middle: Local search navigates samples to nearby high-score regions, but suffers from local maximums and OOD samples. Bottom right: introducing global search explores diverse modes globally, allocating more compute to high-quality samples, and sampling from global optimal modes.
2 Related Works
Here we provide a brief overview of inference-time scaling and diffusion models. For a more in-depth literature review see Appendix. B.
Scaling inference-time compute has its roots in cognitive science, often associated with “System 2” thinking kahneman2011thinking ; anthony_thinking_2017 . Yao et al. yao2023tree conceptualize problem-solving in large language models (LLMs) as a combinatorial tree search problem newell1959report ; newell1972human , proposing the Tree-of-Thoughts framework, which leverages breadth-first and depth-first search for more efficient reasoning.
Inference-time scaling has recently shown impressive performance with LLMs in reasoning and problem-solving tasks. For instance, jaech2024openai ; van2016deep train models to generate long chain-of-thought (CoT) wei2022chain , while others such as snell2024scaling ; wu2024inference incorporate process reward models lightman2023let to guide the generation process more effectively. These works focus on efficient exploration in discrete search spaces, often employing global search strategies. In contrast, diffusion models operate in a continuous, multimodal generative space. Furthermore, the autoregressive nature of LLMs allows for efficient evaluation of intermediate generation steps—something not directly feasible in diffusion models dhariwal2021diffusion ; lu2023contrastive .
Inference-time scaling for diffusion models remains relatively underexplored. Recent works such as ma2025inference ; nakamotosteering introduce BFS-style global search algorithms, resampling candidate samples (or “particles”) at each noise level based on verifier scores (potentials). However, these methods do not effectively perform local search. Alternatively, Du et al. du_learning_2024 propose learning a series of annealed energy landscapes and applying Langevin MCMC for iterative refinement. In this work we propose a general principled framework, encompassing prior methods as instances of local search and global search.
3 Backgrounds
3.1 Diffusion Probabilistic Models
Suppose we have $D$-dimensional random variable ${\bm{x}}_{0}\in\mathbb{R}^{D}$ with distribution $p_{0}({\bm{x}}_{0})$. Diffusion models ho2020denoising ; song2020denoising and the more general flow models lipman2022flow ; albergo2022building are generative models that turn noise into data via a stochastic process $\left\{{\bm{x}}_{t}\right\}_{t=0}^{T}$. The forward “noising” process with $t>s$ can be defined as:
$q({\bm{x}}_{t}|{\bm{x}}_{s})=\mathcal{N}\left({\bm{x}};\frac{\alpha_{t}}{\alpha_{s}}{\bm{x}}_{s},\alpha_{t}^{2}\left(\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}-\frac{\sigma_{s}^{2}}{\alpha_{s}^{2}}\right)\bm{I}\right)\,.$ (1)
where $\alpha_{t}$, $\sigma_{t}$ are referred as the noise schedule with $\alpha_{0}=\sigma_{T}=1$, $\alpha_{T}=\sigma_{0}=0$. We can thus write the random variables ${\bm{x}}_{t}$ as an interpolation between data and noise ma2024sit :
${\bm{x}}_{t}=\alpha_{t}{\bm{x}}_{0}+\sigma_{t}{\bm{\epsilon}}\,,$
and denote $q_{t}({\bm{x}}_{t})$ as the marginal distribution of ${\bm{x}}_{t}$. To model the reverse “denoising” process, we train the model using the denoising objective ho2020denoising :
$L(\theta)=\mathbb{E}_{t,{\bm{x}}_{0},{\bm{\epsilon}}}\left[{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)-{\bm{\epsilon}}\right]\,,$
which is equivalent of learning the score function of $q_{t}({\bm{x}}_{t})$ vincent2011score_matching , as the ground truth of $\epsilon_{\theta}({\bm{x}}_{t},t)$ is $-\sigma_{t}\nabla_{{\bm{x}}_{t}}\log q_{t}({\bm{x}}_{t})$. To generate samples, we transform noise into data via the reverse transition kernel $p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$. In practice, we either sample ${\bm{x}}_{t-1}$ using deterministic samplers like DDIM song2020denoising :
${\bm{x}}_{t-1}=\frac{\alpha_{t-1}}{\alpha_{t}}({\bm{x}}_{t}-\sigma_{t}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t))+\sigma_{t-1}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)$
or stochastic samplers like DDPM ho2020denoising ; nichol2021improved :
$p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})=\mathcal{N}\left({\bm{x}}_{t-1};\bm{\mu}_{\theta}({\bm{x}}_{t},t),\bm{\Sigma}_{\theta}({\bm{x}}_{t},t))\right)\,.$
3.2 Compositional and Controllable Generation of DPMs
Given a base diffusion model with data distribution $p_{0}({\bm{x}}_{0})$, one may wish to sample ${\bm{x}}_{0}$ with some constraints or conditions $f({\bm{x}}_{0})$. Exact diffusion sampling from the composed distribution $\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})f({\bm{x}}_{0})$ would require training a time-dependent $f$ on data generated by $p_{0}$ dhariwal2021diffusion ; lu2023contrastive , which may not be applicable in practice. Thus, we adopt optimization based methods to approximate the target distribution.
Compositional generation via annealed Langevin MCMC
When sampling from a compositional distribution composed of multiple probability distributions, $\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})\hat{p}_{0}({\bm{x}}_{0})$, du2023reduce proposes annealed Langevin MCMC sampling. In this approach, a sequence of annealed distributions $\tilde{q}_{t}({\bm{x}}_{t})\propto q_{t}({\bm{x}}_{t})\hat{q}_{t}({\bm{x}}_{0})$ is constructed, and samples are drawn using Langevin dynamics welling2011bayesian :
${\bm{x}}_{t}^{i+1}={\bm{x}}_{t}^{i}+\eta\nabla_{{\bm{x}}}\log\tilde{q}_{t}({\bm{x}}_{t}^{i})+\sqrt{2\eta}{\bm{\epsilon}}^{i}\,,~{}~{}~{}{\bm{\epsilon}}^{i}\sim\mathcal{N}(\bm{0},\bm{I})\,.$ (2)
Since the distribution of ${\bm{x}}_{t}^{i}$ converges to $\tilde{q}_{t}({\bm{x}}_{t})$ asymptotically as $i\rightarrow\infty$, $\eta\rightarrow 0$, we can sample from $\tilde{p}_{0}({\bm{x}}_{0})$ following the annealing path $\left\{\tilde{q}_{t}\right\}_{t=0}^{T}$ with $\tilde{q}_{0}=\tilde{p}_{0}$. Moreover, since the score of $\tilde{q}_{t}$ can be directly computed by composing the score of two distributions:
$\nabla_{{\bm{x}}}\log\tilde{q}_{t}({\bm{x}}_{t})=\nabla_{{\bm{x}}}\log q_{t}({\bm{x}}_{t})+\nabla_{{\bm{x}}_{t}}\log\hat{q}_{t}({\bm{x}}_{t})\,,$
thus do not require extra training.
Controllable generation through training-free guidance
During the sampling process, training-free guidance propose to update ${\bm{x}}_{t}$ using gradient ascent
$\tilde{{\bm{x}}}_{t}={\bm{x}}_{t}+\bm{\Delta}_{t}\,,~{}~{}\bm{\Delta}_{t}=\rho_{t}\nabla_{{\bm{x}}_{t}}\log f({\bm{x}}_{0|t})+\mu_{t}\alpha_{t}\nabla_{{\bm{x}}_{0|t}}\log f({\bm{x}}_{0|t})\,.$ (3)
where ${\bm{x}}_{0|t}=\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]=\frac{{\bm{x}}_{t}-\sigma_{t}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)}{\alpha_{t}}$. This method approximates the intractable posterior with the posterior mean: $\mathbb{E}_{{\bm{x}}_{0}|{\bm{x}}_{t}}[f({\bm{x}}_{0})]\approx f(\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}])$. To enhance the guidance strength, yu2023freedom propose to apply a recurrence strategy, which first samples ${\bm{x}}_{t-1}$ via $p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$, add the guidance gradient, then add noise back to ${\bm{x}}_{t}$ through the forward process $q_{t}({\bm{x}}_{t}|{\bm{x}}_{t-1})$:
${\bm{x}}_{t-1}^{i}\sim p_{\theta}(\cdot|{\bm{x}}_{t}^{i})\,,~{}~{}\tilde{{\bm{x}}}_{t-1}^{i}={\bm{x}}_{t-1}^{i}+\frac{\alpha_{t-1}}{\alpha_{t}}\bm{\Delta_{t}}\,,~{}~{}{\bm{x}}_{t}^{i+1}\sim q_{t}(\cdot|\tilde{{\bm{x}}}_{t-1}^{i})\,,~{}~{}~{}i=1,2,\cdots,N_{\text{recur}}\,,$ (4)
with $N_{\text{recur}}$ being the total number of recurrence steps.
4 Methods
Problem Formulation
Given a pretrained diffusion model $\epsilon_{\theta}({\bm{x}}_{t},t)$ with a base distribution $p_{0}({\bm{x}}_{0})$, at test-time, we often wish to optimize the generation process to satisfy task-specific objectives. For instance, in reinforcement learning, this might involve generating high-value actions; in image synthesis, we may seek images that satisfy certain constraints such as color or shape; and in trajectory generation, we may want the outputs to conform to physical laws. In this paper, we are interested in how to scale test-time inference to follow such objectives.
We consider an inference-time scaling strategy that adjusts the sampling process based on a verifier function. Specifically, we define a verifier $f({\bm{x}}_{0}):\mathbb{R}^{D}\rightarrow\mathbb{R}^{+}$ which specifies the degree to which samples optimize a specified objective. We then aim to bias sampling toward regions of the sample space where $f({\bm{x}}_{0})$ is high. This leads to the objective of sampling from a compositional distribution that combines the original model distribution with the verifier:
$\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})f({\bm{x}}_{0})^{\lambda}\,,$
where $\lambda$ controls the weight of verifier scores.
Since exact sampling from the distribution is often impractical, we aim to search the manifold for the target samples at inference time, both locally and globally. First, we search the vicinity of the sample using hill-climbing style local search methods, guided by the verifier gradient. However, local search may stuck in a local maximum, where we then propose global graph search algorithms to explore the diverse modes in the complex generative landscape of diffusion models.
4.1 Local Search via Langevin MCMC with Verifier Gradient
To sample from the composition distribution $\tilde{p}_{0}$, we conduct local-search with hill-climbing methods, aiming to find the local maximum with high $\tilde{p}_{0}$. Specifically, we view the sampling problem as compositional optimization in measure space wibisono2018sampling , and follow the gradient flow of KL-divergence, performing Langevin MCMC steps (details see Appendix. C.1).
Similar to annealed Langevin MCMC in du2023reduce , we could construct a series of annealed functions $\hat{f}_{t}({\bm{x}}_{t})$ with $\hat{f}_{0}({\bm{x}}_{0})=f({\bm{x}}_{0})$. Then we sample from the distributions $\tilde{q}_{t}({\bm{x}}_{t})\propto q_{t}({\bm{x}}_{t})\hat{f}({\bm{x}}_{t})$ through Langevin MCMC in Eq. 2 (details see Appendix. C.2). However, it remains unclear how to design such annealed verifier functions $\hat{f}_{t}({\bm{x}}_{t})$ efficiently, especially without training. Alternatively, training-free guidance in Eq. 3 utilizes the gradient of $f({\bm{x}}_{0|t})$ to optimize ${\bm{x}}_{t}$, which can be computed directly using the diffusion model output. However, naive gradient updates have been observed to produce OOD and adversarial samples shen2024understanding . In ye2024tfg , recurrence (Eq. 4) was found to help avoid adversarial samples in challenging guidance tasks, though its theoretical underpinnings remain poorly understood. We unify these two approaches by demonstrating that training-free guidance with recurrence, in the continuous limit, constitutes an instance of Langevin MCMC. For details see Appendix. C.3.
Proposition 1.
In the continuous limit where the number of diffusion denoising steps $T\rightarrow\infty$, training-free guidance with recurrence is equivalent to running Langevin MCMC on a series of annealed distributions $\left\{\tilde{q}_{t}({\bm{x}}_{t})\right\}_{t=0}^{T}$, with $\tilde{q}_{0}({\bm{x}}_{0})=\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})f({\bm{x}}_{0})^{\lambda}$.
Thus, the recurrence step (without guidance) can be interpreted as Langevin MCMC applied to the original distribution of the diffusion model $q_{t}({\bm{x}}_{t})$, effectively guiding the samples back toward the data manifold (details see Appendix. C.4). The guidance term then serves as defining a practical annealing path $\hat{f}_{t}({\bm{x}}_{t})$, biasing the sampling path. We are the first to propose this theoretical unification of the two lines of work, providing insights into efficient local search of diffusion models via gradients.
To implement local search, we parameterize the reverse transition kernel $\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$ as a sequence of Langevin MCMC steps in Eq. 2, followed by a denoising step with DDIM (Eq. 10) or DDPM (Eq.11). Owing to the equivalence between Langevin MCMC and training-free guidance, we can efficiently tune hyper-parameters using the unified design space of ye2024tfg , with details in Appendix. C.5. We scale up the number of local search steps and observe significant improvements across multiple tasks (Fig. 4(d)), highlighting the effectiveness of local search.
4.2 Global Search for Mode Identification
Figure 2: Illustration of global tree search algorithms. Left: Best-of-N sampling denoises multiple samples in parallel and selects the one with the highest verifier score at the final step. Middle Left: BFS-Resampling resamples particles at intermediate noise levels, allocating more children to nodes with higher scores. Middle Right: BFS-Pruning removes particles with low verifier scores during denoising, sampling at most one child per node. Right: DFS denoises a single particle until its verifier score violates the constraint, then backtracks by adding noise to a previous level and continuing the search.
Although Langevin MCMC is asymptotically exact, it often struggles to transition between distant modes latuszynski2025mcmc due to its inherently local nature. This limitation is especially pronounced in the context of the multimodal distributions induced by diffusion models. To mitigate this, we propose identifying high-score modes via a global search strategy.
We model the Markov chain of the denoising process as a fixed-depth tree, where the transition kernel is given by $\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$. As discussed earlier, when applying local search, we parameterize the transition kernel with multiple Langevin MCMC steps, and otherwise we directly adopt conventional diffusion samplers song2020denoising ; ho2020denoising ; lu2022dpm . Hence, our perspective unifies both stochastic and deterministic sampling with guided sampling. This abstraction decouples the transition kernel from the sampling process, enabling the application of classic tree search heuristics to develop compute-efficient exploration methods.
BFS-style linear search
Inspired by breadth-first search (BFS), which expands nodes level-by-level, we denoise a set of particles in parallel at each noise level. The simplest approach is best-of-$N$ sampling: generate $N$ candidate trajectories and select the one with the highest verifier score at the final step. While simple, this strategy ignores information from intermediate stages.
To improve efficiency, similar to best-first search pearl1984heuristics , we evaluate each intermediate particle $\left\{{\bm{x}}_{t}^{k}\right\}_{k=1}^{K_{t}}$ using its verifier score $f({\bm{x}}_{0|t}^{k})$, where $K_{t}$ is the number of particles at timestep $t$. Based on these scores, we dynamically reallocate computation by sampling more children for high-scoring nodes via the reverse transition kernel $\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$. This forms the basis of BFS-Resampling, where each node samples children in proportion to its normalized verifier score
$n_{t}^{k}=\texttt{Round}\left(K_{t}\frac{f({\bm{x}}_{0|t}^{k})^{\tau_{t}}}{\sum_{i=1}^{K_{t}}f({\bm{x}}_{0|t}^{i})^{\tau_{t}}}\right)\,.$
We then sample $n_{t}^{k}$ children from ${\bm{x}}_{t}^{k}$ using $\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$ and proceed to the next denoising step.
When using deterministic ODE solvers lu2022dpm as transition kernels, it is ineffective to sample multiple children from the same parent. To address this, we propose BFS-Pruning, a variant that retains only top-scoring nodes and allows at most one child per parent $n_{t}^{k}=\min\left(1,n_{t}^{k}\right)$. This leads to a gradual narrowing of the search tree as $K_{t}$ decreases over time, enabling us to explore more diverse modes with more initial particles under the same compute.
As shown in Fig. 6, both resampling and pruning significantly improve efficiency over standard best-of-$N$ sampling. The detailed pseudo code is in Alg. 3.
DFS-style non-linear search
Depth-first search (DFS) explores one branch of the search tree as deeply as possible before backtracking. In our setting, this corresponds to iteratively denoising a single particle until its verifier score drops below a predefined threshold: $f({\bm{x}}_{0|t})\leq\delta_{t}$, where $\delta_{t}$ is a scheduled threshold for timestep $t$.
Once the constraint is violated, the algorithm backtracks by reintroducing noise, moving to a higher noise level $t_{\text{next}}=t+\Delta_{T}$ using the forward diffusion process $q({\bm{x}}_{t_{\text{next}}}|{\bm{x}}_{t})$ in Eq. 1. This allows the model to restart the denoising process from a different region of the manifold, encouraging exploration of diverse modes. By placing a cap on the number of backtracking steps, we can explore the space globally while keeping computation bounded.
A key strength of DFS is its ability to allocate compute adaptively: difficult prompts and low-quality trajectories naturally trigger more backtracking and exploration, while easier instances are solved more directly. This dynamic behavior is driven purely by the verifier signal, without needing to know the difficulty in advance. As shown in Fig. 5, this strategy leads to substantial gains in efficiency and performance. The detailed pseudo code is in Alg. 4.
5 Experiments
In this section, we apply inference-time scaling with our search strategy across a range of domains. In Sec.5.1, we show that diffusion models can generate long-horizon plans in complex maze environments with inference-time scaling, given a world model as verifier. In Sec.5.2, we leverage a pretrained Q-function as a verifier to guide diffusion policies, achieving strong performance on D4RL fu2020d4rl locomotion benchmarks. Finally, in Sec. 5.3, we approach the image generation task, and demonstrate that our global search algorithms consistently outperform best-of-N sampling in terms of efficiency, and we reveal interesting properties of BFS and DFS.
5.1 Long Horizon Planning
Figure 4: (Illustration and results for maze planning) Fig. 4(a) shows the maze layout of Giant (left) and Ultra (right), with start: $\bullet$ and goal: $\star$. Fig. 4(b) illustrates the effect of local search, where left shows the directly sampled trajectory that violates the maze block on the bottom left, and right shows the successful sample after local search. Fig. 4(c) shows that inference time search generates diverse solutions. Fig. 4(d) shows the performance with scaling number of local search steps with Langevin MCMC. Fig. 4(e) and Fig. 4(f) shows the scaling behaviors of the global search methods for PointMaze Giant and PointMaze Ultra respectively. Details see Appendix. E.1.
Diffusion models have been widely adopted in planning for trajectory synthesis ubukata2024diffusion . We evaluate long-horizon planning in the PointMaze environment from OGBench park2024ogbench , training the diffusion model following the approach in Diffuser janner2022planning , where the full trajectory is denoised:
$\bm{\tau}=\left[\begin{array}[]{cccc}\bm{s}_{1}&\bm{s}_{2}&\cdots&\bm{s}_{H}\\ \bm{a}_{1}&\bm{a}_{2}&\cdots&\bm{a}_{H}\end{array}\right]\,.$
To highlight the effectiveness of inference-time scaling, we evaluate on the largest PointMaze Giant environment and introduce an even more challenging variant, PointMaze Ultra, following the protocol from marcucci2023motion . The maze layouts are shown in Fig. 4(a). We define the verifier loss for each colliding point of the trajectory as the squared distance to its nearest wall, thus the overall verifier score measures the degree the trajectory violates the environment (details see Appendix. E.1). This verifier functions as a physics-grounded instance of a world model. Importantly, naively maximizing the verifier score does not guarantee a successful plan, and planning remains difficult even with full access to the maze layout luo2024potential ; marcucci2023motion .
In the most challenging scenario—navigating from the bottom-left to the top-right corner—naive sampling strategies rarely succeed, necessitating multiple local search steps (Fig. 4(d)). As illustrated in Fig. 4(b), local search guided by the verifier enables the model to find successful plans that navigate around obstacles. As shown in Fig. 4(e) and Fig. 4(f), incorporating global search alongside local refinement leads to over 90% success rates in both maze environments, with both BFS and DFS significantly improving search efficiency compared to naive best-of-$N$ sampling. Moreover, as shown in Fig. 4(c), global search enables the generation of diverse, high-quality solutions.
5.2 Offline Reinforcement Learning
Table 1: Performance on D4RL locomotion tasks, score normalized as suggested in fu2020d4rl . Following lu2023contrastive we report the mean and standard deviation over 5 independent trials, highlighting the numbers within 5 percent of the maximum in every individual task. To faithfully report our performance, we conduct the hyper-parameter search and evaluation with different seeds. For more details see Appendix. E.2.
Dataset Environment CQL BCQ IQL SfBC DD Diffuser D-QL QGPO TTS(ours)
Medium-Expert HalfCheetah 62.4 64.7 86.7 92.6 90.6 79.8 96.1 93.5 $\mathbf{93.9\pm 0.3}$
Medium-Expert Hopper 98.7 100.9 91.5 108.6 111.8 107.2 110.7 108.0 $104.4\pm 3.1$
Medium-Expert Walker2d 111.0 57.5 109.6 109.8 108.8 108.4 109.7 110.7 $\mathbf{111.4\pm 0.1}$
Medium HalfCheetah 44.4 40.7 47.4 45.9 49.1 44.2 50.6 54.1 $\mathbf{54.8\pm 0.1}$
Medium Hopper 58.0 54.5 66.3 57.1 79.3 58.5 82.4 98.0 $\mathbf{99.5\pm 1.7}$
Medium Walker2d 79.2 53.1 78.3 77.9 82.5 79.7 85.1 86.0 $\mathbf{86.5\pm 0.2}$
Medium-Replay HalfCheetah 46.2 38.2 44.2 37.1 39.3 42.2 47.5 47.6 $\mathbf{47.8\pm 0.4}$
Medium-Replay Hopper 48.6 33.1 94.7 86.2 100.0 96.8 100.7 96.9 $\mathbf{97.4\pm 4.0}$
Medium-Replay Walker2d 26.7 15.0 73.9 65.1 75.0 61.2 94.3 84.4 $79.3\pm 9.7$
Average (Locomotion) 63.9 51.9 76.9 75.6 81.8 75.3 86.3 86.6 86.1
Recently, diffusion models emerged as a powerful action prior in robotics due to its ability to model complex and multimodal distributions chi2023diffusion ; liu2024rdt . However, these diffusion policies are typically trained on offline datasets and struggle to adapt to reinforcement learning or test-time requirements. Following prior work peters2010relative , we formulate the offline RL problem as sampling from a Q-regularized distribution:
$\pi^{*}(\bm{a}|\bm{s})\propto\mu(\bm{a}|\bm{s})e^{\beta Q_{\psi}(\bm{s},\bm{a})}\,,$
where $Q_{\psi}$ is a learned Q-function representing preferences over actions, and $\mu$ is the behavior policy, which we model using a diffusion prior.
In principle, exact sampling from $\pi^{*}$ using diffusion models requires training a noise-dependent energy function on noisy actions $\bm{a}_{t}$ janner2022planning ; lu2023contrastive . However, this approach is often impractical in real-world settings. First, it requires evaluating the pretrained Q-function on samples from the diffusion model, which is costly due to the difficulty of data collection in robotics. Second, many Q-functions rely on foundation models as backbones brohan2023can ; nakamotosteering , and training them on noisy samples risks degrading their pretrained knowledge. In contrast, test-time search (TTS) offers a promising alternative. By decoupling policy and value, TTS enables flexible composition with foundation models without retraining.
We evaluate TTS using both local and global search strategies on the D4RL locomotion benchmarks fu2020d4rl . Following the setup in lu2023contrastive , we adopt their pretrained diffusion policy and ground-truth Q-function. Results are presented in Table 1, where our approach achieves performance competitive with the best training-based methods, including QGPO lu2023contrastive and Diffusion-QL wang2022diffusion . In comparison, our approach is training-free, requiring only an off the shelf Q function and diffusion model.
5.3 Image Generation
In this section we evaluate inference-time scaling in image generation tasks, especially the efficiency of global search.
5.3.1 Compositional Text-to-Image Generation
Figure 5: CompBench huang2023t2icompbench text-to-image results. Fig. 5(a) and Fig. 5(b) show DFS performance on attribute binding and object relationship tasks, achieving up to 2× lower computational cost than Best-of-N. Fig. 5(c) presents the average compute allocated by DFS for prompts of increasing difficulty, evaluated on the color dataset.
We evaluate inference-time scaling for text-to-image generation on the CompBench dataset huang2023t2icompbench using the SSD-1B model gupta2024progressive . For the search verifiers, we use task-specific oracle verifiers: a BLIP-VQA model li2022blip for the attribute binding task and a UniDet model zhou2022simple for the object relationship task. Since these verifiers are non-differentiable, we apply global search exclusively. Detailed setup is in Appendix. E.3.1.
Improved efficiency and adaptivity with DFS
We evaluate Best-of-N and DFS on the five datasets in attribute binding and object relationship tasks, with each dataset consists of 300 prompts. The results are shown in Fig. 5(a) and Fig. 5(b), where DFS consistently outperforms Best-of-N while using up to 2× less compute. To evaluate DFS’s ability to adaptively allocate computation, we divide the 300 prompts in the color dataset into four difficulty bins based on their average score over 8 samples. Using the same set of hyperparameters across all prompts, Fig. 5(c) shows that DFS allocates less compute to easier prompts (bin 1) and progressively more to harder ones with lower scores—demonstrating automatic, difficulty-aware compute allocation. Unlike snell2024scaling , which manually selects different search strategies based on predefined difficulty labels, our method requires no prior knowledge of difficulty. DFS dynamically adjusts compute allocation during inference using a single, unified algorithm and fixed parameters.
5.3.2 Conditional ImageNet Generation
Conditional generation of diffusion models using a unconditional foundation model has important applications ye2024tfg as foundation models obtain increasing expressivity. However, classifier guidance dhariwal2021diffusion requires training a noise-dependent classifier on samples generated by the diffusion model, which is difficult since it may require a lot of attempts for an unconditional model to generate the correct class. In this section, we treat the pretrained classifier as the verifier function and conduct inference-time scaling with both local and global search. The experimental details are in Appendix. E.3.2.
Table 2: Best-of-N results for ImageNet conditional generation, with FID and Accuracy averaged across the two labels.
#Particles Single verifier Double-verifier
[…]
2 188.5 25.6% 178.4 29.7%
4 171.5 31.8% 151.2 37.5%
8 155.7 35.8% 127.8 49.2%
Mitigating verifier hacking with double-verifier
One of the major challenges of guiding the samples using the verifier gradient is that image classifiers are prone to adversarial examples szegedy2013intriguing . Inspired by the double-Q learning in RL van2016deep , we use two verifiers for local search gradient computation and global search respectively. Specifically, we use two different vision transformers xu2023imagereward with input resolution 224 and 384. For accuracy evaluation we use another classifier from touvron2021training . As shown in Table. 2, using double-verifier consistently improves performance with Best-of-N sampling, with up to 2x less compute.
Figure 6: Results for ImageNet conditional generation with BFS Fig. 6(a) and Fig. 6(b) show the FID and class accuracy of conditional samples. BFS consistently outperform Best-of-N sampling in efficiency.
Improved Efficiency with BFS
We investigate breadth-first search (BFS) as a means to improve the computational efficiency of Best-of-N sampling. All experiments use the double-verifier setup, where one verifier guides gradient-based local search and the other handles resampling or pruning. As shown in Fig. 6, both BFS-Resampling and BFS-Pruning outperform Best-of-N in terms of compute efficiency. BFS-Pruning is more effective under limited compute budgets, while BFS-Resampling achieves higher performance with increased compute.
6 Limitations and Conclusion
Limitations
Despite the performance improvement of inference-time scaling, it still requires more hyperparameter tuning than training-based methods. One approach is to design an adaptive hyperparameter tuning framework, which could increase or decrease automatically based on the current sample. For example we can tune the threshold of DFS automatically given the pass rate of samples. Another potential concern is that the generated samples may be used as adversarial examples to attack verifiers. Our method can be used to search for weaknesses of verifiers, which in turn can be used to train a more robust verifier.
Conclusion
In this work we propose a principled framework for inference-time scaling through classic search principles, unlocking the potential of scaling inference-time compute of diffusion models in a variety of downstream tasks. Inference-scaling can also be applied to training models via expert iteration, generating high-quality synthetic data beyond the model’s ability. We leave this as a interesting future direction.
Appendix A Appendix Overview
In Sec. B, we provide a in-depth review of literature related to inference-time scaling and diffusion models. In Sec. C, we elaborate on local search with Langevin MCMC, and in Sec. D we provide the pseudo code and design of global search algorithms BFS and DFS. In Sec. E, we provide the details of all the experiments.
Appendix B Additional Related Works
Inference-time scaling
Scaling compute in inference-time with “slow thinking” has its long history grounded in cognitive science, known as “system 2” thinking [32, 63]. In [52, 53], Newell and his colleagues formalized problem solving as tree search in a combinatorial space, and [93] uses tree-of-thoughts to enable LLM reasoning with multiple exploration paths, using BFS and DFS as strategic search algorithms.
Recently, long chain-of-though (CoT) reasoning has demonstrated remarkable performance for LLM reasoning [30, 21], where the long CoT reasoning ability is incentivized through reinforcement learning [21]. Notably, the CoT process demonstrates reasoning activities such as self-verification, backtracking and self-correction. Using a process reward model [39], we can also conduct explicit tree search without training the language model. [85] propose reward-balanced search (REBASE) which is a special instance of BFS, and [64] applied beam-search to difficult math problems, showing compute-optimal inference can be achieved via selecting different strategy for problems with different difficulty. We refer the readers to [79] for a comprehensive review.
Inference-time scaling could also be used to improve the model itself, known as expert iteration. In [32], they proposed reinforcement learning with expert iteration in games, where the expert is constructed combining the base policy with Monte Carlo Tree Search (MCTS), and [86] applied expert iteration to automated theorem proving. Self improvements can also be achieved through iterative self-training [69, 96, 13, 61].
Inference scaling in diffusion models
The inference-time compute of diffusion models depends heavily on the number of denoising steps. [91] showed that recursive restart sampling can reduce cumulative error during sampling, which can be regarded as scaling the number of denoising steps. More recently, [62] proposed a Sequential Monte Carlo (SMC) [84, 99] style method, known as Feynman-Kac steering, which can be seen as a instance of BFS. Besides image generation, they also applied it to diffusion language models [23, 59]. Additionally, [48] explored inference-time scaling of diffusion models with local zero order search and global search over paths for image generation. They also experimented with different verifiers, such as oracle verifiers, self-supervised verifiers, and studied the verifier-task alignment problem. There zero-order search can be understood as an uninformed version of local search which exhibits low efficiency, and when utilizing gradients, they need to back-propagate through the entire diffusion sampling chain, causing high computation and memory overheads. Compared with their work, we propose efficient gradient-based local search with theoretically grounded Langevin MCMC, which we show is crucial in many tasks. Also, we provide systematic experiments on the compute efficiency of global search methods. Their proposed methods can thus be understood as a instance within our search framework.
Apart from search, [38] exploits the in-context learning abilities of foundation models to provide revision during sampling. Specifically, they leverage the multi-modal capabilities of VLMs to provide feedbacks on past generated images, and train the model to condition on past images and feedback.
Diffusion models and applications
Diffusion models [26, 68, 76] has shown great performance in generative modeling for continuous data domains, such as image [18, 12], videos [28] and molecules [90]. Due to its expressive power on modeling multimodal and complex distributions, they have also been widely used as a decision prior in robotics. [31] proposes the first work on using diffusion models to generate plans. [10] uses the diffusion model for visuo-motor policy in robotics. Recently, a series of robotics foundation models utilize diffusion heads as action experts [5, 73], while [43] trains a end-to-end diffusion transformer for bimanual manipulation. In this work, we demonstrate that inference-time scaling can be especially helpful for decision making tasks with diffusion models.
Control and alignment of diffusion models
To align the diffusion model with flexible objectives, training-free guidance [94, 4, 11, 66, 95, 25] and compositional generation [14, 42] combines the diffusion models with classifiers or other diffusion models at inference time, while RL-based methods [98, 78, 89, 6] finetune the diffusion model using reward or preference signals. ControlNet [97] style approaches have also been used to add additional conditions for sampling, where [71] designs a control block for diffusion transformers, and [88] uses a combination of guidance and controlnet for controllable human motion generation.
Appendix C Details about local search with Langevin MCMC
In this section we provide a comprehensive and detailed overview of (annealed) Langevin MCMC based methods used in local search, as well as proving Proposition. 1.
C.1 Langevin MCMC as gradient flow in measure space
Following [83], the Langevin SDE in sample space corresponds to gradient flow of the KL-divergence in measure space. Here we provide a brief overview.
Define our target distribution that we wish to sample from as $\nu$, and the distribution of our current sample as $\rho$. We define the KL-divergence (relative entropy) as:
$H_{\nu}(\rho)=\int\rho\log\frac{\rho}{\nu}\,.$ (5)
Thus, sampling from $\nu$ can be seen as minimizing $H$, since the minimum of $H$ is achieved at $\rho=\nu$ with $H_{\nu}(\rho)=0$. Furthermore, $\nu$ is the only stationary point of $H$ even for multimodal distributions. Thus we can sample from $\nu$ when optimizing $H$ via gradient based methods.
We have the gradient flow of $H$ in Eq. 5 follows the following PDE:
$\frac{\partial\rho}{\partial t}=\nabla\cdot\left(\rho\nabla(-\log\nu)\right)+\Delta\rho\,,$ (6)
which is known as the Fokker-Planck equation. Here, $\rho=\rho({\bm{x}},t)$ is a smooth positive density evolving through time, driven by the dynamics of the sample ${\bm{x}}$. The dynamics in sample space corresponding to Eq. 6 is the Langevin SDE:
$d{\bm{x}}_{t}=\nabla\log\nu({\bm{x}}_{t})dt+\sqrt{2}d\bm{w}_{t}\,.$ (7)
where $({\bm{x}}_{t})_{t\geq 0}$ is a stochastic process with measure $\rho_{t}$, and $(\bm{w}_{t})_{t\geq 0}$ is standard Brownian motion. That is, if ${\bm{x}}_{t}\sim\rho_{t}$ evolves according to the dynamics in Eq. 7, then the measure $\rho({\bm{x}},t)=\rho_{t}$ evolves according to the PDE in Eq. 6, conducting gradient optimization in measure space.
In practice, we implement Eq. 7 through discretization, which is known as the unadjusted Langevin algorithm (ULA):
${\bm{x}}^{i+1}={\bm{x}}^{i}+\eta\nabla_{{\bm{x}}^{i}}\log\nu({\bm{x}}^{i})+\sqrt{2\eta}{\bm{\epsilon}}^{i}\,,$ (8)
with ${\bm{\epsilon}}^{i}\sim\mathcal{N}(\bm{0},\bm{I})$. When $\eta\rightarrow 0$, the ULA converges to Langevin SDE, providing exact sampling.
Previous works [17, 9] show that ULA can efficiently converge to the target measure $\nu$ if $\nu$ is log-concave and smooth. However, when facing complex and multimodal distributions, we can only guarantee convergence to the concave vicinity.
C.2 Annealed Langevin MCMC Sampling
Langevin MCMC have been used to perform implicit sampling in energy-based models [16] and score-based models [67]. However, these methods suffer from inaccurate score estimation and low density regions [67]. In [67] they propose to perturb the data with gaussian noise, eventually smoothing the data distribution:
$q({\bm{x}}_{t})=\int_{{\bm{x}}_{0}}p_{0}({\bm{x}}_{0})\mathcal{N}({\bm{x}}_{t};{\bm{x}}_{0},\sigma_{t}^{2}\bm{I})\,,$
and creating a sequence of annealed distributions $\left\{q({\bm{x}}_{t})\right\}_{t=0}^{T}$ which converges to $p_{0}({\bm{x}}_{0})$. Since they are smoothed by gaussian noise, we can improve the mixing time of Langevin MCMC on multimodal distributions by sampling from these intermediate distributions, sharing similar spirits with simulated annealing [33].
In [14], they extend this method to compositional generation of diffusion models. Specifically, we consider sampling from a product distribution $p^{\text{prod}}_{0}({\bm{x}}_{0})\propto p_{0}^{1}({\bm{x}}_{0})p_{0}^{2}({\bm{x}}_{0})$, where $p_{0}^{1}({\bm{x}}_{0})$ and $p_{0}^{2}({\bm{x}}_{0})$ are distributions of different diffusion models. Since we have access to the score functions $\nabla_{{\bm{x}}_{t}}\log q_{t}^{1}({\bm{x}}_{t})$ and $\nabla_{{\bm{x}}_{t}}\log q_{t}^{2}({\bm{x}}_{t})$ through the diffusion model, we can construct a sequence of annealing distributions $\tilde{q}_{t}^{\text{prod}}({\bm{x}}_{t})$ such that:
$\nabla_{{\bm{x}}_{t}}\log\tilde{q}_{t}^{\text{prod}}({\bm{x}}_{t})=\nabla_{{\bm{x}}_{t}}\log q_{t}^{1}({\bm{x}}_{t})+\nabla_{{\bm{x}}_{t}}\log q_{t}^{2}({\bm{x}}_{t})\,.$
By sampling from the sequence $\left\{\tilde{q}_{t}^{\text{prod}}({\bm{x}}_{t})\right\}$, we can arrive at $\tilde{q}_{0}^{\text{prod}}({\bm{x}}_{0})$ which is equal to $p^{\text{prod}}_{0}({\bm{x}}_{0})$. A key difference from sampling from $\left\{\tilde{q}_{t}^{\text{prod}}({\bm{x}}_{t})\right\}$ and direct diffusion sampling is that the diffusion process with $p^{\text{prod}}_{0}({\bm{x}}_{0})$ defined as
$q_{t}^{\text{prod}}({\bm{x}}_{t})=\int_{{\bm{x}}_{0}}p^{\text{prod}}_{0}({\bm{x}}_{0})q({\bm{x}}_{t}|{\bm{x}}_{0})$
is different from $\tilde{q}_{t}^{\text{prod}}({\bm{x}}_{t})$. The score of $q_{t}^{\text{prod}}({\bm{x}}_{t})$ can be derived as:
$\nabla_{{\bm{x}}_{t}}\log q_{t}^{\text{prod}}({\bm{x}}_{t})=\nabla_{{\bm{x}}_{t}}\log\left(\int_{{\bm{x}}_{0}}p_{0}^{1}({\bm{x}}_{0})p_{0}^{2}({\bm{x}}_{0})q({\bm{x}}_{t}|{\bm{x}}_{0})\right)\,,$
which is not equal to
$\nabla_{{\bm{x}}_{t}}\log\tilde{q}_{t}^{\text{prod}}({\bm{x}}_{t})=\nabla_{{\bm{x}}_{t}}\log\left(\int_{{\bm{x}}_{0}}p_{0}^{1}({\bm{x}}_{0})q({\bm{x}}_{t}|{\bm{x}}_{0})\right)+\nabla_{{\bm{x}}_{t}}\log\left(\int_{{\bm{x}}_{0}}p_{0}^{2}({\bm{x}}_{0})q({\bm{x}}_{t}|{\bm{x}}_{0})\right)\,,$
and thus intractable to compute directly.
A key distinction between annealed Langevin MCMC sampling and reverse diffusion sampling is that we run multiple Langevin MCMC steps on the same noise level, while reverse diffusion goes from high noise level to low noise level via denoising. A minimal pseudo code is shown in Alg. 1.
Algorithm 1 Annealed Langevin MCMC sampling
  Input: sequence of annealing distributions $\left\{\tilde{q}_{t}({\bm{x}}_{t})\right\}_{t=0}^{T}$, number of MCMC steps $N$, step size $\left\{\eta_{t}\right\}_{t=0}^{T}$. (Optional) reverse transition kernel $\left\{\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})\right\}_{t=0}^{T}$.
  Init: ${\bm{x}}_{T}^{0}\sim\mathcal{N}(\bm{0},\bm{I})$
  for $t=T,\cdots,1$ do
     for $i=0,1,\cdots,N-1$ do
        Perform Langevin MCMC steps:
${\bm{x}}_{t}^{i+1}={\bm{x}}_{t}^{i}+\eta_{t}\nabla_{{\bm{x}}_{t}}\log\tilde{q}_{t}({\bm{x}}_{t}^{i})+\sqrt{2\eta_{t}}{\bm{\epsilon}}_{t}^{i}\,,~{}~{}{\bm{\epsilon}}_{t}^{i}\sim\mathcal{N}(\bm{0},\bm{I})\,.$
     end for
     (Optional) transit to next time step: ${\bm{x}}_{t-1}^{0}\sim\tilde{p}_{\theta}(\cdot|{\bm{x}}_{t}^{N})$. If no reverse kernel initialize ${\bm{x}}_{t-1}^{0}={\bm{x}}_{t}^{N}$.
  end for
  Return ${\bm{x}}_{0}$
C.3 Annealed Langevin MCMC with recurrent training-free guidance
In this section, we prove the connection between annealed Langevin MCMC (Alg. 1) and training-free guidance (Alg. 2) in Proposition. 1. We divide the proof into two parts. In Sec. C.3.1 we prove the equivalence between naive recurrence steps and Langevin MCMC. Then in Sec. C.3.2, we prove that adding the guidance term is defining an annealing path that biases towards high score regions.
C.3.1 Equivalence between Langevin MCMC and naive recurrence
Consider the diffusion process with the following stochastic interpolant [47]:
${\bm{x}}_{t}=\alpha_{t}{\bm{x}}_{0}+\sigma_{t}{\bm{\epsilon}}\,.$
We denote the score function of $q_{t}({\bm{x}}_{t})$ as $\nabla_{{\bm{x}}_{t}}\log q_{t}({\bm{x}}_{t})=s({\bm{x}}_{t},t)$. Recall the forward process in Eq. 1:
${\bm{x}}_{t}=\frac{\alpha_{t}}{\alpha_{t-1}}{\bm{x}}_{t-1}+\sqrt{\alpha_{t}^{2}\left(\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}-\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}}\right)}{\bm{\epsilon}}\,.$ (9)
In a recurrence step in Line. 5, we first solve ${\bm{x}}_{t-1}^{i}$ from ${\bm{x}}_{t}^{i}$ using the learned score function $s({\bm{x}}_{t}^{i},t)$, then add noise to ${\bm{x}}_{t-1}^{i}$ to obtain the recurrent sample ${\bm{x}}_{t}^{i+1}$, where the superscript denotes the recurrence step index: $i=0,1,\cdots,N_{\text{recur}}$. Depending on different solvers, we have different formulations of ${\bm{x}}_{t}^{i+1}$.
DDIM sampler
When using DDIM [65] sampler, we have the reverse step as:
${\bm{x}}_{t-1}=\frac{\alpha_{t-1}}{\alpha_{t}}{\bm{x}}_{t}+\sigma_{t}^{2}\left(\frac{\alpha_{t-1}}{\alpha_{t}}-\frac{\sigma_{t-1}}{\sigma_{t}}\right)s({\bm{x}}_{t},t)\,,$ (10)
where $s({\bm{x}}_{t},t)$ is the score function $\nabla_{{\bm{x}}_{t}}\log q_{t}({\bm{x}}_{t})$. Thus, we have:
${\bm{x}}_{t}^{i+1}$ $=\frac{\alpha_{t}}{\alpha_{t-1}}{\bm{x}}_{t-1}^{i}+\alpha_{t}\sqrt{\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}-\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}}}{\bm{\epsilon}}^{i}$
$={\bm{x}}_{t}^{i}+\sigma_{t}^{2}\left(1-\frac{\alpha_{t}}{\alpha_{t-1}}\frac{\sigma_{t-1}}{\sigma_{t}}\right)s({\bm{x}}_{t}^{i},t)+\sigma_{t}\sqrt{1-\frac{\alpha_{t}^{2}}{\alpha_{t-1}^{2}}\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}}{\bm{\epsilon}}^{i}\,.$
Denote $\lambda_{t}=\log\frac{\alpha_{t}}{\sigma_{t}}$, then we have:
${\bm{x}}_{t}^{i+1}$ $={\bm{x}}_{t}^{i}+\sigma_{t}^{2}\left(1-e^{\lambda_{t}-\lambda_{t-1}}\right)s({\bm{x}}_{t}^{i},t)+\sigma_{t}\sqrt{1-e^{2(\lambda_{t}-\lambda_{t-1})}}{\bm{\epsilon}}^{i}$
$={\bm{x}}_{t}^{i}+\sigma_{t}^{2}\left(1-e^{\lambda_{t}-\lambda_{t-1}}\right)s({\bm{x}}_{t}^{i},t)+\sigma_{t}\sqrt{\left(1-e^{\lambda_{t}-\lambda_{t-1}}\right)\left(1+e^{\lambda_{t}-\lambda_{t-1}}\right)}{\bm{\epsilon}}^{i}\,,$
where $1+e^{\lambda_{t}-\lambda_{t-1}}\rightarrow 2$ when $T\rightarrow\infty$ and denoising step size approaches 00, as $\lambda_{t}-\lambda_{t-1}\rightarrow 0$.
DDPM sampler
In DDPM [26], we parametrize the posterior distribution as:
$p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})=\mathcal{N}\left({\bm{x}}_{t-1};\mu_{\theta}({\bm{x}}_{t},t),\Sigma_{\theta}({\bm{x}}_{t},t)\right)\,,$ (11)
where the posterior mean is:
$\mu_{\theta}({\bm{x}}_{t},t)=\frac{\alpha_{t-1}}{\alpha_{t}}{\bm{x}}_{t}+\left(\sigma_{t}^{2}\frac{\alpha_{t-1}}{\alpha_{t}}-\sigma_{t-1}^{2}\frac{\alpha_{t}}{\alpha_{t-1}}\right)s({\bm{x}}_{t},t)\,.$
[26] parameterizes the posterior variance as $\Sigma_{\theta}({\bm{x}}_{t},t)=\beta_{t}\bm{I}$ or $\Sigma_{\theta}({\bm{x}}_{t},t)=\tilde{\beta}_{t}\bm{I}$:
$\beta_{t}$ $=\alpha_{t}^{2}\left(\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}-\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}}\right)\,,$
$\tilde{\beta}_{t}$ $=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\beta_{t}\,,$
while [54] propose to train the posterior variance as $\Sigma_{\theta}({\bm{x}}_{t},t)=\exp\left(v\log\beta_{t}+(1-v)\log\tilde{\beta}_{t}\right)$.
Thus, a backward step can be written as:
${\bm{x}}_{t-1}=\frac{\alpha_{t-1}}{\alpha_{t}}{\bm{x}}_{t}+\left(\sigma_{t}^{2}\frac{\alpha_{t-1}}{\alpha_{t}}-\sigma_{t-1}^{2}\frac{\alpha_{t}}{\alpha_{t-1}}\right)s({\bm{x}}_{t},t)+\Sigma_{\theta}^{1/2}({\bm{x}}_{t},t){\bm{\epsilon}}_{\text{post}}\,,$
where ${\bm{\epsilon}}_{\text{post}}$ denotes the noise added in the posterior sampling step. Then, we can write the recurrence step as:
${\bm{x}}_{t}^{i+1}$ $={\bm{x}}_{t}^{i}+\left(\sigma_{t}^{2}-\sigma_{t-1}^{2}\frac{\alpha_{t}^{2}}{\alpha_{t-1}^{2}}\right)s({\bm{x}}_{t}^{i},t)+\Sigma_{\theta}^{1/2}({\bm{x}}_{t}^{i},t){\bm{\epsilon}}_{\text{post}}^{i}+\alpha_{t}\sqrt{\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}-\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}}}{\bm{\epsilon}}^{i}_{\text{forward}}$
$={\bm{x}}_{t}^{i}+\beta_{t}s({\bm{x}}_{t}^{i},t)+\sqrt{\Sigma_{\theta}({\bm{x}}_{t},t)+\beta_{t}\bm{I}}{\bm{\epsilon}}^{i}\,,$
where $\Sigma_{\theta}({\bm{x}}_{t},t)\rightarrow\beta_{t}\bm{I}$ when $T\rightarrow\infty$, and the denoising step size approaches 00.
Putting together
In general, we can write the recurrence step as:
${\bm{x}}_{t}^{i+1}={\bm{x}}_{t}^{i}+a_{t}s({\bm{x}}_{t}^{i},t)+b_{t}{\bm{\epsilon}}^{i}\,.$ (12)
with $a_{t}\rightarrow 0$, $b_{t}\rightarrow 0$ and $2a_{t}\rightarrow b_{t}^{2}$ as the denoising step size approaches 00. Thus, it can be seen as a approximation of the ULA in Eq. 8, and also a discretization of the Langevin SDE in Eq. 7.
C.3.2 Annealed Langevin MCMC with guidance
When applying training free guidance [94] during the recurrence, we have:
${\bm{x}}_{t}^{i+1}={\bm{x}}_{t}^{i}+a_{t}s({\bm{x}}_{t}^{i},t)+b_{t}{\bm{\epsilon}}^{i}+\bm{\Delta}({\bm{x}}_{t},t)\,,$
where $a_{t}$, $b_{t}$ are the coefficients of the recurrence equation in Eq. 12 without guidance. In general, $\bm{\Delta_{t}}=\rho_{t}\nabla_{{\bm{x}}_{t}}\log f({\bm{x}}_{0|t})+\mu_{t}\alpha_{t}\nabla_{{\bm{x}}_{0|t}}\log f({\bm{x}}_{0|t})$, where $\rho_{t},\mu_{t}$ controls the guidance strength. We then show that the guidance term can be considered as the score function of a set of annealed verifiers $\left\{\hat{f}({\bm{x}}_{t})\right\}_{t=0}^{T}$.
When considering ‘variance guidance’ in Line. 7, we have $\bm{\Delta}_{\text{var}}=\rho_{t}\nabla_{{\bm{x}}_{t}}\log f({\bm{x}}_{0|t})$. Thus, we can define $\hat{f}_{t}^{\text{var}}({\bm{x}}_{t})=f({\bm{x}}_{0|t})$, which satisfies $\hat{f}_{0}^{\text{var}}({\bm{x}}_{0})=f({\bm{x}}_{0})$. Similarly, for ‘mean guidance’ in Line. 8, we have
$\bm{\Delta}_{\text{mean}}$ $=\mu_{t}\alpha_{t}\nabla_{{\bm{x}}_{0|t}}\log f({\bm{x}}_{0|t})$
$=\mu_{t}\frac{\sigma_{t}^{2}}{\bm{\Sigma}_{0|t}}\nabla_{{\bm{x}}_{t}}\log f({\bm{x}}_{0|t})\,,$
where the second Equation follows from Lemma 3.3 in [94]. Thus, there exists a set of functions $\hat{f}_{t}^{\text{mean}}({\bm{x}}_{t})$ such that $\nabla_{{\bm{x}}_{t}}\log\hat{f}_{t}^{\text{mean}}({\bm{x}}_{t})=\frac{\sigma_{t}^{2}}{\bm{\Sigma}_{0|t}}\nabla_{{\bm{x}}_{t}}\log f({\bm{x}}_{0|t})$, and we can see that when $t\rightarrow 0$,$\nabla_{{\bm{x}}_{t}}\log\hat{f}_{t}^{\text{mean}}({\bm{x}}_{t})=\nabla_{{\bm{x}}_{0}}\log f({\bm{x}}_{0})$. If we additionally incorporate the ‘implicit dynamics’ in Line. 4, our arguments still stands since the smoothed objective $\tilde{f}({\bm{x}})=\mathbb{E}_{\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})}f({\bm{x}}+\bar{\gamma}{\sigma_{t}}\bm{\delta})$ converges to $f$ with $t\rightarrow 0$ and $\sigma_{t}\rightarrow 0$.
Combining the two terms together, we have $\bm{\Delta}_{t}=c_{t}\nabla_{{\bm{x}}_{t}}\log\hat{f}_{t}({\bm{x}}_{t})$ with $\hat{f}_{t}=\hat{f}_{t}^{\text{var}}\cdot\hat{f}_{t}^{\text{mean}}$. Thus, recurrence with guidance can be written as:
${\bm{x}}_{t}^{i+1}$ $={\bm{x}}_{t}^{i}+a_{t}s({\bm{x}}_{t}^{i},t)+b_{t}{\bm{\epsilon}}^{i}+c_{t}\nabla_{{\bm{x}}_{t}}\log\hat{f}_{t}({\bm{x}}_{t})\,,$
which, in the continuous limit of $a_{t},b_{t},c_{t}\rightarrow 0$, converges to the Langevin SDE:
$d{\bm{x}}_{t}^{i}=\nabla_{{\bm{x}}_{t}^{i}}\log\left(q_{t}({\bm{x}}_{t}^{i})\hat{f}_{t}({\bm{x}}_{t}^{i})\right)di+\sqrt{2}d\bm{w}^{i}\,,$
where $i$ is the time index of recurrence steps. Thus, we have defined the annealing path as $\tilde{q}_{t}({\bm{x}}_{t})=q_{t}({\bm{x}}_{t})\hat{f}_{t}({\bm{x}}_{t})\,,~{}t=1,2,\cdots,T$.
C.4 Relationship between Langevin MCMC and gradient ascent
In training-free guidance, most prior works only apply gradient ascent without recurrence. Here we provide a theoretical analysis of both methods.
Recall the KL-divergence objective in Eq. 5, which can be further decomposed when we are sampling from a compositional distribution of $p_{0}({\bm{x}}_{0})$ and verifier $f({\bm{x}}_{0})$, with $\nu\propto p_{0}\cdot f$:
$H_{\nu}(\rho)=\mathbb{E}_{\rho}[-\log f]+H_{p_{0}}(\rho)+\log Z\,.$
where $Z=\int p_{0}f$ is a normalization constant. Thus, gradient ascent is optimizing the verifier objective $\mathbb{E}_{\rho}[-\log f]$, while Langevin MCMC in Eq. 12 is optimizing the divergence between current sample and base distribution $H_{p_{0}}(\rho)$. This explains why naive gradient updates leads to OOD samples, and recurrence effectively mitigates this issue, acting as a contraction force pulling the sample back to the original manifold. However, since we start from $p_{0}$ as the distribution of our initial sample, sometimes we can omit the recurrence if the guidance strength is small. But if we wish to traverse different modes with multiple gradient updates, introducing recurrence helps to avoid OOD during optimization.
C.5 Implementing Local Search with TFG hyper-parameter space
Due to the equivalence between annealed Langevin MCMC and training-free guidance with recurrence, we can implement local search with Langevin MCMC using the TFG framework of [94], efficiently searching the hyperparameters. Here we provide a overview of the algorithm and design space. Following Sec. C.3, every iteration of recurrence in Line. 5 is equivalent to an annealed Langevin MCMC step, thus $N_{\text{recur}}$ is equal to the number of local search steps.
Algorithm 2 Training-Free Guidance
1:  Input: Unconditional diffusion model ${\bm{\epsilon}}_{\theta}$, verifier $f$, guidance strength $\bm{\rho},\bm{\mu},\bar{\gamma}$, number of steps $T,N_{\text{recur}},N_{\text{iter}}$
2:  ${\bm{x}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
3:  for $t=T,\cdots,1$ do
4:     Define function $\tilde{f}({\bm{x}})=\mathbb{E}_{\bm{\delta}\sim\mathcal{N}(\bm{0},\bm{I})}f({\bm{x}}+\bar{\gamma}{\sigma_{t}}\bm{\delta})$
5:     for $r=1,\cdots,N_{\text{recur}}$ do
6:        ${\bm{x}}_{0|t}=({\bm{x}}_{t}-\sigma_{t}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t))/\alpha_{t}$
7:        $\bm{\Delta}_{\text{var}}=\rho_{t}\nabla_{{\bm{x}}_{t}}\log\tilde{f}({\bm{x}}_{0|t})$
8:        $\bm{\Delta}_{\text{mean}}=\bm{\Delta}_{\text{mean}}+\mu_{t}\alpha_{t}\nabla_{{\bm{x}}_{0|t}}\log\tilde{f}({\bm{x}}_{0|t}+\bm{\Delta}_{\text{mean}})$trianglerightIterate $N_{\text{iter}}$ times starting from $\bm{\Delta}_{\text{mean}}=\bm{0}$
9:        ${\bm{x}}_{t-1}=\text{Sample}({\bm{x}}_{t},{\bm{x}}_{0|t},t)+\frac{\alpha_{t-1}}{\alpha_{t}}\left(\bm{\Delta}_{\text{var}}+\bm{\Delta}_{\text{mean}}\right)$triangleright Sample follows DDIM or DDPM
10:        ${\bm{x}}_{t}\sim\mathcal{N}\left(\frac{\alpha_{t}}{\alpha_{t-1}}{\bm{x}}_{t-1},\alpha_{t}^{2}\left(\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}-\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}}\right)\bm{I}\right)$triangleright Recurrent strategy
11:     end for
12:  end for
13:  Output: Conditional sample ${\bm{x}}_{0}$
For time varying schedules $\rho_{t}$, $\mu_{t}$, we follow [94] and propose to use either the ‘increase’ schedule:
$s_{t}=T\frac{\alpha_{t}/\alpha_{t-1}}{\sum_{t=1}^{T}\alpha_{t}/\alpha_{t-1}}\,,$ (13)
where we increase the guidance strength as we denoise: $s_{T}<s_{T-1}<\cdots<s_{1}$; or the ‘constant’ schedule
$s_{t}=1\,,$ (14)
which uses constant parameters throughout the denoising process. Thus, the time-varying schedules can be computed as $\rho_{t}=s_{t}\bar{\rho}$ and $\mu_{t}=s_{t}\bar{\mu}$, and we only need to determine the average $\bar{\rho}$ and $\bar{\mu}$.
Appendix D Global Search of Denoising Diffusion Models
In this section, we provide details about the global search algorithms: BFS and DFS.
D.1 BFS-Based Search
We present the pseudo code for BFS in Alg. 3. The core hyperparameters of BFS are the temperature $\tau_{t}$ and the evaluation time steps $\mathcal{S}$. These parameters enable a smooth transition from Best-of-N to best-first search. In general, a higher $\tau_{t}$ and evaluation at earlier time steps bias the process toward high-score samples, while a lower temperature and evaluation at later (smaller) time steps approximate the Best-of-N approach.
In practice, we typically evaluate at fractions of the sampling steps, with $\mathcal{S}=\left\{\frac{3}{4}T,\frac{1}{2}T,\frac{1}{4}T\right\}$. We perform a grid search over the temperature values to determine an appropriate magnitude using a small number of particles, and then apply this configuration when scaling up the number of particles. We use the ‘increase’ schedule in Eq. 13 for $\tau_{t}$, so that $\tau_{T}\leq\tau_{T-1}\leq\cdots\leq\tau_{0}$, reflecting increased confidence in the verifier evaluation at smaller time steps.
Algorithm 3 Diffusion BFS
  Diffusion input: diffusion model ${\bm{\epsilon}}_{\theta}$ with diffusion time steps $T$ and proposal transition kernel $\left\{\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})\right\}_{t=1}^{N}$. Verifier $f$.
  BFS input: Set of evaluation time steps $\mathcal{S}$. Temperature schedule $\tau_{t}$. Budget of particles $K$.
  Init: Random sample $K$ particles ${\bm{x}}_{T}^{k}\sim\mathcal{N}(\bm{0},\bm{I}),k=1,2,\cdots K$. Set the number of particles $K_{T}=K$
  for $t=T,\cdots,1$ do
     if $t\in\mathcal{S}$ then
        for $k=1,2,\cdots,K_{t}$ do
           Estimate the conditional mean: ${\bm{x}}_{0|t}^{k}=\frac{{\bm{x}}_{t}^{k}-\sigma_{t}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t}^{k},t)}{\alpha_{t}}$. Compute the verifier score $f({\bm{x}}_{0|t}^{k})$.
           Compute the number of children $n_{t}^{k}$:
           if Resample then
              $n_{t}^{k}=\texttt{Round}\left(K_{t}\frac{f({\bm{x}}_{0|t}^{k})^{\tau_{t}}}{\sum_{i=1}^{K_{t}}f({\bm{x}}_{0|t}^{i})^{\tau_{t}}}\right)$
           else if Prune then
              $n_{t}^{k}=\min\left(1,\texttt{Round}\left(K_{t}\frac{f({\bm{x}}_{0|t}^{k})^{\tau_{t}}}{\sum_{i=1}^{K_{t}}f({\bm{x}}_{0|t}^{i})^{\tau_{t}}}\right)\right)$
           end if
        end for
     else
        $n_{t}^{k}=1$
     end if
     for $k=1,\cdots,K_{t}$ do
        Sample $n_{t}^{k}$ particles from ${\bm{x}}_{t}^{k}$: ${\bm{x}}_{t-1}^{j}\sim\tilde{p}_{\theta}(\cdot|{\bm{x}}_{t}^{k})\,,j=1,2,\cdots,n_{t}^{k}$
     end for
     The number of particles at next noise level: $K_{t-1}=\sum_{k=1}^{K_{t}}n_{t}^{k}$
  end for
  Return ${\bm{x}}_{0}=\operatorname*{argmax}_{k=1,\cdots,K_{0}}f({\bm{x}}_{0}^{k})$
D.2 DFS-based search
In this section, we provide the details and pseudo code for DFS in Alg. 4. To better utilize previously explored sampling paths, we employ a buffer to store prior results. When no particles pass the threshold constraint, we retrieve the best sample from the buffer.
Similar to BFS, controlling the set of evaluation steps allows a trade-off between efficiency and accuracy. Evaluating at earlier time steps introduces higher uncertainty but enables earlier resampling. Additionally, adjusting the backtracking depth $\Delta_{T}$ governs the search scope: a small $\Delta_{T}$ reduces computation and favors local search, while a larger $\Delta_{T}$ enables broader exploration at the cost of increased computation.
In practice, we set the evaluation steps to $\mathcal{S}=\left\{\frac{1}{2}T,\frac{1}{4}T\right\}$ for image experiments to save compute, and to $\mathcal{S}=\left\{\frac{3}{4}T,\frac{3}{4}T-1,\cdots,1\right\}$ for PointMaze experiments. We set the recurrence depth to $T/2$ for image tasks and $T/4$ for PointMaze, corresponding to the denoised steps at which samples are first evaluated. The threshold schedule $\delta_{t}$ is also set to ‘increase’ as in Eq.13, enforcing tighter constraints for samples with lower noise.
In our experiments, we observed that when backtracking to $t_{\text{next}}=T$—thus fully restarting—the nonzero terminal SNR $\alpha_{T}/\sigma_{T}$ in many diffusion schedules [40] can lead to cumulative errors with repeated backtracking. Therefore, when backtracking to $t_{\text{next}}=T$, we initialize with fresh Gaussian noise.
Algorithm 4 Diffusion DFS
  Diffusion input: diffusion model ${\bm{\epsilon}}_{\theta}$ with diffusion time steps $T$ and proposal transition kernel $\left\{\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})\right\}_{t=1}^{N}$. Verifier $f$.
  DFS input Budget for total number of backtracking $B=K$, backtracking depth $\Delta_{T}$ and threshold $\left\{\delta_{t}\right\}_{t=1}^{T}$. Set of evaluation time steps $\mathcal{S}$.
  Init ${\bm{x}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$, $t=T$. Init buffer with empty sets: $\texttt{buffer}(t)\leftarrow\left\{\right\}\,,t=1,2,\cdots,T$.
  while $t>0$ do
     if $t\in\mathcal{S}$ then
        Estimate the conditional mean and verifier score: ${\bm{x}}_{0|t}=\frac{{\bm{x}}_{t}-\sigma_{t}\epsilon_{\theta}({\bm{x}}_{t},t)}{\alpha_{t}}$, $f({\bm{x}}_{0|t})$.
        if $f({\bm{x}}_{0|t})<\delta_{t}$ and budget $B>0$ then
           Backtrack: $t_{\text{next}}\leftarrow\min(t+\Delta_{T},T)$, ${\bm{x}}_{t_{\text{next}}}\sim q({\bm{x}}_{t_{\text{next}}}|{\bm{x}}_{t})$ with $q$ in Eq. 1
           Decrease the budget: $B\leftarrow B-1$
           Add the score-value pair to the buffer: $\texttt{buffer}(t).\texttt{add}\left(f({\bm{x}}_{0|t}):{\bm{x}}_{t}\right)$
        else
           if $f({\bm{x}}_{0|t})<\delta_{t}$ and budget $B=0$ then
              Pop the best sample from buffer: ${\bm{x}}_{t}\leftarrow\texttt{buffer}(t).\texttt{max}$triangleright select the best sample from past explorations
           end if
           Sample posterior: $t_{\text{next}}\leftarrow t-1$, ${\bm{x}}_{t_{\text{next}}}\sim\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$
        end if
     else
        Sample posterior: $t_{\text{next}}\leftarrow t-1$, ${\bm{x}}_{t_{\text{next}}}\sim\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$
     end if
     $t\leftarrow t_{\text{next}}$, ${\bm{x}}_{t}\leftarrow{\bm{x}}_{\text{next}}$
  end while
  Return ${\bm{x}}_{0}$
Appendix E Experiment Details
In this section we provide the details of experimental setup and implementation for all our experiments. We run our experiments on clusters with Nvidia A100 GPUs, with over 1000 GPU hours used.
E.1 Long Horizon Maze Planning
Figure 7: Additional examples for PointMaze Ultra Fig. 7(a) shows tha maze layout and a example trajectory rollout. Fig. 7(b) shows the failure case of model generated plan. Fig. 7(c) shows a successful generated plan after inference-time scaling.
Maze environment
For all our maze experiments we use the OGBench PointMaze [55]. PointMaze Giant was directly adapted from their datasets, and we create PointMaze Ultra using the same protocol of Figure 5 in [49] [^1] https://github.com/mpetersen94/gcs/blob/main/models/maze.py, but with a smaller size of 20x20 cells. Dataset for Ultra is collected following the protocal in OGBench [55]. A larger layout and example trajectory of PointMaze Ultra is shown in Fig. 7. We evaluate the model on the default task 1 of OGBench [55], which is navigating from bottom left to top right. Empirically we discover that the diffusion model can perform well on short-horizon tasks without extra inference compute, but struggles heavily in the long horizon tasks.
Model Training
We train the model following diffuser [31], where we use a temporal U-Net to denoise the trajectory. Since our objective start and goal is more distant than trajectories in dataset, we sample at longer horizons than training, which is enabled by the temporal U-Net architecture. Specifically, for PointMaze Giant, we train at horizon 400 and sample at 600; and for PointMaze Ultra, we train at horizon 2000 and sample at 2800. For the PointMaze Ultra model we extend the depth of the U-Net for 1 more layer, and keep all other parameters the same. We train the model for 1.2M steps using the same configuration as [31].
Inference
We found that the model performance saturates with 16 denoising steps, which we use for all our experiments. For all the data points we report the average success rate with over 40 samples.
For verifier design, we use the ground-truth maze layout, and calculate the violation of each point in the trajectory using the position coordinates. Specifically, if a point $(x,y)$ is inside a maze wall box with center $(c_{x},c_{y})$ and half-width $d$, then the point loss can be calculated as the minimum distance from the point to box walls:
$L(x,y)=\min\left(x-(c_{x}-d),(c_{x}+d)-x,y-(c_{y}-d),(c_{y}+d)-y\right)\,.$
and the total verifier score is computed as:
$f(\bm{\tau})=\exp\left(-\sum_{i=1}^{H}L(x_{i},y_{i})^{2}\right)\,.$
So if all the points are free of violation in the trajectory, then $f(\bm{\tau})=1$. We point out that this does not indicate a successful plan as the connection between consecutive points $(x_{i},y_{i})\rightarrow(x_{i+1},y_{i+1})$ may violate the maze layout, and using only the verifier function can not generate a successful plan.
For local search, we search the hyper-parameters $\bar{\rho}$ and $\bar{\mu}$ in Sec. C.5 with $\bar{\gamma}=0$, and we set the local search steps (recurrence steps) to $2$ for Giant and $6$ for Ultra. We set $N_{\text{iter}}=1$ as we observe they do not contribute to the performance gains, while the number of Langevin MCMC steps has the most impact on performance. For global search with BFS we evaluate at steps $\left\{12,8,4\right\}$, and for DFS we evaluate at $\left\{12,11,\cdots,1\right\}$ with backtracking depth $\Delta_{T}=4$. We also observe that increasing backtracking depth to $12$ and evaluate at smaller time steps $\left\{4,3,2,1\right\}$ helps to scale up the performance with more compute. We use grid search to determine temperature and threshold, with a step size of $[1,5,10]$ for each magnitude.
E.2 Offline RL
Setup
We follow the setup in [44], and we directly use their pretrained diffusion model and Q-function, omitting the time-dependent energy function. The diffusion model was trained to generate action $\bm{a}$ given state $\bm{s}$, and we sample with 15 steps of DDIM.
For hyper-parameter search, we disable the implicit dynamics and set $\bar{\gamma}=0$, and use the ‘increase’ schedule for $\bm{\rho}$ and $\bm{\mu}$. For strength parameters $\bar{\rho}$ and $\bar{\mu}$, we first search for the right magnitude. Then, we also follow [44] and search with step size [1,2,3,5,8,10] within the magnitude. Same as [44], we use 5 different seeds with 10 samples per seed for each task. To avoid over fitting, we use different seeds for parameter search and evaluation. We report the hyper-parameters and the performance within the parameter-searching dataset and evaluation dataset.
For global search, we use 4 particles for Medium-Expert and Medium datasets, and 2 particles for Medium-Replay datasets. Since the number of particles are small, we do not carry out BFS or DFS methods and simply use Best-of-N. We point out that the number of particles we use are much smaller than the 50 particles in [80] and the 32 particles in [8], highlighting the effectiveness of local search.
Baseline
We compare our method to a variety of baselines, including traditional state-of-the-art methods like CQL [35], BCQ [20] and IQL [34], diffusion-based policies such as diffuser [31], decision-diffuser (DD) [1], Diffusion-QL (D-QL) [80], SfBS [8] and QGPO [44]. We directly take the numbers from [44].
Among the baseline diffusion-based methods, both Diffuser [31] and QGPO [44] requires training a noise-dependent guidance function, and D-QL [80] requires updating the diffusion model during training using the Q-function iteratively, which needs to back-propagate through the diffusion sampling chain, introducing high computation and memory overheads. DD [1] uses classifier-free guidance [27] to generate high-return trajectories, which requires training a return-conditional model on labeled datasets, which can be expensive to obtain in robotics where only demonstration data is available [43].
Dataset Environment particles $N_{\text{recur}}$ $N_{\text{iter}}$ $\bar{\rho}$ $\bar{\mu}$ Eval set Search set
Medium-Expert HalfCheetah 4 1 1 0.008 0.02 $93.9\pm 0.3$ $\mathbf{94.3\pm 0.5}$
Medium-Expert Hopper 4 1 4 0.001 0.00 $104.4\pm 3.1$ $\mathbf{109.4\pm 5.2}$
Medium-Expert Walker2d 4 1 1 0.005 0.10 $\mathbf{111.4\pm 0.1}$ $111.4\pm 0.2$
Medium HalfCheetah 4 1 4 0.003 0.05 $\mathbf{54.8\pm 0.1}$ $54.8\pm 0.2$
Medium Hopper 4 4 4 0.003 0.02 $99.5\pm 1.7$ $\mathbf{100.1\pm 0.1}$
Medium Walker2d 4 1 6 0.003 0.08 $\mathbf{86.5\pm 0.2}$ $85.2\pm 3.2$
Medium-Replay HalfCheetah 2 1 6 0.005 0.03 $47.8\pm 0.4$ $\mathbf{48.4\pm 0.1}$
Medium-Replay Hopper 2 1 1 0.003 0.20 $97.4\pm 4.0$ $\mathbf{100.4\pm 2.2}$
Medium-Replay Walker2d 2 2 4 0.003 0.03 $79.3\pm 9.7$ $\mathbf{83.2\pm 2.8}$
Average $86.1$ $\mathbf{87.5}$
Table 3: Hyper-parameters on D4RL locomotion tasks with test-time scaling. We report the performance on hyper-parameter search dataset and the evaluation dataset, highlighting the best number.
E.3 Image Generation
E.3.1 Text-to-Image Compositional Generation
Setup
We use the SSD-1B model [^2] https://huggingface.co/segmind/SSD-1B which is distilled from SDXL, and we use the default sampling configuration with 50 steps of DDIM sampler. We use the oracle verifier of each task following CompBench [29].
Global search with DFS
For DFS we evaluate at time steps $\left\{25,35,45\right\}$ and set the backtrack depth $\Delta_{T}=25$. We set the threshold for our experiments to $0.5$, $0.7$ and $0.9$, with the ‘constant’ schedule in Eq. 14. We typically use a higher threshold for high computation budgets, representing using more compute for higher quality standard.
We provide example visualizations in Fig. 8.
(a) 1 Particle
(b) 2 Particles
(c) 4 Particles
(d) 8 Particles
Figure 8: Example of inference-time scaling with Numeracy dataset Above shows the results of Best-of-N with different particles, with prompt: “four dogs played with two toys”
E.3.2 ImageNet Conditional Generation
Setup
We use the 256x256 DDPM ImageNet model from [12], and for local search we use the 224px ViT classifier [^3] https://huggingface.co/google/vit-base-patch16-224 and for global search we use the 384px version [^4] https://huggingface.co/google/vit-base-patch16-384 when using double-verifier. We evaluate on class label 222 and 444, with 256 samples per label. We sample with 100 DDPM steps. When evaluating the class accuracy we use the DeiT model [^5] https://huggingface.co/facebook/deit-small-patch16-224, and we evaluate the FID againt the images of the corresponding class in ImageNet.
Global search with BFS
Due to computation constraints we set the local search steps to $1$ in all experiments. For BFS we set the evaluation steps as $\left\{75,50,25\right\}$, and use $\bar{\tau}=1.0$ for label 222 and $\bar{\tau}=0.5$ for label 444 with the ‘increase’ schedule.
We provide example visualizations in Fig. 9.
(a) 2 Particles
(b) 4 Particles
(c) 8 Particles
(d) 12 Particles
Figure 9: Visualizations of ImageNet Conditional Generation Above shows the uncurated samples for class label 222 with BFS with different number of particles

### v2

@@
 We propose a general framework that orchestrates local and global search to efficiently navigate the generative space.
-It employs a theoretically grounded local search via annealed Langevin MCMC and performs compute-efficient global exploration using breadth-first and depth-first tree search.
-We evaluate our approach on a range of challenging domains, including planning, offline reinforcement learning, and image generation.
-Across all tasks, we observe significant gains in both performance and efficiency.
-These results show that classical search provides a principled and practical foundation for inference-time scaling in diffusion models.
+It performs compute-efficient global exploration using breadth-first and depth-first tree search and employs a theoretically grounded, scalable local search via annealed Langevin MCMC.
+We evaluate our approach on a range of challenging domains, including planning, offline reinforcement learning, and image generation, and observe significant gains in both performance and efficiency over baseline methods.
+These results demonstrate that classical search offers a principled and practical foundation for inference-time scaling in diffusion models.
+By jointly scaling local and global search for the first time, our framework establishes a new Pareto frontier across challenging decision-making domains.
 Project page at diffusion-inference-scaling.github.io.
 1 Introduction
-Classical search algorithms have laid the foundation for modern artificial intelligence 10.5555/1671238 .
+Classical search algorithms have laid the foundation for modern artificial intelligence (Russell & Norvig, 2009).
 In discrete settings, graph search algorithms are widely used to explore the state space.
-Breadth-first search (BFS) moore1959shortest and depth-first search (DFS) tarjan1972depth traverse the search tree in a fixed order.
-To better leverage problem-specific information, best-first search methods pearl1984heuristics , such as A* hart1968formal , use a heuristic to evaluate and prioritize states.
-Alternatively, local search methods, such as hill-climbing (10.5555/1671238, , Sec.
+Breadth-first search (BFS) (Moore, 1959) and depth-first search (DFS) (Tarjan, 1972) traverse the search tree in a fixed order.
+To better leverage problem-specific information, best-first search methods (Pearl, 1984), such as A* (Hart et al., 1968), use a heuristic to evaluate and prioritize states.
+Alternatively, local search methods, such as hill-climbing (Russell & Norvig, 2009, Sec.
 4.1), explore neighboring states.
 More recent techniques like gradient descent and Markov Chain Monte Carlo (MCMC) have become widely adopted in optimization and probabilistic inference, underpinning many modern AI models.
-Diffusion models ho2020denoising have shown impressive performance in generative modeling for continuous domains such as images dhariwal2021diffusion , videos ho2022video , and world modeling yanglearning .
-They are also increasingly used in robotics and decision-making liu2024rdt ; black2024pi_0 ; team2024octo to generate diverse actions chi2023diffusion .
-However, generated samples may not always align with physical laws song2023loss or human intent wallace2024diffusion , and the vast generative space often necessitates multiple trials to produce satisfactory outputs xie2025sana .
+Diffusion models (Ho et al., 2020) have shown impressive performance in generative modeling for continuous domains such as images (Dhariwal & Nichol, 2021), videos (Ho et al., 2022).
+They are also increasingly used in robotics and decision-making (Liu et al., 2024; Black et al., 2024; Team et al., 2024) to generate diverse actions (Chi et al., 2023).
+However, generated samples may not always align with physical laws (Song et al., 2023) or human intent (Wallace et al., 2024), and the vast generative space often necessitates multiple trials to produce satisfactory outputs (Xie et al., 2025).
 To address this, we scale up inference-time compute using strategic search methods that navigate the generative manifold for high-quality samples.
 We formalize sample evaluation using a verifier function $f({\bm{x}}_{0})$ defined on ground truth samples, which measures the quality of the sample.
-Such verifiers could be reward functions xu2023imagereward , Q-functions lu2023contrastive , classifier conditions $p(c|{\bm{x}}_{0})$ ye2024tfg ; dhariwal2021diffusion , and multi-modal LLMs huang2023t2icompbench .
-To efficiently search the generative space induced by diffusion, we revisit classical search principles.
-Inspired by local search, when the verifier $f$ is differentiable, we use Langevin MCMC to explore neighborhoods of current samples, guided by the verifier gradient and “score function” vincent2011score_matching from the diffusion model.
-To explore diverse modes in the complex distribution generated by diffusion models, we search the Markov chain of the “denoising process” with global search methods.
-Viewing sampling as a search tree, we use BFS and DFS to gradually explore the states while denoising.
-Like best-first search methods, we can evaluate intermediate states ${\bm{x}}_{t}$ using the verifier $f({\bm{x}}_{0|t})$, expanding high-quality paths more effectively.
-An illustration of local and global search is shown in Fig.
+Such verifiers could be reward functions (Xu et al., 2023a), Q-functions (Lu et al., 2023), classifier conditions $p(c|{\bm{x}}_{0})$ (Ye et al., 2024; Dhariwal & Nichol, 2021), and VLMs (Huang et al., 2023).
+To efficiently search the generative space of diffusion models, we revisit classical search principles.
+To capture diverse modes in the complex distributions generated by diffusion models, we view sampling as traversing a search tree, employing BFS and DFS to progressively explore states during denoising.
+Similar to best-first search, we evaluate intermediate states ${\bm{x}}_{t}$ with the verifier $f({\bm{x}}_{0|t})$, prioritizing high-quality paths and efficiently allocating compute via branching and backtracking.
+To go beyond the base model and obtain higher-quality samples, we perform local search via Langevin MCMC, exploring the neighborhood of current samples under guidance from both the verifier gradient and the diffusion model’s “score function” (Vincent, 2011).
+By jointly optimizing the compositional objective of the diffusion model and the verifier (Du et al., 2023), our local search surpasses the capabilities of the base model.
+An overview of our framework is shown in Fig.
 1.
-Recently, Ma et al.
-ma2025inference also explored inference-time scaling of diffusion models using search-based techniques, focusing on image generation tasks.
-Their zero-order search approach can be interpreted as a instance of local search, while their path-based strategies a combination of both local and global search.
-However, their method does not effectively leverage verifier gradients, and we introduce a local search strategy with annealed Langevin MCMC, efficiently steering sampling toward high-score regions.
-Moreover, we systematically evaluate the compute efficiency of search methods.
-Finally, we extend inference-time scaling beyond image generation to a broader set of domains.
-To this end, we summarize our key contributions as follows:
+Recent works scale diffusion model inference via particle-based SMC (Kim et al., 2025; Singhal et al., 2025a; Wu et al., 2023) and tree-based methods (Guo et al., 2025), typically as BFS with fixed schedules.
+We generalize these with a BFS-based framework, clarifying prior design choices and providing an improved BFS surpassing prior methods.
+Inspired by DFS, we add adaptive backtracking to allocate compute adaptively, surpassing BFS in efficiency and adaptivity.
+While global search remains limited to base-distribution modes, scaling local search with Langevin MCMC explores high-reward regions beyond the model, proving effective in challenging decision-making tasks.
+Finally, we jointly scale local and global search with the classical AI search principles, demonstrating superior performance over prior methods, which scale local and global search in isolation.
+Our key contributions are summarized as follows:
   • i)
-    We propose a general and unified framework for efficient inference-time search in diffusion models, drawing inspiration from classical search algorithms.
-Specifically, we introduce a local search method based on Langevin MCMC that theoretically unifies prior gradient-based approaches, and we develop efficient global search strategies using tree-based algorithms such as BFS and DFS.
+    For global tree search, we elucidate the design space of prior BFS-style methods and provide an improved BFS design.
+We further present the first adaptive DFS algorithm for diffusion inference scaling, offering superior efficiency and adaptivity.
   • ii)
-    We implement inference-time scaling across a range of tasks—including planning, offline reinforcement learning, and image generation—and introduce task-specific verifiers.
-Our results demonstrate that inference-time scaling with search against verifiers can significantly improve the performance on many downstream tasks.
-Throughout the experiments, we show that our proposed approaches consistently surpass baselines in terms of compute efficiency.
+    We introduce a theoretically grounded local search method using annealed Langevin MCMC, demonstrating superior performance in challenging decision-making domains.
+  • iii)
+    We propose a unified framework for inference scaling that integrates both global and local search within the classical AI search paradigm.
+By jointly scaling local and global search for the first time, we advance the Pareto frontier of inference-time scaling for diffusion models across diverse domains.
 Figure 1: Illustration of our search framework.
 Bottom left: direct sampling results in samples with low verifier scores.
-Bottom middle: Local search navigates samples to nearby high-score regions, but suffers from local maximums and OOD samples.
-Bottom right: introducing global search explores diverse modes globally, allocating more compute to high-quality samples, and sampling from global optimal modes.
+Bottom middle: global search identifies high score modes within the base distribution.
+Bottom right: local search further optimizes the samples for higher quality, driven by the gradient signal.
 2 Related Works
-Here we provide a brief overview of inference-time scaling and diffusion models. For a more in-depth literature review see Appendix. B.
-Scaling inference-time compute has its roots in cognitive science, often associated with “System 2” thinking kahneman2011thinking ; anthony_thinking_2017 .
-Yao et al.
-yao2023tree conceptualize problem-solving in large language models (LLMs) as a combinatorial tree search problem newell1959report ; newell1972human , proposing the Tree-of-Thoughts framework, which leverages breadth-first and depth-first search for more efficient reasoning.
-Inference-time scaling has recently shown impressive performance with LLMs in reasoning and problem-solving tasks.
-For instance, jaech2024openai ; van2016deep train models to generate long chain-of-thought (CoT) wei2022chain , while others such as snell2024scaling ; wu2024inference incorporate process reward models lightman2023let to guide the generation process more effectively.
-These works focus on efficient exploration in discrete search spaces, often employing global search strategies.
-In contrast, diffusion models operate in a continuous, multimodal generative space.
-Furthermore, the autoregressive nature of LLMs allows for efficient evaluation of intermediate generation steps—something not directly feasible in diffusion models dhariwal2021diffusion ; lu2023contrastive .
-Inference-time scaling for diffusion models remains relatively underexplored.
-Recent works such as ma2025inference ; nakamotosteering introduce BFS-style global search algorithms, resampling candidate samples (or “particles”) at each noise level based on verifier scores (potentials).
-However, these methods do not effectively perform local search.
-Alternatively, Du et al.
-du_learning_2024 propose learning a series of annealed energy landscapes and applying Langevin MCMC for iterative refinement.
-In this work we propose a general principled framework, encompassing prior methods as instances of local search and global search.
+Here, we provide a brief overview of inference-time scaling with diffusion models. For a more comprehensive literature review and discussion of concurrent works, see Appendix B.
+Particle-based scaling methods
+Recent works (Kim et al., 2025; Singhal et al., 2025a) propose SMC-style particle filtering methods, scaling inference compute by increasing the number of particles and improving efficiency via resampling.
+Similarly, tree-search-based methods (Li et al., 2024; Guo et al., 2025; Ma et al., 2025) evaluate intermediate nodes and expand promising candidates.
+Both approaches can be seen as special cases of our BFS tree search framework, which denoise a set of particles in parallel, applying branching and filtering based on intermediate reward signals.
+Additional to linear BFS-style algorithms, we also propose DFS-style algorithms with non-linear adaptive backtracking, demonstrating superior performance over prior methods.
+Gradient-based guidance methods
+To sample from a conditional distribution, classifier guidance (Dhariwal & Nichol, 2021) utilizes the gradient from a trained noise-dependent classifier to compute the conditional score function (Song et al., 2020b).
+However, such noise-dependent classifier requires additional training and data collection.
+To use classifiers defined on clean samples, training-free guidance methods (Ye et al., 2024; Chung & Ye, 2022; Song et al., 2023; Yu et al., 2023; He et al., 2023) approximates noisy conditional probability using the denoised output ${\bm{x}}_{0|t}$.
+Such methods are inheritability biased due to their first order approximation.
+Different from prior methods that rely on the conditional diffusion process, we sample from the compositional distribution via annealed Langevin MCMC, which provides asymptotically exact sampling without any training.
 3 Backgrounds
@@
 Suppose we have $D$-dimensional random variable ${\bm{x}}_{0}\in\mathbb{R}^{D}$ with distribution $p_{0}({\bm{x}}_{0})$.
-Diffusion models ho2020denoising ; song2020denoising and the more general flow models lipman2022flow ; albergo2022building are generative models that turn noise into data via a stochastic process $\left\{{\bm{x}}_{t}\right\}_{t=0}^{T}$.
+Diffusion models (Ho et al., 2020; Song et al., 2020a) and the more general flow models (Lipman et al., 2022; Albergo & Vanden-Eijnden, 2022) are generative models that turn noise into data via a stochastic process $\left\{{\bm{x}}_{t}\right\}_{t=0}^{T}$.
 The forward “noising” process with $t>s$ can be defined as:
@@
 where $\alpha_{t}$, $\sigma_{t}$ are referred as the noise schedule with $\alpha_{0}=\sigma_{T}=1$, $\alpha_{T}=\sigma_{0}=0$.
-We can thus write the random variables ${\bm{x}}_{t}$ as an interpolation between data and noise ma2024sit :
+We can thus write the random variables ${\bm{x}}_{t}$ as an interpolation between data and noise (Ma et al., 2024):
 ${\bm{x}}_{t}=\alpha_{t}{\bm{x}}_{0}+\sigma_{t}{\bm{\epsilon}}\,$,
-and denote $q_{t}({\bm{x}}_{t})$ as the marginal distribution of ${\bm{x}}_{t}$. To model the reverse “denoising” process, we train the model using the denoising objective ho2020denoising :
+and denote $q_{t}({\bm{x}}_{t})$ as the marginal distribution of ${\bm{x}}_{t}$. To model the reverse “denoising” process, we train the model using the denoising objective (Ho et al., 2020):
 $L(\theta)=\mathbb{E}_{t,{\bm{x}}_{0},{\bm{\epsilon}}}\left[{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)-{\bm{\epsilon}}\right]\,$,
-which is equivalent of learning the score function of $q_{t}({\bm{x}}_{t})$ vincent2011score_matching , as the ground truth of $\epsilon_{\theta}({\bm{x}}_{t},t)$ is $-\sigma_{t}\nabla_{{\bm{x}}_{t}}\log q_{t}({\bm{x}}_{t})$.
+which is equivalent of learning the score function of $q_{t}({\bm{x}}_{t})$ (Vincent, 2011), as the ground truth of $\epsilon_{\theta}({\bm{x}}_{t},t)$ is $\sigma_{t}\nabla_{{\bm{x}}_{t}}\log q_{t}({\bm{x}}_{t})$.
 To generate samples, we transform noise into data via the reverse transition kernel $p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$.
-In practice, we either sample ${\bm{x}}_{t-1}$ using deterministic samplers like DDIM song2020denoising :
+In practice, we either sample ${\bm{x}}_{t-1}$ using deterministic samplers like DDIM (Song et al., 2020a):
 ${\bm{x}}_{t-1}=\frac{\alpha_{t-1}}{\alpha_{t}}({\bm{x}}_{t}-\sigma_{t}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t))+\sigma_{t-1}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)$
-or stochastic samplers like DDPM ho2020denoising ; nichol2021improved :
+or stochastic samplers like DDPM (Ho et al., 2020; Nichol & Dhariwal, 2021):
 $p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})=\mathcal{N}\left({\bm{x}}_{t-1};\bm{\mu}_{\theta}({\bm{x}}_{t},t),\bm{\Sigma}_{\theta}({\bm{x}}_{t},t))\right)\,.$
@@
 Given a base diffusion model with data distribution $p_{0}({\bm{x}}_{0})$, one may wish to sample ${\bm{x}}_{0}$ with some constraints or conditions $f({\bm{x}}_{0})$.
-Exact diffusion sampling from the composed distribution $\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})f({\bm{x}}_{0})$ would require training a time-dependent $f$ on data generated by $p_{0}$ dhariwal2021diffusion ; lu2023contrastive , which may not be applicable in practice.
+Exact diffusion sampling from the composed distribution $\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})f({\bm{x}}_{0})$ would require training a time-dependent $f$ on data generated by $p_{0}$ (Dhariwal & Nichol, 2021; Lu et al., 2023), which may not be applicable in practice.
 Thus, we adopt optimization based methods to approximate the target distribution.
 Compositional generation via annealed Langevin MCMC
-When sampling from a compositional distribution composed of multiple probability distributions, $\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})\hat{p}_{0}({\bm{x}}_{0})$, du2023reduce proposes annealed Langevin MCMC sampling.
-In this approach, a sequence of annealed distributions $\tilde{q}_{t}({\bm{x}}_{t})\propto q_{t}({\bm{x}}_{t})\hat{q}_{t}({\bm{x}}_{0})$ is constructed, and samples are drawn using Langevin dynamics welling2011bayesian :
+When sampling from a compositional distribution composed of multiple probability distributions, $\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})\hat{p}_{0}({\bm{x}}_{0})$, Du et al.
+(2023) proposes annealed Langevin MCMC sampling.
+In this approach, a sequence of annealed distributions $\tilde{q}_{t}({\bm{x}}_{t})\propto q_{t}({\bm{x}}_{t})\hat{q}_{t}({\bm{x}}_{0})$ is constructed, and samples are drawn using Langevin dynamics (Welling & Teh, 2011):
 ${\bm{x}}_{t}^{i+1}={\bm{x}}_{t}^{i}+\eta\nabla_{{\bm{x}}}\log\tilde{q}_{t}({\bm{x}}_{t}^{i})+\sqrt{2\eta}{\bm{\epsilon}}^{i}\,,~~~{\bm{\epsilon}}^{i}\sim\mathcal{N}(\bm{0},\bm{I})\,.$ (2)
@@
 This method approximates the intractable posterior with the posterior mean: $\mathbb{E}_{{\bm{x}}_{0}|{\bm{x}}_{t}}[f({\bm{x}}_{0})]\approx f(\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}])$.
-To enhance the guidance strength, yu2023freedom propose to apply a recurrence strategy, which first samples ${\bm{x}}_{t-1}$ via $p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$, add the guidance gradient, then add noise back to ${\bm{x}}_{t}$ through the forward process $q_{t}({\bm{x}}_{t}|{\bm{x}}_{t-1})$:
+To enhance the guidance strength, Yu et al.
+(2023) propose to apply a recurrence strategy, which first samples ${\bm{x}}_{t-1}$ via $p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$, add the guidance gradient, then add noise back to ${\bm{x}}_{t}$ through the forward process $q_{t}({\bm{x}}_{t}|{\bm{x}}_{t-1})$:
 ${\bm{x}}_{t-1}^{i}\sim p_{\theta}(\cdot|{\bm{x}}_{t}^{i})\,,~~\tilde{{\bm{x}}}_{t-1}^{i}={\bm{x}}_{t-1}^{i}+\frac{\alpha_{t-1}}{\alpha_{t}}\bm{\Delta_{t}}\,,~~{\bm{x}}_{t}^{i+1}\sim q_{t}(\cdot|\tilde{{\bm{x}}}_{t-1}^{i})\,,~~~i=1,2,\cdots,N_{\text{recur}}\,$, (4)
@@
 Given a pretrained diffusion model $\epsilon_{\theta}({\bm{x}}_{t},t)$ with a base distribution $p_{0}({\bm{x}}_{0})$, at test-time, we often wish to optimize the generation process to satisfy task-specific objectives.
-For instance, in reinforcement learning, this might involve generating high-value actions; in image synthesis, we may seek images that satisfy certain constraints such as color or shape; and in trajectory generation, we may want the outputs to conform to physical laws.
+For example, RL may require generating high-value actions, image synthesis may seek constraint-satisfying images, and trajectory generation may demand physically valid outputs.
 In this paper, we are interested in how to scale test-time inference to follow such objectives.
@@
 This leads to the objective of sampling from a compositional distribution that combines the original model distribution with the verifier:
-$\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})f({\bm{x}}_{0})^{\lambda}\,,$
+$\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})f({\bm{x}}_{0})^{\lambda}\,$, (5)
 where $\lambda$ controls the weight of verifier scores.
-Since exact sampling from the distribution is often impractical, we aim to search the manifold for the target samples at inference time, both locally and globally.
-First, we search the vicinity of the sample using hill-climbing style local search methods, guided by the verifier gradient.
-However, local search may stuck in a local maximum, where we then propose global graph search algorithms to explore the diverse modes in the complex generative landscape of diffusion models.
-4.1 Local Search via Langevin MCMC with Verifier Gradient
-To sample from the composition distribution $\tilde{p}_{0}$, we conduct local-search with hill-climbing methods, aiming to find the local maximum with high $\tilde{p}_{0}$.
-Specifically, we view the sampling problem as compositional optimization in measure space wibisono2018sampling , and follow the gradient flow of KL-divergence, performing Langevin MCMC steps (details see Appendix.
+Since exact sampling from the distribution is often impractical, we aim to search the manifold for the target samples at inference time, both globally and locally.
+First, we explore the diverse modes in the complex generative landscape of diffusion models through global graph search algorithms.
+However, global search alone can not generate samples beyond the pretrained model.
+We then propose to search the vicinity of the sample using hill-climbing style local search methods, guided by the verifier gradient.
+By combining local search with global search, we can avoid being stuck in local optima, advancing the performance of the base model by sampling beyond the pretrained modes.
+4.1 Global Search for Mode Identification
+To efficiently explore the modes of a diffusion model, we represent the Markov chain of its denoising process as a fixed-depth tree, where the transition kernel $\tilde{p}_{\theta}({\bm{x}}{t-1}|{\bm{x}}_{t})$ may correspond to either deterministic or stochastic samplers.
+This abstraction enables the use of classical tree-search heuristics to design compute-efficient exploration methods.
+A simple baseline is best-of-$N$ (BoN) sampling: generate $N$ independent trajectories and select the one with the highest verifier score at the final step.
+While straightforward, this approach discards valuable information from intermediate stages.
+To improve efficiency, inspired by best-first search (Pearl, 1984), we evaluate intermediate particles using the denoised estimate ${\bm{x}}_{0|t}$.
+This allows dynamic allocation of computation based on particle quality, using techniques such as branching and backtracking.
+We next describe two representative efficient search strategies—breadth-first search (BFS) and depth-first search (DFS)—adapted for diffusion model sampling.
+By prioritizing the expansion of nodes with higher score estimates and backtracking from low-quality regions of the tree, we can more effectively navigate the generative space and identify high-quality modes.
+An illustration of the tree search methods is provided in Fig.
+2.
+4.1.1 Unified BFS-style linear search
+Figure 2: Illustration of global tree search algorithms.
+Inspired by breadth-first search (BFS), which expands nodes level by level, we denoise a set of particles in parallel at each noise level.
+We score each intermediate particle $\left\{{\bm{x}}_{t}^{k}\right\}_{k=1}^{N}$ using estimates of its verifier score $f({\bm{x}}_{0|t}^{k})$, and dynamically reallocate computation by sampling more children $n_{t}^{k}$ for high-scoring nodes.
+We provide a general design space for schedules, scoring functions, and resampling strategies that unifies previous tree-search-based and particle-based baselines such as SVDD (Li et al., 2024), DAS (Kim et al., 2025), and FK-steering (Singhal et al., 2025b).
+The pseudocode is shown in Alg.
+3, and more details can be found in Sec.
+D.1.
+Tempering
+To mitigate score estimation bias in early steps, Kim et al.
+(2025) increases weights on smaller time steps with $\tau_{T}<\tau_{T-1}\cdots<\tau_{0}$, re-weighting scores with $\tau_{t}f({\bm{x}}_{0|t}^{k})$.
+Li et al.
+(2024) samples only from the top-scoring particle, i.e., $\tau_{t}=\infty$.
+We consider three different tempering design choices: $\textbf{Constant}:\tau_{t}=\tau,~\textbf{Increase}:\tau_{t}=\left((1+\gamma)^{T-t}-1\right)\tau,~\textbf{Inf}:\tau_{t}=\infty$~.
+Scoring
+Following Kim et al.
+(2025); Singhal et al.
+(2025b), we propose to evaluate intermediate particles $\widehat{f}({\bm{x}}_{t}^{k})$ using functions of the estimated rewards $f({\bm{x}}_{0|s}^{k})$ along the sampling path.
+This approach accounts for reward variation during sampling by considering not only the current reward but also its trajectory.
+Specifically, we define the following scoring functions $\widehat{f}({\bm{x}}_{t}^{k})$ based on $f({\bm{x}}_{0|s}^{k})$ for $s\geq t$: $\textbf{Current}:\tau_{t}f({\bm{x}}_{0|t}^{k}),~\textbf{Difference}:\tau_{t}f({\bm{x}}_{0|t}^{k})-\tau_{t+1}f({\bm{x}}_{0|t+1}^{k}),~\textbf{Max}:\max_{s\geq t}\tau_{s}f({\bm{x}}_{0|s}^{k}).$
+Resampling
+Given $\widehat{f}({\bm{x}}_{t}^{k})$, we allocate particles as $\left(n_{t}^{1},\cdots,n_{t}^{N}\right)=\text{Resample}\left(N;w_{t}^{1},\cdots,w_{t}^{N}\right)$, where $n_{t}^{k}$ is the number of children for ${\bm{x}}_{t}^{k}$ and $w_{t}^{k}=\text{softmax}{\widehat{f}({\bm{x}}_{t}^{k})}$ is the softmax score of particle $k$.
+Multinomial resampling (Singhal et al., 2025b) samples $n_{t}^{k}$ independently from the multinomial distribution $w_{t}^{k}$, but suffers from large variance.
+Kim et al.
+(2025) adopts the Srinivasan sampling process (SSP) resampling (Alg.
+4) for reduced variance.
+We compare the baseline Multinomial resampling and the variance-reduced SSP (Kim et al., 2025); see Gerber et al.
+(2020) for detailed discussions of other resampling methods.
+Prior methods are special cases of BFS: SVDD (Li et al., 2024) = BFS (Inf, Current, Multinomial); DAS (Kim et al., 2025) = BFS (Increase, Difference, SSP); FK-steering (Singhal et al., 2025a) = BFS (Constant, Max, Multinomial).
+Ablations (Sec.
+5.1) show SSP resampling is key for performance, and our improved BFS (Increase, Max, SSP) consistently outperforms prior methods in efficiency.
+4.1.2 DFS-style non-linear search
+Depth-first search (DFS) explores one branch of the search tree as deeply as possible before backtracking.
+In our setting, this corresponds to iteratively denoising a single particle until its verifier score drops below a predefined threshold: $f({\bm{x}}_{0|t})\leq\delta_{t}$, where $\delta_{t}$ is a threshold representing the quality requirement of the users.
+Once the constraint is violated, the algorithm backtracks by reintroducing noise, moving to a higher noise level $t_{\text{next}}=t+\Delta_{T}$ using the forward diffusion process $q({\bm{x}}_{t_{\text{next}}}|{\bm{x}}_{t})$ in Eq.
+1.
+This allows the model to restart the denoising process from a different region of the manifold, encouraging exploration of diverse modes.
+Unlike the small noise injection and fixed schedule used in SoP (Ma et al., 2025) for local exploration, DFS performs global exploration with large backtracking depth $\Delta_{T}\geq\frac{T}{4}$, and adopts an score-adaptive exploration strategy .
+A key strength of DFS is its ability to allocate compute adaptively: difficult prompts and low-quality trajectories naturally trigger more backtracking and exploration, while easier instances are solved more directly.
+This dynamic behavior is driven purely by the verifier signal, without needing to know the difficulty in advance as in Snell et al.
+(2024).
+Also, the threshold acts as a control knob for users to balance output quality and computation resources, where higher threshold automatically scales compute for better output.
+As shown in Sec.
+5.2, this adaptive strategy leads to substantial gains in efficiency and performance over prior methods, and even our improved BFS implementation.
+The pseudocode is in Alg.
+5 and more details can be found in Sec.
+D.2.
+4.2 Scaling Local Search via Langevin MCMC with Verifier Gradient
+Global search can efficiently discover the high score modes from the base diffusion model, but can not generate higher quality samples that exceed the pretrained model.
+Thus, we aim to sample from the compositional distribution $\tilde{p}_{0}$ in Eq.
+5 for higher quality samples.
+To optimize the compositional objective, we conduct local-search with hill-climbing methods, aiming to find the local maximum with high $\tilde{p}_{0}$.
+Specifically, we view the sampling problem as compositional optimization in measure space (Wibisono, 2018), and follow the gradient flow of KL-divergence, performing Langevin MCMC steps (details see Appendix.
 C.1).
-Similar to annealed Langevin MCMC in du2023reduce , we could construct a series of annealed functions $\hat{f}_{t}({\bm{x}}_{t})$ with $\hat{f}_{0}({\bm{x}}_{0})=f({\bm{x}}_{0})$.
+Similar to annealed Langevin MCMC in Du et al.
+(2023), we could construct a series of annealed functions $\hat{f}_{t}({\bm{x}}_{t})$ with $\hat{f}_{0}({\bm{x}}_{0})=f({\bm{x}}_{0})$.
 Then we sample from the distributions $\tilde{q}_{t}({\bm{x}}_{t})\propto q_{t}({\bm{x}}_{t})\hat{f}({\bm{x}}_{t})$ through Langevin MCMC in Eq.
@@
 C.2).
-However, it remains unclear how to design such annealed verifier functions $\hat{f}_{t}({\bm{x}}_{t})$ efficiently, especially without training.
 Alternatively, training-free guidance in Eq.
 3 utilizes the gradient of $f({\bm{x}}_{0|t})$ to optimize ${\bm{x}}_{t}$, which can be computed directly using the diffusion model output.
-However, naive gradient updates have been observed to produce OOD and adversarial samples shen2024understanding .
-In ye2024tfg , recurrence (Eq.
+However, naive gradient updates have been observed to produce OOD and adversarial samples (Shen et al., 2024).
+In Ye et al.
+(2024), recurrence (Eq.
 4) was found to help avoid adversarial samples in challenging guidance tasks, though its theoretical underpinnings remain poorly understood.
@@
 For details see Appendix.
-C.3.
+C.3, and a rigorous convergence bound is in Theorem.
+1.
 Proposition 1.
 In the continuous limit where the number of diffusion denoising steps $T\rightarrow\infty$, training-free guidance with recurrence is equivalent to running Langevin MCMC on a series of annealed distributions $\left\{\tilde{q}_{t}({\bm{x}}_{t})\right\}_{t=0}^{T}$, with $\tilde{q}_{0}({\bm{x}}_{0})=\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})f({\bm{x}}_{0})^{\lambda}$.
-Thus, the recurrence step (without guidance) can be interpreted as Langevin MCMC applied to the original distribution of the diffusion model $q_{t}({\bm{x}}_{t})$, effectively guiding the samples back toward the data manifold (details see Appendix.
-C.4).
-The guidance term then serves as defining a practical annealing path $\hat{f}_{t}({\bm{x}}_{t})$, biasing the sampling path.
+Thus, the recurrence step (without guidance) can be interpreted as Langevin MCMC applied to the original distribution of the diffusion model $q_{t}({\bm{x}}_{t})$, and the guidance term $\bm{\Delta}_{t}$ in Eq.
+3 then serves as defining a practical annealing path $\hat{f}_{t}({\bm{x}}_{t})$ that bias the sampling path towards high reward areas beyond the modes of the base model.
 We are the first to propose this theoretical unification of the two lines of work, providing insights into efficient local search of diffusion models via gradients.
-To implement local search, we parameterize the reverse transition kernel $\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$ as a sequence of Langevin MCMC steps in Eq.
-2, followed by a denoising step with DDIM (Eq.
-10) or DDPM (Eq.11).
-Owing to the equivalence between Langevin MCMC and training-free guidance, we can efficiently tune hyper-parameters using the unified design space of ye2024tfg , with details in Appendix.
-C.5.
-We scale up the number of local search steps and observe significant improvements across multiple tasks (Fig.
-4(d)), highlighting the effectiveness of local search.
-4.2 Global Search for Mode Identification
-Figure 2: Illustration of global tree search algorithms.
-Left: Best-of-N sampling denoises multiple samples in parallel and selects the one with the highest verifier score at the final step.
-Middle Left: BFS-Resampling resamples particles at intermediate noise levels, allocating more children to nodes with higher scores.
-Middle Right: BFS-Pruning removes particles with low verifier scores during denoising, sampling at most one child per node.
-Right: DFS denoises a single particle until its verifier score violates the constraint, then backtracks by adding noise to a previous level and continuing the search.
-Although Langevin MCMC is asymptotically exact, it often struggles to transition between distant modes latuszynski2025mcmc due to its inherently local nature.
-This limitation is especially pronounced in the context of the multimodal distributions induced by diffusion models.
-To mitigate this, we propose identifying high-score modes via a global search strategy.
-We model the Markov chain of the denoising process as a fixed-depth tree, where the transition kernel is given by $\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$.
-As discussed earlier, when applying local search, we parameterize the transition kernel with multiple Langevin MCMC steps, and otherwise we directly adopt conventional diffusion samplers song2020denoising ; ho2020denoising ; lu2022dpm .
-Hence, our perspective unifies both stochastic and deterministic sampling with guided sampling.
-This abstraction decouples the transition kernel from the sampling process, enabling the application of classic tree search heuristics to develop compute-efficient exploration methods.
-BFS-style linear search
-Inspired by breadth-first search (BFS), which expands nodes level-by-level, we denoise a set of particles in parallel at each noise level.
-The simplest approach is best-of-$N$ sampling: generate $N$ candidate trajectories and select the one with the highest verifier score at the final step.
-While simple, this strategy ignores information from intermediate stages.
-To improve efficiency, similar to best-first search pearl1984heuristics , we evaluate each intermediate particle $\left\{{\bm{x}}_{t}^{k}\right\}_{k=1}^{K_{t}}$ using its verifier score $f({\bm{x}}_{0|t}^{k})$, where $K_{t}$ is the number of particles at timestep $t$.
-Based on these scores, we dynamically reallocate computation by sampling more children for high-scoring nodes via the reverse transition kernel $\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$.
-This forms the basis of BFS-Resampling, where each node samples children in proportion to its normalized verifier score
-$n_{t}^{k}=\texttt{Round}\left(K_{t}\frac{f({\bm{x}}_{0|t}^{k})^{\tau_{t}}}{\sum_{i=1}^{K_{t}}f({\bm{x}}_{0|t}^{i})^{\tau_{t}}}\right)\,.$
-We then sample $n_{t}^{k}$ children from ${\bm{x}}_{t}^{k}$ using $\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$ and proceed to the next denoising step.
-When using deterministic ODE solvers lu2022dpm as transition kernels, it is ineffective to sample multiple children from the same parent.
-To address this, we propose BFS-Pruning, a variant that retains only top-scoring nodes and allows at most one child per parent $n_{t}^{k}=\min\left(1,n_{t}^{k}\right)$.
-This leads to a gradual narrowing of the search tree as $K_{t}$ decreases over time, enabling us to explore more diverse modes with more initial particles under the same compute.
-As shown in Fig. 6, both resampling and pruning significantly improve efficiency over standard best-of-$N$ sampling. The detailed pseudo code is in Alg. 3.
-DFS-style non-linear search
-Depth-first search (DFS) explores one branch of the search tree as deeply as possible before backtracking.
-In our setting, this corresponds to iteratively denoising a single particle until its verifier score drops below a predefined threshold: $f({\bm{x}}_{0|t})\leq\delta_{t}$, where $\delta_{t}$ is a scheduled threshold for timestep $t$.
-Once the constraint is violated, the algorithm backtracks by reintroducing noise, moving to a higher noise level $t_{\text{next}}=t+\Delta_{T}$ using the forward diffusion process $q({\bm{x}}_{t_{\text{next}}}|{\bm{x}}_{t})$ in Eq.
-1.
-This allows the model to restart the denoising process from a different region of the manifold, encouraging exploration of diverse modes.
-By placing a cap on the number of backtracking steps, we can explore the space globally while keeping computation bounded.
-A key strength of DFS is its ability to allocate compute adaptively: difficult prompts and low-quality trajectories naturally trigger more backtracking and exploration, while easier instances are solved more directly.
-This dynamic behavior is driven purely by the verifier signal, without needing to know the difficulty in advance.
-As shown in Fig.
-5, this strategy leads to substantial gains in efficiency and performance.
-The detailed pseudo code is in Alg.
-4.
+We implement local search by parameterizing the reverse transition kernel $\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$ as a sequence of Langevin MCMC steps (Eq.
+2), followed by a denoising step using DDIM (Eq.
+11) or DDPM (Eq.
+12); see Appendix C.5 for details.
+Unlike classifier-guidance or naive training-free guidance, which apply only gradient guidance in the denoising step, our approach incorporates explicit Langevin MCMC steps.
+In Sec.
+5.3, we scale the number of local search steps for the first time and observe substantial improvements over pretrained models across multiple tasks.
 5 Experiments
 In this section, we apply inference-time scaling with our search strategy across a range of domains.
-In Sec.5.1, we show that diffusion models can generate long-horizon plans in complex maze environments with inference-time scaling, given a world model as verifier.
-In Sec.5.2, we leverage a pretrained Q-function as a verifier to guide diffusion policies, achieving strong performance on D4RL fu2020d4rl locomotion benchmarks.
-Finally, in Sec.
-5.3, we approach the image generation task, and demonstrate that our global search algorithms consistently outperform best-of-N sampling in terms of efficiency, and we reveal interesting properties of BFS and DFS.
-5.1 Long Horizon Planning
-Figure 4: (Illustration and results for maze planning) Fig.
-4(a) shows the maze layout of Giant (left) and Ultra (right), with start: $\bullet$ and goal: $\star$.
-Fig.
-4(b) illustrates the effect of local search, where left shows the directly sampled trajectory that violates the maze block on the bottom left, and right shows the successful sample after local search.
-Fig.
-4(c) shows that inference time search generates diverse solutions.
-Fig.
-4(d) shows the performance with scaling number of local search steps with Langevin MCMC.
-Fig.
-4(e) and Fig.
-4(f) shows the scaling behaviors of the global search methods for PointMaze Giant and PointMaze Ultra respectively.
-Details see Appendix.
-E.1.
-Diffusion models have been widely adopted in planning for trajectory synthesis ubukata2024diffusion .
-We evaluate long-horizon planning in the PointMaze environment from OGBench park2024ogbench , training the diffusion model following the approach in Diffuser janner2022planning , where the full trajectory is denoised:
-$\bm{\tau}=\left[\begin{array}[]{cccc}\bm{s}_{1}&\bm{s}_{2}&\cdots&\bm{s}_{H}\\ \bm{a}_{1}&\bm{a}_{2}&\cdots&\bm{a}_{H}\end{array}\right]\,.$
-To highlight the effectiveness of inference-time scaling, we evaluate on the largest PointMaze Giant environment and introduce an even more challenging variant, PointMaze Ultra, following the protocol from marcucci2023motion .
-The maze layouts are shown in Fig.
-4(a).
-We define the verifier loss for each colliding point of the trajectory as the squared distance to its nearest wall, thus the overall verifier score measures the degree the trajectory violates the environment (details see Appendix.
-E.1).
-This verifier functions as a physics-grounded instance of a world model.
-Importantly, naively maximizing the verifier score does not guarantee a successful plan, and planning remains difficult even with full access to the maze layout luo2024potential ; marcucci2023motion .
-In the most challenging scenario—navigating from the bottom-left to the top-right corner—naive sampling strategies rarely succeed, necessitating multiple local search steps (Fig.
-4(d)).
-As illustrated in Fig.
-4(b), local search guided by the verifier enables the model to find successful plans that navigate around obstacles.
+In Sec.
+5.1, we present a strengthened BFS baseline that outperforms previous particle-based methods.
+In Sec.
+5.2, we demonstrate the adaptivity and efficiency of our DFS method.
+In Sec.
+5.3, we scale up local search in challenging decision-making domains, highlighting the importance of jointly scaling local and global search.
+In Sec.
+5.4, we propose a double-verifier strategy to mitigate the reward hacking problem in inference-scaling.
+5.1 Elucidating the Design Space of BFS
+In this section, we explore the design choices of BFS, and present an improved BFS (Increase, Max, SSP) over prior particle-based or tree-search-based methods (Singhal et al., 2025b; Kim et al., 2025; Guo et al., 2025; Li et al., 2024).
+To ensure a fair comparison, we directly use the official implementation of FK-steering (Singhal et al., 2025a) with the ImageReward (Xu et al., 2023a) verifier and the SD v1.5 model.
+For details, see Appendix E.1.
+Table 1: Ablation of BFS design choices.
+For each section, we vary one design choice while keeping the others fixed.
+Left: Resampling methods (with Constant tempering, Max scoring).
+Center: Scoring functions (with SSP resampling, Constant tempering).
+Right: Tempering schedules (with SSP resampling, Max scoring).
+Resampling Scoring Tempering
+N BoN Multinomial SSP Current Difference Max Increase Inf Constant
+4 $0.702\pm 0.057$ $0.743\pm 0.037$ $\mathbf{0.834\pm 0.041}$ $0.812\pm 0.037$ $0.823\pm 0.036$ $\mathbf{0.834\pm 0.041}$ $\mathbf{0.882\pm 0.029}$ $0.667\pm 0.076$ $0.834\pm 0.041$
+8 $0.896\pm 0.031$ $0.926\pm 0.042$ $\mathbf{1.032\pm 0.035}$ $0.996\pm 0.029$ $1.013\pm 0.032$ $\mathbf{1.032\pm 0.035}$ $\mathbf{1.087\pm 0.031}$ $0.775\pm 0.087$ $1.032\pm 0.035$
+As mentioned in Sec.4.1.1, we experiment with different resampling, scoring, and tempering design choices.
+We begin with BFS (Constant, Max, Multinomial) and evaluate various resampling strategies.
+As shown in Table 1 (left), SSP significantly outperforms naive multinomial resampling, and we adopt it in our design.
+We then assess scoring functions and tempering schedules under SSP resampling, as reported in Table 1 (center) and 1 (right).
+Here, we observe modest improvements with Max scoring and Increase tempering, leading to our final design: BFS (Increase, Max, SSP).
+Model N BoN FKSinghal et al. (2025a) DASKim et al. (2025) TreeG Guo et al. (2025) SVDDLi et al. (2024) BFS (ours)
+SD v1.5 4 $0.702\pm 0.057$ $0.743\pm 0.037$ $0.878\pm 0.028$ $0.860\pm 0.033$ $0.667\pm 0.076$ $\mathbf{0.882\pm 0.029}$
+SD v1.5 8 $0.896\pm 0.031$ $0.926\pm 0.042$ $1.052\pm 0.033$ $1.023\pm 0.018$ $0.775\pm 0.087$ $\mathbf{1.087\pm 0.031}$
+[…]
+SDXL 8 $1.198\pm 0.021$ $1.251\pm 0.011$ $1.265\pm 0.019$ $1.261\pm 0.021$ $1.225\pm 0.027$ $\mathbf{1.291\pm 0.018}$
+Table 2: Comparison of our BFS with prior methods
+To compare our improved BFS (Increase, Max, SSP) with prior baselines, we additionally experiment with the SDXL model (Podell et al., 2023), which differs from the model used in our ablations.
+As shown in Table 2, our improved BFS consistently outperforms previous methods across compute budgets and models.
+In the following experiments, we use the improved BFS as our baseline.
+5.2 Adaptive and Efficient Inference-Scaling with DFS
+Figure 3: CompBench text-to-image results with DFS.
+Left: the Pareto curve of DFS, with DFS-$\delta$ denotes DFS with threshold $\delta_{t}=\delta$.
+Right: average compute allocation by DFS for prompts of increasing difficulty.
+In this section, we evaluate the adaptivity and efficiency of DFS on the CompBench dataset (Huang et al., 2023), using the SSD-1B model (Gupta et al., 2024) and a VLM (Li et al., 2022) as our verifier.
+The detailed setup is provided in Appendix E.2.
+Through these experiments, we address the following questions:
+  • Can DFS outperform Best-of-N and prior particle-based methods?
 As shown in Fig.
-4(e) and Fig.
-4(f), incorporating global search alongside local refinement leads to over 90% success rates in both maze environments, with both BFS and DFS significantly improving search efficiency compared to naive best-of-$N$ sampling.
-Moreover, as shown in Fig.
-4(c), global search enables the generation of diverse, high-quality solutions.
-5.2 Offline Reinforcement Learning
-Table 1: Performance on D4RL locomotion tasks, score normalized as suggested in fu2020d4rl .
-Following lu2023contrastive we report the mean and standard deviation over 5 independent trials, highlighting the numbers within 5 percent of the maximum in every individual task.
-To faithfully report our performance, we conduct the hyper-parameter search and evaluation with different seeds.
-For more details see Appendix.
-E.2.
-Dataset Environment CQL BCQ IQL SfBC DD Diffuser D-QL QGPO TTS(ours)
-Medium-Expert HalfCheetah 62.4 64.7 86.7 92.6 90.6 79.8 96.1 93.5 $\mathbf{93.9\pm 0.3}$
-Medium-Expert Hopper 98.7 100.9 91.5 108.6 111.8 107.2 110.7 108.0 $104.4\pm 3.1$
-Medium-Expert Walker2d 111.0 57.5 109.6 109.8 108.8 108.4 109.7 110.7 $\mathbf{111.4\pm 0.1}$
-Medium HalfCheetah 44.4 40.7 47.4 45.9 49.1 44.2 50.6 54.1 $\mathbf{54.8\pm 0.1}$
-Medium Hopper 58.0 54.5 66.3 57.1 79.3 58.5 82.4 98.0 $\mathbf{99.5\pm 1.7}$
-Medium Walker2d 79.2 53.1 78.3 77.9 82.5 79.7 85.1 86.0 $\mathbf{86.5\pm 0.2}$
-Medium-Replay HalfCheetah 46.2 38.2 44.2 37.1 39.3 42.2 47.5 47.6 $\mathbf{47.8\pm 0.4}$
-Medium-Replay Hopper 48.6 33.1 94.7 86.2 100.0 96.8 100.7 96.9 $\mathbf{97.4\pm 4.0}$
-Medium-Replay Walker2d 26.7 15.0 73.9 65.1 75.0 61.2 94.3 84.4 $79.3\pm 9.7$
-Average (Locomotion) 63.9 51.9 76.9 75.6 81.8 75.3 86.3 86.6 86.1
-Recently, diffusion models emerged as a powerful action prior in robotics due to its ability to model complex and multimodal distributions chi2023diffusion ; liu2024rdt .
+E.2 (left), DFS consistently outperforms Best-of-N and also our improved BFS design across different compute budgets, with up to $2\times$ less computational cost.
+  • Can DFS adjust compute allocation with different thresholds?
+We evaluate DFS across a wide range of practical threshold values ($0.5$, $0.7$, and $0.9$) and find that lower thresholds automatically allocate less compute, while higher thresholds scale up compute for better quality.
+DFS consistently outperforms baseline methods across all threshold choices, demonstrating the robustness of our method for different hyperparameters.
+  • Can DFS dynamically adjust compute allocation for different instances?
+We measure the computational cost of DFS on prompts of varying difficulty under fixed thresholds, where the difficulty of a prompt is defined as the average score over four random samples.
+As shown in Fig.
+3 (right), harder prompts with lower scores naturally consume more compute, without requiring prior knowledge of difficulty as in Snell et al.
+(2024).
+Unlike linear-search methods that use a fixed exploration schedule, DFS offers higher efficiency and adaptivity, which may be of independent interest to the broader community.
+5.3 Joint Scaling Local and Global Search
+Although global search methods such as BFS and DFS can efficiently explore the generative space of the diffusion model, they are restricted to the modes of the base distribution and therefore cannot exceed the capabilities of the base model.
+To optimize the compositional objective in Eq.
+5 and sample from high-reward regions beyond the base model, we propose scaling up local search steps via annealed Langevin MCMC, introducing a new scaling dimension for diffusion models.
+We validate the effectiveness of scaling local search in challenging decision-making domains, such as long-horizon planning and offline RL.
+Baselines
+To evaluate the effectiveness of scaling local search steps, we compare our method with particle-based DAS (Kim et al., 2025), which also leverages verifier gradients but only incorporates gradient guidance without recurrent local search.
+We also compare against the state-of-the-art training-free guidance method TFG (Ye et al., 2024), which performs multiple recurrence steps but lacks any global search.
+For fairness, compute is measured as the total NFEs of both local and global search.
+As shown in the following experiments, scaling local and global search separately yields suboptimal performance, whereas our joint scaling strategy establishes a new Pareto frontier.
+5.3.1 Long Horizon Planning
+Figure 4: Pareto curves of local search.
+Left: Pareto curves of best-of-N with different local search steps, where BoN-$i$ denotes $i$ local search steps.
+Right: Pareto curves of BFS and DFS with 6 local search steps.
+Figure 5: Illustration of Maze layout and task, with failed trajectory (left) and successfual sample (right).
+Diffusion models have been widely adopted in planning for trajectory synthesis (Ubukata et al., 2024).
+We evaluate long-horizon planning in a challenging PointMaze environment, using the base model trained following Diffuser (Janner et al., 2022), with the verifier defined as the total number of collisions between the trajectory and maze walls (see Appendix E.3 for details).
+(Luo et al., 2024; Marcucci et al., 2023).
+As shown in Fig.
+5 (left), naive sampling without local search often resulted in failed trajectories that violates the maze layout in some of the segments.
+When scaling up local search steps as in Fig.
+5 (right), we are able to generate successful plans free of violation.
+To evaluate the compute efficiency of recurrent steps, we observe in Fig.
+4 (left) that scaling up local search with BoN improves the overall Pareto frontier and significantly outperforms baseline methods.
+Scaling local search only in TFG (Ye et al., 2024) is efficient with a low compute budget but fails to scale with increased compute, as local search alone can become trapped in local optima.
+DAS (Kim et al., 2025) is more efficient than the corresponding BoN-0 baseline without local search, but underperforms best-of-N when more local search steps are used.
+In Fig.
+4 (right), we show that local search can be combined with global search techniques such as BFS and DFS to further improve scaling efficiency, demonstrating the flexibility of our framework.
+5.3.2 Offline Reinforcement Learning
+Table 3: Performance on D4RL locomotion tasks, highlighting results within 5% of the maximum.
+Dataset Environment IQL SfBC $D$ Diffuser D-QL QGPO TFG DAS TTS(ours)
+Medium-Expert HalfCheetah 86.7 92.6 90.6 79.8 96.1 93.5 $90.2\pm 0.2$ $\mathbf{93.3\pm 0.3}$ $\mathbf{93.9\pm 0.3}$
+Medium-Expert Hopper 91.5 108.6 111.8 107.2 110.7 108.0 $100.2\pm 3.5$ $105.4\pm 5.1$ $104.4\pm 3.1$
+Medium-Expert Walker2d 109.6 109.8 108.8 108.4 109.7 110.7 $\mathbf{108.1\pm 0.1}$ $\mathbf{111.4\pm 0.1}$ $\mathbf{111.4\pm 0.1}$
+Medium HalfCheetah 47.4 45.9 49.1 44.2 50.6 54.1 $\mathbf{53.1\pm 0.1}$ $\mathbf{53.4\pm 0.1}$ $\mathbf{54.8\pm 0.1}$
+Medium Hopper 66.3 57.1 79.3 58.5 82.4 98.0 $\mathbf{96.2\pm 0.5}$ $71.3\pm 2.7$ $\mathbf{99.5\pm 1.7}$
+Medium Walker2d 78.3 77.9 82.5 79.7 85.1 86.0 $\mathbf{83.2\pm 1.4}$ $\mathbf{83.9\pm 0.9}$ $\mathbf{86.5\pm 0.2}$
+Medium-Replay HalfCheetah 44.2 37.1 39.3 42.2 47.5 47.6 $\mathbf{45.0\pm 0.3}$ $42.2\pm 0.1$ $\mathbf{47.8\pm 0.4}$
+Medium-Replay Hopper 94.7 86.2 100.0 96.8 100.7 96.9 $93.1\pm 0.1$ $\mathbf{96.7\pm 3.0}$ $\mathbf{97.4\pm 4.0}$
+Medium-Replay Walker2d 73.9 65.1 75.0 61.2 94.3 84.4 $69.8\pm 4.0$ $63.8\pm 2.0$ $79.3\pm 9.7$
+Average (Locomotion) 76.9 75.6 81.8 75.3 86.3 86.6 82.1 80.2 86.1
+Recently, diffusion models have emerged as a powerful action prior in robotics due to their ability to model complex and multimodal distributions (Chi et al., 2023).
 However, these diffusion policies are typically trained on offline datasets and struggle to adapt to reinforcement learning or test-time requirements.
-Following prior work peters2010relative , we formulate the offline RL problem as sampling from a Q-regularized distribution:
-$\pi^{*}(\bm{a}|\bm{s})\propto\mu(\bm{a}|\bm{s})e^{\beta Q_{\psi}(\bm{s},\bm{a})}\,,$
-where $Q_{\psi}$ is a learned Q-function representing preferences over actions, and $\mu$ is the behavior policy, which we model using a diffusion prior.
-In principle, exact sampling from $\pi^{*}$ using diffusion models requires training a noise-dependent energy function on noisy actions $\bm{a}_{t}$ janner2022planning ; lu2023contrastive .
-However, this approach is often impractical in real-world settings.
-First, it requires evaluating the pretrained Q-function on samples from the diffusion model, which is costly due to the difficulty of data collection in robotics.
-Second, many Q-functions rely on foundation models as backbones brohan2023can ; nakamotosteering , and training them on noisy samples risks degrading their pretrained knowledge.
-In contrast, test-time search (TTS) offers a promising alternative.
-By decoupling policy and value, TTS enables flexible composition with foundation models without retraining.
-We evaluate TTS using both local and global search strategies on the D4RL locomotion benchmarks fu2020d4rl .
-Following the setup in lu2023contrastive , we adopt their pretrained diffusion policy and ground-truth Q-function.
-Results are presented in Table 1, where our approach achieves performance competitive with the best training-based methods, including QGPO lu2023contrastive and Diffusion-QL wang2022diffusion .
-In comparison, our approach is training-free, requiring only an off the shelf Q function and diffusion model.
-5.3 Image Generation
-In this section we evaluate inference-time scaling in image generation tasks, especially the efficiency of global search.
-5.3.1 Compositional Text-to-Image Generation
-Figure 5: CompBench huang2023t2icompbench text-to-image results.
-Fig.
-5(a) and Fig.
-5(b) show DFS performance on attribute binding and object relationship tasks, achieving up to 2× lower computational cost than Best-of-N.
-Fig.
-5(c) presents the average compute allocated by DFS for prompts of increasing difficulty, evaluated on the color dataset.
-We evaluate inference-time scaling for text-to-image generation on the CompBench dataset huang2023t2icompbench using the SSD-1B model gupta2024progressive .
-For the search verifiers, we use task-specific oracle verifiers: a BLIP-VQA model li2022blip for the attribute binding task and a UniDet model zhou2022simple for the object relationship task.
-Since these verifiers are non-differentiable, we apply global search exclusively.
-Detailed setup is in Appendix.
-E.3.1.
-Improved efficiency and adaptivity with DFS
-We evaluate Best-of-N and DFS on the five datasets in attribute binding and object relationship tasks, with each dataset consists of 300 prompts.
-The results are shown in Fig.
-5(a) and Fig.
-5(b), where DFS consistently outperforms Best-of-N while using up to 2× less compute.
-To evaluate DFS’s ability to adaptively allocate computation, we divide the 300 prompts in the color dataset into four difficulty bins based on their average score over 8 samples.
-Using the same set of hyperparameters across all prompts, Fig.
-5(c) shows that DFS allocates less compute to easier prompts (bin 1) and progressively more to harder ones with lower scores—demonstrating automatic, difficulty-aware compute allocation.
-Unlike snell2024scaling , which manually selects different search strategies based on predefined difficulty labels, our method requires no prior knowledge of difficulty.
-DFS dynamically adjusts compute allocation during inference using a single, unified algorithm and fixed parameters.
-5.3.2 Conditional ImageNet Generation
-Conditional generation of diffusion models using a unconditional foundation model has important applications ye2024tfg as foundation models obtain increasing expressivity.
-However, classifier guidance dhariwal2021diffusion requires training a noise-dependent classifier on samples generated by the diffusion model, which is difficult since it may require a lot of attempts for an unconditional model to generate the correct class.
-In this section, we treat the pretrained classifier as the verifier function and conduct inference-time scaling with both local and global search.
-The experimental details are in Appendix.
-E.3.2.
-Table 2: Best-of-N results for ImageNet conditional generation, with FID and Accuracy averaged across the two labels.
-#Particles Single verifier Double-verifier
+Following Peters et al.
+(2010), we formulate the offline RL problem as sampling from a Q-regularized distribution: $\pi^{*}(\bm{a}|\bm{s})\propto\mu(\bm{a}|\bm{s})e^{\beta Q_{\psi}(\bm{s},\bm{a})}$, where $Q_{\psi}$ is a learned Q-function representing preferences over actions, and $\mu$ is the behavior policy, which we model using a diffusion prior.
+We approach this problem from the inference-scaling perspective, composing off-the-shelf pretrained diffusion policy with ground-truth Q-functions without training.
+Among the baselines, Diffuser (Janner et al., 2022), QGPO (Lu et al., 2023), and D-QL (Wang et al., 2022) require additional joint training of the diffusion model and Q-function.
+To demonstrate the effectiveness of our method test-time search (TTS), we allow TFG (Ye et al., 2024) and DAS (Kim et al., 2025) to use up to twice the compute of TTS.
+As shown in Table 3, TTS achieves performance comparable to training-based baselines, while DAS and TFG struggles on the Medium and Medium-Replay datasets where the model’s capabilities are limited.
+Details see Sec.
+E.4.
+Environment TTS-distill DPPO
+Halfcheetah $\mathbf{51.6\pm 0.7}$ 47.8
+Hopper $\mathbf{98.8\pm 0.8}$ 92.8
+Walker2d $\mathbf{86.3\pm 0.5}$ 82.7
+Table 4: Results for policy distillation with TTS
+Advancing Online Learning with Policy Distillation
+To enhance the capabilities of the base model, we fine-tune it on actions generated by TTS and subsequently evaluate its one-shot sampling performance.
+We adapt the Medium datasets by replacing the original actions with those produced by TTS, and fine-tune the corresponding models from Lu et al.
+(2023) (see Sec.
+E.4.1 for details).
+Perhaps surprisingly, this simple offline distillation outperforms SOTA diffusion-based online RL methods such as DPPO (Ren et al., 2024), as shown in Table 4.
+To the best of our knowledge, this is the first work to demonstrate expert iteration with inference scaling for diffusion models, a direction that may be of independent interest.
+5.4 Mitigating reward hacking with double verifier
+In this section, we show that reward hacking and over-optimization can be mitigated by a double-verifier strategy: employing separate verifiers for local and global search.
+As observed in Shen et al.
+(2024), training-free guidance with verifier gradients is vulnerable to adversarial reward hacking: gradient guidance often over-optimize the verifier, causing it to classify them as belonging to the target class despite being out-of-distribution (OOD).
+Inspired by double-Q learning in reinforcement learning (Van Hasselt et al., 2016), we propose a double-verifier approach, assigning distinct verifiers to local and global search to efficiently reduce overestimation.
+Table 5: Results for ImageNet conditional generation with double-verifier using BoN and BFS.
+BoN-single BoN-double BFS-single BFS-double
 […]
-2 188.5 25.6% 178.4 29.7%
-4 171.5 31.8% 151.2 37.5%
-8 155.7 35.8% 127.8 49.2%
-Mitigating verifier hacking with double-verifier
-One of the major challenges of guiding the samples using the verifier gradient is that image classifiers are prone to adversarial examples szegedy2013intriguing .
-Inspired by the double-Q learning in RL van2016deep , we use two verifiers for local search gradient computation and global search respectively.
-Specifically, we use two different vision transformers xu2023imagereward with input resolution 224 and 384.
-For accuracy evaluation we use another classifier from touvron2021training .
-As shown in Table.
-2, using double-verifier consistently improves performance with Best-of-N sampling, with up to 2x less compute.
-Figure 6: Results for ImageNet conditional generation with BFS Fig.
-6(a) and Fig.
-6(b) show the FID and class accuracy of conditional samples.
-BFS consistently outperform Best-of-N sampling in efficiency.
-Improved Efficiency with BFS
-We investigate breadth-first search (BFS) as a means to improve the computational efficiency of Best-of-N sampling.
-All experiments use the double-verifier setup, where one verifier guides gradient-based local search and the other handles resampling or pruning.
-As shown in Fig.
-6, both BFS-Resampling and BFS-Pruning outperform Best-of-N in terms of compute efficiency.
-BFS-Pruning is more effective under limited compute budgets, while BFS-Resampling achieves higher performance with increased compute.
+4 171.5 31.8% 0.161 151.2 37.5% 0.164 156.2 36.1% 0.169 145.5 44.3% 0.181
+8 155.7 35.8% 0.165 127.8 49.2% 0.184 133.3 46.5% 0.183 118.2 55.9% 0.214
+We evaluate the proposed double-verifier on the challenging conditional ImageNet generation task, generating target-class samples from an unconditional model guided by a pretrained classifier defined only on clean samples.
+We report the FID and class accuracy of 256 samples.
+We also report the MSP score (Hendrycks & Gimpel, 2016), with higher MSP score indicating less OOD samples.
+As shown in Table  5, using double-verifier significantly improves performance over single verifier with 2x less compute.
+Also, double-verifier significantly reduces OOD samples indicated by the higher MSP score, improving robustness alongside efficiency.
+Visualizations and details see Appendix.
+E.5.
 6 Limitations and Conclusion
-Limitations
-Despite the performance improvement of inference-time scaling, it still requires more hyperparameter tuning than training-based methods.
-One approach is to design an adaptive hyperparameter tuning framework, which could increase or decrease automatically based on the current sample.
-For example we can tune the threshold of DFS automatically given the pass rate of samples.
-Another potential concern is that the generated samples may be used as adversarial examples to attack verifiers.
-Our method can be used to search for weaknesses of verifiers, which in turn can be used to train a more robust verifier.
-Conclusion
-In this work we propose a principled framework for inference-time scaling through classic search principles, unlocking the potential of scaling inference-time compute of diffusion models in a variety of downstream tasks.
-Inference-scaling can also be applied to training models via expert iteration, generating high-quality synthetic data beyond the model’s ability.
-We leave this as a interesting future direction.
+In this work, we present a unified and principled framework for inference-time scaling of diffusion models, jointly scaling local and global search.
+A potential limitation of our approach is that inference-time scaling still requires additional hyperparameters.
+To address this, heuristics such as evolutionary search can be adopted during hyperparameter tuning, which we leave for future work.
 Appendix A Appendix Overview
@@
 Appendix B Additional Related Works
-Inference-time scaling
-Scaling compute in inference-time with “slow thinking” has its long history grounded in cognitive science, known as “system 2” thinking [32, 63].
-In [52, 53], Newell and his colleagues formalized problem solving as tree search in a combinatorial space, and [93] uses tree-of-thoughts to enable LLM reasoning with multiple exploration paths, using BFS and DFS as strategic search algorithms.
-Recently, long chain-of-though (CoT) reasoning has demonstrated remarkable performance for LLM reasoning [30, 21], where the long CoT reasoning ability is incentivized through reinforcement learning [21].
-Notably, the CoT process demonstrates reasoning activities such as self-verification, backtracking and self-correction.
-Using a process reward model [39], we can also conduct explicit tree search without training the language model.
-[85] propose reward-balanced search (REBASE) which is a special instance of BFS, and [64] applied beam-search to difficult math problems, showing compute-optimal inference can be achieved via selecting different strategy for problems with different difficulty.
-We refer the readers to [79] for a comprehensive review.
-Inference-time scaling could also be used to improve the model itself, known as expert iteration.
-In [32], they proposed reinforcement learning with expert iteration in games, where the expert is constructed combining the base policy with Monte Carlo Tree Search (MCTS), and [86] applied expert iteration to automated theorem proving.
-Self improvements can also be achieved through iterative self-training [69, 96, 13, 61].
-Inference scaling in diffusion models
-The inference-time compute of diffusion models depends heavily on the number of denoising steps.
-[91] showed that recursive restart sampling can reduce cumulative error during sampling, which can be regarded as scaling the number of denoising steps.
-More recently, [62] proposed a Sequential Monte Carlo (SMC) [84, 99] style method, known as Feynman-Kac steering, which can be seen as a instance of BFS.
-Besides image generation, they also applied it to diffusion language models [23, 59].
-Additionally, [48] explored inference-time scaling of diffusion models with local zero order search and global search over paths for image generation.
-They also experimented with different verifiers, such as oracle verifiers, self-supervised verifiers, and studied the verifier-task alignment problem.
-There zero-order search can be understood as an uninformed version of local search which exhibits low efficiency, and when utilizing gradients, they need to back-propagate through the entire diffusion sampling chain, causing high computation and memory overheads.
-Compared with their work, we propose efficient gradient-based local search with theoretically grounded Langevin MCMC, which we show is crucial in many tasks.
-Also, we provide systematic experiments on the compute efficiency of global search methods.
-Their proposed methods can thus be understood as a instance within our search framework.
-Apart from search, [38] exploits the in-context learning abilities of foundation models to provide revision during sampling.
-Specifically, they leverage the multi-modal capabilities of VLMs to provide feedbacks on past generated images, and train the model to condition on past images and feedback.
-Diffusion models and applications
-Diffusion models [26, 68, 76] has shown great performance in generative modeling for continuous data domains, such as image [18, 12], videos [28] and molecules [90].
-Due to its expressive power on modeling multimodal and complex distributions, they have also been widely used as a decision prior in robotics.
-[31] proposes the first work on using diffusion models to generate plans.
-[10] uses the diffusion model for visuo-motor policy in robotics.
-Recently, a series of robotics foundation models utilize diffusion heads as action experts [5, 73], while [43] trains a end-to-end diffusion transformer for bimanual manipulation.
-In this work, we demonstrate that inference-time scaling can be especially helpful for decision making tasks with diffusion models.
-Control and alignment of diffusion models
-To align the diffusion model with flexible objectives, training-free guidance [94, 4, 11, 66, 95, 25] and compositional generation [14, 42] combines the diffusion models with classifiers or other diffusion models at inference time, while RL-based methods [98, 78, 89, 6] finetune the diffusion model using reward or preference signals.
-ControlNet [97] style approaches have also been used to add additional conditions for sampling, where [71] designs a control block for diffusion transformers, and [88] uses a combination of guidance and controlnet for controllable human motion generation.
+B.1 Discussions on Concurrent Works
+Adaptive Bi-directional Cyclic Diffusion (ABCD) (Lee et al., 2025) introduces a search-based inference-scaling framework that combines elements of both DFS and BFS.
+Unlike DFS, which relies on a quality threshold to determine backtracking, ABCD maintains a set of particles and backtracks by propagating them across all noise levels.
+The termination criterion is based on whether backtracking to higher noise levels improves sample quality.
+Compared with DFS, ABCD achieves smaller score estimation errors since particles are evaluated after full denoising, and it explores the generative space more thoroughly by combining BFS and DFS.
+However, ABCD cannot adaptively allocate compute across different instances due to its fixed termination condition, and it expends unnecessary compute on easier instances because it denoises a full set of particles regardless of sample quality.
+EvoSearch (He et al., 2025) proposes evolutionary search to scale inference compute for image and video generation.
+At selected timesteps, particles are fully denoised; high-scoring particles are retained, while low-scoring ones are perturbed via noise injection.
+This method improves over the FK-steering baseline (Singhal et al., 2025a).
+In contrast, our work develops a uniform and principled framework grounded in classical search heuristics.
+BFS and DFS serve as fundamental building blocks for more advanced search methods, and, crucially, we are the first to scale local search using annealed Langevin MCMC.
+This local search can be seamlessly combined with global search strategies, a joint scaling approach that we show is critical for success in challenging decision-making domains.
+B.2 Inference-Time Scaling with Classical Search
+We now review the use of classical search for inference-time scaling, including applications beyond diffusion models.
+Inference-time compute scaling, often viewed as “slow thinking,” has long-standing roots in cognitive science, where it corresponds to “system 2” reasoning (Kahneman, 2011; Sloman, 1996).
+In early work, Newell et al.
+(1959; 1972) formalized problem solving as tree search in combinatorial spaces.
+More recently, Yao et al.
+(2023) proposed tree-of-thoughts reasoning for LLMs, enabling exploration along multiple reasoning paths using BFS and DFS as strategic search primitives.
+B.3 Inference Scaling in Diffusion Models without Differentiable Optimization
+We next provide a more detailed overview of gradient-free inference-scaling approaches in diffusion models.
+Inference-time compute for diffusion models is largely determined by the number of denoising steps.
+Xu et al.
+(2023b) proposed recursive restart sampling to mitigate cumulative errors, effectively scaling the number of denoising steps.
+More recently, Singhal et al.
+(2025a); Kim et al.
+(2025) introduced Sequential Monte Carlo (SMC) approaches (Wu et al., 2023; Zhao et al., 2024), interpretable as BFS-style algorithms.
+Their implementations differ mainly in the choice of scoring functions and resampling strategies, with Kim et al.
+(2025) proposing an increasing tempering schedule.
+Ma et al.
+(2025) further explored inference-time scaling for image generation via a combination of local zero-order search and global path search, experimenting with different verifiers (oracle, self-supervised, and task-aligned) and analyzing verifier-task alignment.
+However, they did not systematically evaluate and improve the compute efficiency of search methods, while we demonstrate significantly improved efficiency with our novel design in search methods.
+Additionally, Yoon et al.
+(2025) proposed a Monte Carlo Tree Search-based diffusion framework tailored for trajectory synthesis, which denoises different trajectory segments at distinct steps.
+Beyond search-based methods, Li et al.
+(2025) leverage the in-context learning abilities of foundation models to enable revision during sampling.
+Their approach exploits multimodal VLM capabilities to provide feedback on intermediate generations and trains the model to condition on both prior samples and corresponding feedback.
+B.4 Gradient Guidance of Diffusion Models
+We review guidance methods for diffusion models that leverage gradients from verifiers or classifiers.
+To sample from a conditional distribution
+$\tilde{p}_{0}({\bm{x}}_{0})\propto p_{0}({\bm{x}}_{0})\,p(c|{\bm{x}}_{0}),$
+where $p$ denotes a classifier, one can define a conditional diffusion process $\{\tilde{q}_{t}({\bm{x}}_{t})\}_{t=0}^{T}$ with initial distribution $\tilde{p}_{0}$.
+The corresponding conditional score function is
+$\nabla_{{\bm{x}}_{t}}\log\tilde{q}_{t}({\bm{x}}_{t})=\nabla_{{\bm{x}}_{t}}\log q_{t}({\bm{x}}_{t})+\nabla_{{\bm{x}}_{t}}\log p_{t}(c|{\bm{x}}_{t}),$
+where $p_{t}(c|{\bm{x}}_{t})=\mathbb{E}_{{\bm{x}}_{0}|{\bm{x}}_{t}}[p(c|{\bm{x}}_{0})]$.
+Classifier guidance (Dhariwal & Nichol, 2021) trains a noise-dependent classifier $p_{t}(c|{\bm{x}}_{t})$ on noisy samples, using data from the diffusion model together with the base classifier.
+However, this training procedure is computationally expensive and impractical when the base diffusion model cannot generate target-class samples.
+Training-free guidance methods (Ye et al., 2024; Chung & Ye, 2022; Song et al., 2023; Yu et al., 2023; He et al., 2023) approximate the intractable $p_{t}(c|{\bm{x}}_{t})$ with $p(c|{\bm{x}}_{0|t})=p(c|\mathbb{E}_{{\bm{x}}_{0}|{\bm{x}}_{t}}[{\bm{x}}_{0}])$, which can be interpreted as a first-order approximation.
+These methods introduce bias, though recurrence has been shown to mitigate out-of-distribution artifacts and improve performance in classifier-guided generation (Ye et al., 2024).
+Nevertheless, the theoretical foundations and scaling behavior of recurrent guidance remain largely unexplored.
+In this work, we instead adopt Langevin MCMC sampling in place of score-based ODE sampling.
+This approach guarantees asymptotic exactness using only the base verifier, without requiring additional training.
+Relatedly, Du et al.
+(2024) proposed annealed energy-based models (EBMs) combined with iterative MCMC for inference scaling and reasoning, but their method cannot be applied to pretrained diffusion models and verifiers.
+For the first time, we establish a connection between annealed Langevin MCMC and recurrent guidance, thereby enabling inference scaling with pretrained foundation models.
 Appendix C Details about local search with Langevin MCMC
@@
 C.1 Langevin MCMC as gradient flow in measure space
-Following [83], the Langevin SDE in sample space corresponds to gradient flow of the KL-divergence in measure space. Here we provide a brief overview.
+Following Wibisono (2018), the Langevin SDE in sample space corresponds to gradient flow of the KL-divergence in measure space. Here we provide a brief overview.
 Define our target distribution that we wish to sample from as $\nu$, and the distribution of our current sample as $\rho$. We define the KL-divergence (relative entropy) as:
@@
 That is, if ${\bm{x}}_{t}\sim\rho_{t}$ evolves according to the dynamics in Eq.
-7, then the measure $\rho({\bm{x}},t)=\rho_{t}$ evolves according to the PDE in Eq.
-6, conducting gradient optimization in measure space.
+8, then the measure $\rho({\bm{x}},t)=\rho_{t}$ evolves according to the PDE in Eq.
+7, conducting gradient optimization in measure space.
 In practice, we implement Eq. 8 through discretization, which is known as the unadjusted Langevin algorithm (ULA):
@@
 with ${\bm{\epsilon}}^{i}\sim\mathcal{N}(\bm{0},\bm{I})$. When $\eta\rightarrow 0$, the ULA converges to Langevin SDE, providing exact sampling.
-Previous works [17, 9] show that ULA can efficiently converge to the target measure $\nu$ if $\nu$ is log-concave and smooth.
+Previous works (Durmus & Moulines, 2015; Cheng & Bartlett, 2018) show that ULA can efficiently converge to the target measure $\nu$ if $\nu$ is log-concave and smooth.
 However, when facing complex and multimodal distributions, we can only guarantee convergence to the concave vicinity.
 C.2 Annealed Langevin MCMC Sampling
-Langevin MCMC have been used to perform implicit sampling in energy-based models [16] and score-based models [67].
-However, these methods suffer from inaccurate score estimation and low density regions [67].
-In [67] they propose to perturb the data with gaussian noise, eventually smoothing the data distribution:
+Langevin MCMC have been used to perform implicit sampling in energy-based models (Du & Mordatch, 2019) and score-based models (Song & Ermon, 2020).
+However, these methods suffer from inaccurate score estimation and low density regions (Song & Ermon, 2020).
+In Song & Ermon (2020) they propose to perturb the data with gaussian noise, eventually smoothing the data distribution:
 $q({\bm{x}}_{t})=\int_{{\bm{x}}_{0}}p_{0}({\bm{x}}_{0})\mathcal{N}({\bm{x}}_{t};{\bm{x}}_{0},\sigma_{t}^{2}\bm{I})\,$,
 and creating a sequence of annealed distributions $\left\{q({\bm{x}}_{t})\right\}_{t=0}^{T}$ which converges to $p_{0}({\bm{x}}_{0})$.
-Since they are smoothed by gaussian noise, we can improve the mixing time of Langevin MCMC on multimodal distributions by sampling from these intermediate distributions, sharing similar spirits with simulated annealing [33].
-In [14], they extend this method to compositional generation of diffusion models.
+Since they are smoothed by gaussian noise, we can improve the mixing time of Langevin MCMC on multimodal distributions by sampling from these intermediate distributions, sharing similar spirits with simulated annealing (Kirkpatrick et al., 1983).
+In Du et al.
+(2023), they extend this method to compositional generation of diffusion models.
 Specifically, we consider sampling from a product distribution $p^{\text{prod}}_{0}({\bm{x}}_{0})\propto p_{0}^{1}({\bm{x}}_{0})p_{0}^{2}({\bm{x}}_{0})$, where $p_{0}^{1}({\bm{x}}_{0})$ and $p_{0}^{2}({\bm{x}}_{0})$ are distributions of different diffusion models.
@@
 C.3.2, we prove that adding the guidance term is defining an annealing path that biases towards high score regions.
+Finally, we provide a rigorous convergence analysis in Sec.
+C.3.3.
 C.3.1 Equivalence between Langevin MCMC and naive recurrence
-Consider the diffusion process with the following stochastic interpolant [47]:
+Consider the diffusion process with the following stochastic interpolant (Ma et al., 2024):
 ${\bm{x}}_{t}=\alpha_{t}{\bm{x}}_{0}+\sigma_{t}{\bm{\epsilon}}\,.$
@@
 DDIM sampler
-When using DDIM [65] sampler, we have the reverse step as:
+When using DDIM (Song et al., 2020a) sampler, we have the reverse step as:
 ${\bm{x}}_{t-1}=\frac{\alpha_{t-1}}{\alpha_{t}}{\bm{x}}_{t}+\sigma_{t}^{2}\left(\frac{\alpha_{t-1}}{\alpha_{t}}-\frac{\sigma_{t-1}}{\sigma_{t}}\right)s({\bm{x}}_{t},t)\,$, (11)
@@
 $={\bm{x}}_{t}^{i}+\sigma_{t}^{2}\left(1-e^{\lambda_{t}-\lambda_{t-1}}\right)s({\bm{x}}_{t}^{i},t)+\sigma_{t}\sqrt{\left(1-e^{\lambda_{t}-\lambda_{t-1}}\right)\left(1+e^{\lambda_{t}-\lambda_{t-1}}\right)}{\bm{\epsilon}}^{i}\,$,
-where $1+e^{\lambda_{t}-\lambda_{t-1}}\rightarrow 2$ when $T\rightarrow\infty$ and denoising step size approaches 00, as $\lambda_{t}-\lambda_{t-1}\rightarrow 0$.
+where $1+e^{\lambda_{t}-\lambda_{t-1}}\rightarrow 2$ when $T\rightarrow\infty$ and denoising step size approaches 0, as $\lambda_{t}-\lambda_{t-1}\rightarrow 0$.
 DDPM sampler
-In DDPM [26], we parametrize the posterior distribution as:
+In DDPM (Ho et al., 2020), we parametrize the posterior distribution as:
 $p_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})=\mathcal{N}\left({\bm{x}}_{t-1};\mu_{\theta}({\bm{x}}_{t},t),\Sigma_{\theta}({\bm{x}}_{t},t)\right)\,$, (12)
@@
 $\mu_{\theta}({\bm{x}}_{t},t)=\frac{\alpha_{t-1}}{\alpha_{t}}{\bm{x}}_{t}+\left(\sigma_{t}^{2}\frac{\alpha_{t-1}}{\alpha_{t}}-\sigma_{t-1}^{2}\frac{\alpha_{t}}{\alpha_{t-1}}\right)s({\bm{x}}_{t},t)\,.$
-[26] parameterizes the posterior variance as $\Sigma_{\theta}({\bm{x}}_{t},t)=\beta_{t}\bm{I}$ or $\Sigma_{\theta}({\bm{x}}_{t},t)=\tilde{\beta}_{t}\bm{I}$:
+Ho et al. (2020) parameterizes the posterior variance as $\Sigma_{\theta}({\bm{x}}_{t},t)=\beta_{t}\bm{I}$ or $\Sigma_{\theta}({\bm{x}}_{t},t)=\tilde{\beta}_{t}\bm{I}$:
 $\beta_{t}$ $=\alpha_{t}^{2}\left(\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}-\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}}\right)\,$,
 $\tilde{\beta}_{t}$ $=\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\beta_{t}\,$,
-while [54] propose to train the posterior variance as $\Sigma_{\theta}({\bm{x}}_{t},t)=\exp\left(v\log\beta_{t}+(1-v)\log\tilde{\beta}_{t}\right)$.
+while Nichol & Dhariwal (2021) propose to train the posterior variance as $\Sigma_{\theta}({\bm{x}}_{t},t)=\exp\left(v\log\beta_{t}+(1-v)\log\tilde{\beta}_{t}\right)$.
 Thus, a backward step can be written as:
@@
 $={\bm{x}}_{t}^{i}+\beta_{t}s({\bm{x}}_{t}^{i},t)+\sqrt{\Sigma_{\theta}({\bm{x}}_{t},t)+\beta_{t}\bm{I}}{\bm{\epsilon}}^{i}\,$,
-where $\Sigma_{\theta}({\bm{x}}_{t},t)\rightarrow\beta_{t}\bm{I}$ when $T\rightarrow\infty$, and the denoising step size approaches 00.
+where $\Sigma_{\theta}({\bm{x}}_{t},t)\rightarrow\beta_{t}\bm{I}$ when $T\rightarrow\infty$, and the denoising step size approaches 0.
 Putting together
 In general, we can write the recurrence step as:
-${\bm{x}}_{t}^{i+1}={\bm{x}}_{t}^{i}+a_{t}s({\bm{x}}_{t}^{i},t)+b_{t}{\bm{\epsilon}}^{i}\,.$ (12)
-with $a_{t}\rightarrow 0$, $b_{t}\rightarrow 0$ and $2a_{t}\rightarrow b_{t}^{2}$ as the denoising step size approaches 00.
-Thus, it can be seen as a approximation of the ULA in Eq.
-8, and also a discretization of the Langevin SDE in Eq.
-7.
+${\bm{x}}_{t}^{i+1}={\bm{x}}_{t}^{i}+a_{t}r_{t}s({\bm{x}}_{t}^{i},t)+\sqrt{2a_{t}}{\bm{\epsilon}}^{i}\,.$ (13)
+with $a_{t}\rightarrow 0$ and $r_{t}\rightarrow 1$ as the denoising step size approaches 0:
+  • For DDIM sampler, we have $a_{t}=\frac{1}{2}\alpha_{t}^{2}\left(\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}-\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}}\right)$ and $r_{t}=\frac{2}{1+e^{\lambda_{t}-\lambda_{t-1}}}$.
+  • For DDPM sampler, we have $a_{t}=\frac{1}{2}\alpha_{t}^{2}\left(\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}-\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}}\right)$ and $1\leq r_{t}\leq\frac{2}{1+\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}}$.
+Thus, it can be seen as a approximation of the ULA in Eq. 9, and also a discretization of the Langevin SDE in Eq. 8.
 C.3.2 Annealed Langevin MCMC with guidance
-When applying training free guidance [94] during the recurrence, we have:
-${\bm{x}}_{t}^{i+1}={\bm{x}}_{t}^{i}+a_{t}s({\bm{x}}_{t}^{i},t)+b_{t}{\bm{\epsilon}}^{i}+\bm{\Delta}({\bm{x}}_{t},t)\,,$
+When applying training free guidance (Ye et al., 2024) during the recurrence, we have:
+${\bm{x}}_{t}^{i+1}={\bm{x}}_{t}^{i}+a_{t}r_{t}s({\bm{x}}_{t}^{i},t)+\sqrt{2a_{t}}{\bm{\epsilon}}^{i}+\bm{\Delta}({\bm{x}}_{t},t)\,$,
 where $a_{t}$, $b_{t}$ are the coefficients of the recurrence equation in Eq.
@@
 $=\mu_{t}\frac{\sigma_{t}^{2}}{\bm{\Sigma}_{0|t}}\nabla_{{\bm{x}}_{t}}\log f({\bm{x}}_{0|t})\,$,
-where the second Equation follows from Lemma 3.3 in [94].
-Thus, there exists a set of functions $\hat{f}_{t}^{\text{mean}}({\bm{x}}_{t})$ such that $\nabla_{{\bm{x}}_{t}}\log\hat{f}_{t}^{\text{mean}}({\bm{x}}_{t})=\frac{\sigma_{t}^{2}}{\bm{\Sigma}_{0|t}}\nabla_{{\bm{x}}_{t}}\log f({\bm{x}}_{0|t})$, and we can see that when $t\rightarrow 0$,$\nabla_{{\bm{x}}_{t}}\log\hat{f}_{t}^{\text{mean}}({\bm{x}}_{t})=\nabla_{{\bm{x}}_{0}}\log f({\bm{x}}_{0})$.
+where the second Equation follows from Lemma 3.3 in Ye et al.
+(2024).
+Thus, there exists a set of functions $\hat{f}_{t}^{\text{mean}}({\bm{x}}_{t})$ such that $\nabla_{{\bm{x}}_{t}}\log\hat{f}_{t}^{\text{mean}}({\bm{x}}_{t})=\frac{\sigma_{t}^{2}}{\bm{\Sigma}_{0|t}}\nabla_{{\bm{x}}_{t}}\log f({\bm{x}}_{0|t})$, and we can see that when $t\rightarrow 0$$\nabla_{{\bm{x}}_{t}}\log\hat{f}_{t}^{\text{mean}}({\bm{x}}_{t})=\nabla_{{\bm{x}}_{0}}\log f({\bm{x}}_{0})$.
 If we additionally incorporate the ‘implicit dynamics’ in Line.
@@
 Thus, recurrence with guidance can be written as:
-${\bm{x}}_{t}^{i+1}$ $={\bm{x}}_{t}^{i}+a_{t}s({\bm{x}}_{t}^{i},t)+b_{t}{\bm{\epsilon}}^{i}+c_{t}\nabla_{{\bm{x}}_{t}}\log\hat{f}_{t}({\bm{x}}_{t})\,,$
-which, in the continuous limit of $a_{t},b_{t},c_{t}\rightarrow 0$, converges to the Langevin SDE:
-$d{\bm{x}}_{t}^{i}=\nabla_{{\bm{x}}_{t}^{i}}\log\left(q_{t}({\bm{x}}_{t}^{i})\hat{f}_{t}({\bm{x}}_{t}^{i})\right)di+\sqrt{2}d\bm{w}^{i}\,,$
-where $i$ is the time index of recurrence steps. Thus, we have defined the annealing path as $\tilde{q}_{t}({\bm{x}}_{t})=q_{t}({\bm{x}}_{t})\hat{f}_{t}({\bm{x}}_{t})\,,~{}t=1,2,\cdots,T$.
+${\bm{x}}_{t}^{i+1}$ $={\bm{x}}_{t}^{i}+a_{t}r_{t}s({\bm{x}}_{t}^{i},t)+\sqrt{2a_{t}}{\bm{\epsilon}}^{i}+c_{t}\nabla_{{\bm{x}}_{t}}\log\hat{f}_{t}({\bm{x}}_{t})$
+$={\bm{x}}_{t}^{i}+a_{t}r_{t}\nabla_{{\bm{x}}_{t}}\log q_{t}({\bm{x}}_{t})\hat{f}_{t}({\bm{x}}_{t})^{c_{t}/a_{t}r_{t}}+\sqrt{2a_{t}}{\bm{\epsilon}}^{i}\,$,
+Thus, we have defined the annealing path as $\tilde{q}_{t}({\bm{x}}_{t})=q_{t}({\bm{x}}_{t})\hat{f}_{t}({\bm{x}}_{t})^{c_{t}/a_{t}r_{t}}\,,~t=1,2,\cdots,T$.
+C.3.3 Convergence analysis
+In this section, we provide a rigorous convergence analysis of recurrence to the target distribution $\tilde{q}_{t}({\bm{x}}_{t})$.
+Theorem 1.
+Suppose $\tilde{q}_{t}({\bm{x}}_{t})$ has bounded support, is $\alpha$-strongly log-concave and $L$-log-smooth, and $\nabla^{2}\log\tilde{q}_{t}$ is $M$-Lipschitz.
+Denote ${\bm{x}}_{t}^{N_{\text{recur}}}$ as the sample after $N_{\text{recur}}$} steps of recurrence, we can bound the Wasserstein distance between the distribution of ${\bm{x}}_{t}^{N_{\text{recur}}}$ and $\tilde{q}_{t}$ as:
+$W_{2}(p({\bm{x}}_{t}^{N_{\text{recur}}}),\tilde{q}_{t})=\mathcal{O}\left(\sqrt{\lambda_{t-1}-\lambda_{t}}+e^{-2\lambda_{t}}-e^{-2\lambda_{t-1}}+(1-e^{-2\lambda_{t}}+e^{-2\lambda_{t-1}})^{N_{\text{recur}}}\right)\,$,
+where $\lambda_{t}=\log\frac{\alpha_{t}}{\sigma_{t}}$ is half of the log SNR.
+Proof.
+Recall recurrence is equivalent to the following recursion equation, with $a_{t}$, $r_{t}$ being sampler-dependent parameters in Eq. 13:
+${\bm{x}}_{t}^{i+1}$ $={\bm{x}}_{t}^{i}+a_{t}r_{t}\nabla_{{\bm{x}}_{t}}\log\tilde{q}_{t}({\bm{x}}_{t}^{i})+\sqrt{2a_{t}}{\bm{\epsilon}}^{i}$
+$={\bm{x}}_{t}^{i}+a_{t}\nabla_{{\bm{x}}_{t}}\log\tilde{q}_{t}({\bm{x}}_{t}^{i})^{r_{t}}+\sqrt{2a_{t}}{\bm{\epsilon}}^{i}\,.$
+Thus, recurrence is equivalent to running unadjusted Langevin algorithm (ULA) on the tempered distribution $p^{\text{tempered}}\propto\tilde{q}_{t}^{r_{t}}$.
+Using Lemma 1 and Lemma 2 from Wibisono (2018), given the regularity conditions on $\tilde{q}_{t}$, we can bound the discretization error from ULA as:
+$W_{2}(p^{\text{tempered}},p({\bm{x}}_{t}^{N_{\text{recur}}}))$ $=\mathcal{O}\left(a_{t}+(1-a_{t})^{N_{\text{recur}}}\right)$
+$=\mathcal{O}\left(\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}-\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}}+(1-\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}+\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}})^{N_{\text{recur}}}\right)$
+$=\mathcal{O}\left(e^{-2\lambda_{t}}-e^{-2\lambda_{t-1}}+(1-e^{-2\lambda_{t}}+e^{-2\lambda_{t-1}})^{N_{\text{recur}}}\right)\,.$
+To bound $W_{2}(p^{\text{tempered}},\tilde{q}_{t})$, we first bound the TV distance as $\text{TV}(p^{\text{tempered}},\tilde{q}_{t}$).
+Denote $Z(r_{t})=\int\tilde{q}_{t}(x)^{r_{t}}dx$, and $\psi(r)=\log Z(r)$. We have:
+$\text{KL}(p^{\text{tempered}},\tilde{q}_{t})$
+$=\mathbb{E}_{p^{\text{tempered}}}\left[r_{t}\log\tilde{q}_{t}-\psi(r_{t})-\log\tilde{q}_{t}\right]$
+$\leq\mathcal{O}\left((r_{t}-1)\mathbb{E}_{p^{\text{tempered}}}\left[\log\tilde{q}_{t}-\psi^{\prime}(r)\right]+\psi^{\prime\prime}(r)(r_{t}-1)^{2}\right)\,$,
+where the last step is using Taylor expansion.
+Since $\psi^{\prime}(r)=\mathbb{E}_{p^{\text{tempered}}}[\log\tilde{q}_{t}]$ and $\psi^{\prime\prime}(r)=\text{Var}_{p^{\text{tempered}}}[\log\tilde{q}_{t}]$.
+Using the Pinsker inequality, we have $\text{TV}(p^{\text{tempered}},\tilde{q}_{t})\leq\frac{1}{2}\sqrt{\text{KL}(p^{\text{tempered}},\tilde{q}_{t})}\leq\mathcal{O}\left((r_{t}-1)\sqrt{\text{Var}_{p^{\text{tempered}}}[\log\tilde{q}_{t}]}\right)=\mathcal{O}(r_{t}-1)$, given the finiteness of $\text{Var}_{p^{\text{tempered}}}[\log\tilde{q}_{t}]$ under bounded support.
+Following Proposition 7.10 in Villani (2003) for distributions with bounded support, we have:
+$W_{2}(p^{\text{tempered}},\tilde{q}_{t})$
+$=\mathcal{O}\left(\sqrt{\text{TV}(p^{\text{tempered}},\tilde{q}_{t})}\right)$
+$=\mathcal{O}\left(\sqrt{r_{t}-1}\right)$
+$=\mathcal{O}\left(\sqrt{1-\min\left(\frac{\alpha_{t}\sigma_{t-1}}{\alpha_{t-1}\sigma_{t}},\frac{\sigma_{t-1}^{2}}{\sigma_{t}^{2}}\right)}\right)$
+$=\mathcal{O}\left(\sqrt{\log\frac{\sigma_{t}}{\sigma_{t-1}}+\max\left(\log\frac{\alpha_{t-1}}{\alpha_{t}},\log\frac{\sigma_{t}}{\sigma_{t-1}}\right)}\right)$
+$=\mathcal{O}\left(\sqrt{\log\frac{\alpha_{t-1}}{\alpha_{t}}+\log\frac{\sigma_{t}}{\sigma_{t-1}}}\right)$
+$=\mathcal{O}\left(\sqrt{\lambda_{t-1}-\lambda_{t}}\right)\,$,
+where we recall the definition of $r_{t}$ in Eq. 13.
+Putting together the bound on $W_{2}(p^{\text{tempered}},p({\bm{x}}_{t}^{N_{\text{recur}}}))$ and $W_{2}(p^{\text{tempered}},\tilde{q}_{t})$, we obtain our desired bound. ∎
 C.4 Relationship between Langevin MCMC and gradient ascent
@@
 Recall the KL-divergence objective in Eq.
-5, which can be further decomposed when we are sampling from a compositional distribution of $p_{0}({\bm{x}}_{0})$ and verifier $f({\bm{x}}_{0})$, with $\nu\propto p_{0}\cdot f$:
+6, which can be further decomposed when we are sampling from a compositional distribution of $p_{0}({\bm{x}}_{0})$ and verifier $f({\bm{x}}_{0})$, with $\nu\propto p_{0}\cdot f$:
 $H_{\nu}(\rho)=\mathbb{E}_{\rho}[-\log f]+H_{p_{0}}(\rho)+\log Z\,.$
@@
 C.5 Implementing Local Search with TFG hyper-parameter space
-Due to the equivalence between annealed Langevin MCMC and training-free guidance with recurrence, we can implement local search with Langevin MCMC using the TFG framework of [94], efficiently searching the hyperparameters.
+Due to the equivalence between annealed Langevin MCMC and training-free guidance with recurrence, we can implement local search with Langevin MCMC using the TFG framework of Ye et al.
+(2024), efficiently searching the hyperparameters.
 Here we provide a overview of the algorithm and design space.
@@
 1:Input: Unconditional diffusion model ${\bm{\epsilon}}_{\theta}$, verifier $f$, guidance strength $\bm{\rho},\bm{\mu},\bar{\gamma}$, number of steps $T,N_{\text{recur}},N_{\text{iter}}$
-2:  ${\bm{x}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
+[…]
 3:for $t=T,\cdots,1$ do
@@
 7:   $\bm{\Delta}_{\text{var}}=\rho_{t}\nabla_{{\bm{x}}_{t}}\log\tilde{f}({\bm{x}}_{0|t})$
-8:        $\bm{\Delta}_{\text{mean}}=\bm{\Delta}_{\text{mean}}+\mu_{t}\alpha_{t}\nabla_{{\bm{x}}_{0|t}}\log\tilde{f}({\bm{x}}_{0|t}+\bm{\Delta}_{\text{mean}})$trianglerightIterate $N_{\text{iter}}$ times starting from $\bm{\Delta}_{\text{mean}}=\bm{0}$
-9:        ${\bm{x}}_{t-1}=\text{Sample}({\bm{x}}_{t},{\bm{x}}_{0|t},t)+\frac{\alpha_{t-1}}{\alpha_{t}}\left(\bm{\Delta}_{\text{var}}+\bm{\Delta}_{\text{mean}}\right)$triangleright Sample follows DDIM or DDPM
-10:        ${\bm{x}}_{t}\sim\mathcal{N}\left(\frac{\alpha_{t}}{\alpha_{t-1}}{\bm{x}}_{t-1},\alpha_{t}^{2}\left(\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}-\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}}\right)\bm{I}\right)$triangleright Recurrent strategy
+[…]
+9:   ${\bm{x}}_{t-1}=\text{Sample}({\bm{x}}_{t},{\bm{x}}_{0|t},t)+\frac{\alpha_{t-1}}{\alpha_{t}}\left(\bm{\Delta}_{\text{var}}+\bm{\Delta}_{\text{mean}}\right)$ $\triangleright$ Sample follows DDIM or DDPM
+10:   ${\bm{x}}_{t}\sim\mathcal{N}\left(\frac{\alpha_{t}}{\alpha_{t-1}}{\bm{x}}_{t-1},\alpha_{t}^{2}\left(\frac{\sigma_{t}^{2}}{\alpha_{t}^{2}}-\frac{\sigma_{t-1}^{2}}{\alpha_{t-1}^{2}}\right)\bm{I}\right)$ $\triangleright$ Recurrent strategy
 11:  end for
@@
 13:Output: Conditional sample ${\bm{x}}_{0}$
-For time varying schedules $\rho_{t}$, $\mu_{t}$, we follow [94] and propose to use either the ‘increase’ schedule:
+For time varying schedules $\rho_{t}$, $\mu_{t}$, we follow Ye et al. (2024) and propose to use either the ‘increase’ schedule:
 $s_{t}=T\frac{\alpha_{t}/\alpha_{t-1}}{\sum_{t=1}^{T}\alpha_{t}/\alpha_{t-1}}\,$, (14)
@@
 D.1 BFS-Based Search
-We present the pseudo code for BFS in Alg.
-3.
-The core hyperparameters of BFS are the temperature $\tau_{t}$ and the evaluation time steps $\mathcal{S}$.
-These parameters enable a smooth transition from Best-of-N to best-first search.
-In general, a higher $\tau_{t}$ and evaluation at earlier time steps bias the process toward high-score samples, while a lower temperature and evaluation at later (smaller) time steps approximate the Best-of-N approach.
-In practice, we typically evaluate at fractions of the sampling steps, with $\mathcal{S}=\left\{\frac{3}{4}T,\frac{1}{2}T,\frac{1}{4}T\right\}$.
-We perform a grid search over the temperature values to determine an appropriate magnitude using a small number of particles, and then apply this configuration when scaling up the number of particles.
-We use the ‘increase’ schedule in Eq.
-13 for $\tau_{t}$, so that $\tau_{T}\leq\tau_{T-1}\leq\cdots\leq\tau_{0}$, reflecting increased confidence in the verifier evaluation at smaller time steps.
+We present the pseudo code for BFS in Alg. 3. In practice, we can evaluate and resample the particles at a fixed subset of time steps, with detailed hyper-parameters in the details of each experiment.
 Algorithm 3 Diffusion BFS
@@
 Verifier $f$.
-  BFS input: Set of evaluation time steps $\mathcal{S}$. Temperature schedule $\tau_{t}$. Budget of particles $K$.
-  Init: Random sample $K$ particles ${\bm{x}}_{T}^{k}\sim\mathcal{N}(\bm{0},\bm{I}),k=1,2,\cdots K$. Set the number of particles $K_{T}=K$
+BFS input: Tempering schedule $\tau_{t}$. Budget of particles NN. Scoring function and resampling method.
+Init: Random sample NN particles ${\bm{x}}_{T}^{k}\sim\mathcal{N}(\bm{0},\bm{I}),k=1,2,\cdots N$.
 for $t=T,\cdots,1$ do
-     if $t\in\mathcal{S}$ then
-        for $k=1,2,\cdots,K_{t}$ do
+  for $k=1,2,\cdots,N$ do
+Estimation
 Estimate the conditional mean: ${\bm{x}}_{0|t}^{k}=\frac{{\bm{x}}_{t}^{k}-\sigma_{t}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t}^{k},t)}{\alpha_{t}}$.
-Compute the verifier score $f({\bm{x}}_{0|t}^{k})$.
-           Compute the number of children $n_{t}^{k}$:
-           if Resample then
-              $n_{t}^{k}=\texttt{Round}\left(K_{t}\frac{f({\bm{x}}_{0|t}^{k})^{\tau_{t}}}{\sum_{i=1}^{K_{t}}f({\bm{x}}_{0|t}^{i})^{\tau_{t}}}\right)$
-           else if Prune then
-              $n_{t}^{k}=\min\left(1,\texttt{Round}\left(K_{t}\frac{f({\bm{x}}_{0|t}^{k})^{\tau_{t}}}{\sum_{i=1}^{K_{t}}f({\bm{x}}_{0|t}^{i})^{\tau_{t}}}\right)\right)$
-           end if
+Compute the verifier score estimate $f({\bm{x}}_{0|t}^{k})$.
+Scoring
+Score the particles according to the scoring functions:
+$\textbf{Current}:\widehat{f}({\bm{x}}_{t}^{k})=\tau_{t}f({\bm{x}}_{0|t}^{k}),$
+$\textbf{Difference}:\widehat{f}({\bm{x}}_{t}^{k})=\tau_{t}f({\bm{x}}_{0|t}^{k})-\widehat{f}_{\text{prev}}^{k},$
+$\textbf{Max}:\widehat{f}({\bm{x}}_{t}^{k})=\max\left(\tau_{t}f({\bm{x}}_{0|t}^{k}),\widehat{f}_{\text{prev}}^{k}\right).$
+Resampling
+Compute the weights
+$(w_{t}^{1},w_{t}^{2},\cdots,w_{t}^{N})=\text{softmax}\left(\widehat{f}({\bm{x}}_{t}^{1}),\widehat{f}({\bm{x}}_{t}^{2}),\cdots\widehat{f}({\bm{x}}_{t}^{N})\right)\,$,
+and resample the children
+$(n_{t}^{1},n_{t}^{2}\cdots,n_{t}^{N})=\text{Resample}(N;w_{t}^{1},w_{t}^{2},\cdots,w_{t}^{N}),$
+where Resample can be Multinomial resampling or SSP resampling (Alg.4).
   end for
-     else
-        $n_{t}^{k}=1$
-     end if
-     for $k=1,\cdots,K_{t}$ do
-        Sample $n_{t}^{k}$ particles from ${\bm{x}}_{t}^{k}$: ${\bm{x}}_{t-1}^{j}\sim\tilde{p}_{\theta}(\cdot|{\bm{x}}_{t}^{k})\,,j=1,2,\cdots,n_{t}^{k}$
+  for $k=1,\cdots,N$ do
+   Sample $n_{t}^{k}$ children particles from ${\bm{x}}_{t}^{k}$: ${\bm{x}}_{t-1}^{j}\sim\tilde{p}_{\theta}(\cdot|{\bm{x}}_{t}^{k})\,,j=1,2,\cdots,n_{t}^{k}$
   end for
-     The number of particles at next noise level: $K_{t-1}=\sum_{k=1}^{K_{t}}n_{t}^{k}$
+  Update the score buffers for computing the Difference and Max score at next timestep.
+$\widehat{f}_{\text{prev}}^{k}=\left\{\begin{aligned} &\widehat{f}\left({\bm{x}}_{t}^{\text{parent(k)}}\right)\quad\text{if}~~\text{scoring function}~~\text{is}~~\textbf{Max}\\ &\tau_{t}f\left({\bm{x}}_{0|t}^{\text{parent(k)}}\right)\quad\text{if}~~\text{scoring function}~~\text{is}~~\textbf{Difference}\end{aligned}\right$.
 end for
-  Return ${\bm{x}}_{0}=\operatorname*{argmax}_{k=1,\cdots,K_{0}}f({\bm{x}}_{0}^{k})$
+Return ${\bm{x}}_{0}=\operatorname*{argmax}_{{\bm{x}}_{0}^{1},\cdots,{\bm{x}}_{0}^{N}}f({\bm{x}}_{0}^{k})$
+D.1.1 Overview of design space
+Tempering
+When estimating the particle score via the denoised output $f({\bm{x}}_{0|t}^{k})$, for larger $t$ the estimation often has larger bias, due to the fact that $\mathbb{E}[f({\bm{x}}_{0})|{\bm{x}}_{t}]\neq f(\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}])=f({\bm{x}}_{0|t})$.
+Thus, we propose to re-weight the score estimates with a tempering schedule $\tau_{t}$ before scoring the particles.
+For a increasing schedule $\tau_{T}<\tau_{T-1}<\cdots<\tau_{1}<\tau_{0}$, the estimated scores have more influence at smaller time steps with higher weights, where the estimation has less bias.
+For the increasing tempering schedule, we simply adopt the design in DAS (Kim et al., 2025) with $\tau_{t}=\left((1+\gamma)^{T-t}-1\right)\tau$, with $\gamma$ searched in $\left\{0.008,0.024\right\}$.
+We also consider the deterministic particle selection in SVDD (Li et al., 2024) where we only retain the highest scoring particle, which can be seen as $\tau_{t}=\infty$.
+For the constant tempering schedule we simply set $\tau_{t}=\tau$, without any tempering on different time steps.
+Scoring
+Given the score estimates $f({\bm{x}}_{0|t})$, one can construct scoring rules based on the scores of the entire path $\left\{f({\bm{x}}_{0|s}^{k})\right\}_{s=T}^{t}$.
+In traditional sequential-monte-carlo (SMC) based methods, the score is computed as the difference between two consecutive evaluations: $\widehat{f}({\bm{x}}_{t}^{k})=\tau_{t}f({\bm{x}}_{0|t}^{k})-\tau_{t+1}f({\bm{x}}_{0|t+1}^{k})$.
+However, in FK-steering (Singhal et al., 2025a), they pointed out that the difference potential is unfair for sample path that saturates earlier, reaching a high reward at earlier time steps.
+Thus they propose the max scoring: $\widehat{f}({\bm{x}}_{t}^{k})=\max_{s\geq t}\tau_{t}f({\bm{x}}_{0|s}^{k})$.
+We also consider the simple baseline of current score: $\widehat{f}({\bm{x}}_{t}^{k})=\tau_{t}f({\bm{x}}_{0|t}^{k})$.
+Resampling
+Given the scores $\widehat{f}({\bm{x}}_{t}^{k})$, we can compute the softmax logits of particles via $\left\{w_{t}^{k}\right\}_{k=1}^{N}=\text{softmax}\left\{\widehat{f}({\bm{x}}_{t}^{k})\right\}_{k=1}^{N}$.
+Then, we resample the particles according to the logits, and obtain the number of children for each particle $n_{t}^{k}$.
+The sets $\left\{n_{t}^{k}\right\}_{k=1}^{N}$ should satisfy the following constraints:
+$\sum_{k=1}^{N}n_{t}^{k}=$ $N$
+$\mathbb{E}[n_{t}^{k}]=$ $Nw_{t}^{k}$
+Most previous approaches simply adopt the multinomial resampling, with $\left\{n_{t}^{k}\right\}_{k=1}^{N}\sim\text{Multinomial}\left(N;\left\{w_{t}^{k}\right\}_{k=1}^{N}\right)$.
+However, this sampling process introduces large variance since every component is sampled independently.
+To reduce variance, for example, residual resampling only resamples the residuals $\tilde{n}_{t}^{k}$ from NRN-R with weights $\tilde{w}_{t}^{k}=\frac{Nw_{t}^{k}-[Nw_{t}^{k}]}{N-R}$, where $R=\sum_{k=1}^{N}[Nw_{t}^{k}]$.
+Then, we obtain $n_{t}^{k}=[Nw_{t}^{k}]+\tilde{n}_{t}^{k}$.
+This way we reduce the randomness in $n_{t}^{k}$.
+The SSP resampling views the sampling process as randomized rounding on the expectations $Nw_{t}^{k}$, with pseudo code from Gerber et al. (2020) shown below in Alg. 4.
+Algorithm 4 SSP resampling
+Inputs: $u\in[0,1]^{{N}}$ and $(\xi_{1},\dots,\xi_{N})\in\mathbb{R}_{+}^{N}$ such that $\xi_{n}=Nw_{t}^{n}$ with $w_{t}^{n}$ being the softmax weight of the $n$-th particle, and we have $\sum_{n=1}^{N}\xi_{n}=N\in\mathbb{N}$.
+Output: $\left(Y_{\text{ssp}}^{1}(u),\dots,Y_{\text{ssp}}^{N}(u)\right)\in\mathbb{N}^{N}$ such that $\sum_{n=1}^{N}Y_{\text{ssp}}^{n}(u)=\sum_{n=1}^{N}\xi_{n}$.
+Initialization: $\left(Y_{\text{ssp}}^{1}(u),\dots,Y_{\text{ssp}}^{N}(u)\right)\leftarrow(\xi_{1},\dots,\xi_{N})$, $(n,m,k)\leftarrow(1,2,1)$
+Iterate the following steps until $\left(Y_{\text{ssp}}^{1}(u),\dots,Y_{\text{ssp}}^{N}(u)\right)\in\mathbb{N}^{N}$:
+(1) Let $\delta$ be the smallest number in $(0,1)$ such that at least one of $Y_{\text{ssp}}^{n}(u)+\delta$ or $Y_{\text{ssp}}^{m}(u)-\delta$ is an integer, and let $\epsilon$ be the smallest number in $(0,1)$ such that at least one of $Y_{\text{ssp}}^{n}(u)-\epsilon$ or $Y_{\text{ssp}}^{m}(u)+\epsilon$ is an integer.
+(2) If $u_{k}\leq\epsilon/(\delta+\epsilon)$ set $(Y_{\text{ssp}}^{n}(u),Y_{\text{ssp}}^{m}(u))\leftarrow(Y_{\text{ssp}}^{n}(u)+\delta,Y_{\text{ssp}}^{m}(u)-\delta)$; otherwise set $(Y_{\text{ssp}}^{n}(u),Y_{\text{ssp}}^{m}(u))\leftarrow(Y_{\text{ssp}}^{n}(u)-\epsilon,Y_{\text{ssp}}^{m}(u)+\epsilon)$.
+(3) Update $n$ and $m$ as follows:
+  1. 1.
+    If $\left(Y_{\text{ssp}}^{n}(u),Y_{\text{ssp}}^{m}(u)\right)\in\mathbb{N}^{2}$, $(n,m)\leftarrow(m+1,m+2)$;
+  2. 2.
+    If $Y_{\text{ssp}}^{n}(u)\in\mathbb{N}$ and $Y_{\text{ssp}}^{m}(u)\not\in\mathbb{N}$ set $(n,m)\leftarrow(m,m+1)$;
+  3. 3.
+    if $Y_{\text{ssp}}^{n}(u)\not\in\mathbb{N}$ and $Y_{\text{ssp}}^{m}(u)\in\mathbb{N}$ set $(n,m)\leftarrow(n,m+1)$.
+(4) $k\leftarrow k+1$
+There are more variance reduced sampling methods such as systematic resampling and stratified resampling, and we adopt the Srinivasan sampling process (SSP) resampling as our design choice.
+For other variance-reduced methods, the performance is similar to SSP resampling, all outperforming naive multinomial resampling.
+D.1.2 Overview of prior methods
+Here, we provide an overview of prior methods.
+SVDD (Li et al., 2024)
+In SVDD, the best sample is selected at each time step, from which $M$ children are generated.
+This approach can be viewed as a variant of BFS with $\tau=\infty$ and $M$ particles.
+Nodes are evaluated using the current score estimate $f({\bm{x}}_{0|t})$.
+TreeG (Guo et al., 2025)
+In TreeG, particles are ranked and the top $M$ are either selected directly or resampled based on their scores to obtain $M$ samples.
+From each selected particle, $K$ children are sampled, resulting in an effective tree width of $KM$.
+Particles are evaluated using their current score $f({\bm{x}}_{0|t})$.
+They also consider adding Gaussian noise to ${\bm{x}}_{0|t}$ to approximate the posterior distribution $p({\bm{x}}_{0}|{\bm{x}}_{t})$.
+However, the posterior distribution is intractable as it is not a Gaussian distribution, and exact Gaussian approximation requires computing the variance using the Jacobian of the score function (see Ye et al.
+(2024)).
+DAS (Kim et al., 2025)
+In DAS, the authors propose an exponentially increasing tempering schedule as the default, given by $\tau_{t}=(1+\gamma)^{T-t}-1$, and also introduce an adaptive tempering schedule.
+They adopt advanced SSP resampling instead of multinomial resampling, and evaluate particles based on the difference in rewards from the previous evaluation, similar to SMC methods.
+FK-steering (Singhal et al., 2025a)
+In FK, the authors propose several options for evaluating intermediate particles, including difference, max, and sum, with max adopted as the default.
+In the official implementation, multinomial resampling is used, which may lead to suboptimal performance.
 D.2 DFS-based search
 In this section, we provide the details and pseudo code for DFS in Alg.
-4.
+5.
 To better utilize previously explored sampling paths, we employ a buffer to store prior results.
-When no particles pass the threshold constraint, we retrieve the best sample from the buffer.
+When the budget is exhausted and no more backtracking is allowed, we retrieve the best sample from the buffer.
 Similar to BFS, controlling the set of evaluation steps allows a trade-off between efficiency and accuracy.
-Evaluating at earlier time steps introduces higher uncertainty but enables earlier resampling.
+Evaluating at earlier time steps introduces higher uncertainty but enables backtracking.
 Additionally, adjusting the backtracking depth $\Delta_{T}$ governs the search scope: a small $\Delta_{T}$ reduces computation and favors local search, while a larger $\Delta_{T}$ enables broader exploration at the cost of increased computation.
@@
 We set the recurrence depth to $T/2$ for image tasks and $T/4$ for PointMaze, corresponding to the denoised steps at which samples are first evaluated.
-The threshold schedule $\delta_{t}$ is also set to ‘increase’ as in Eq.13, enforcing tighter constraints for samples with lower noise.
-In our experiments, we observed that when backtracking to $t_{\text{next}}=T$—thus fully restarting—the nonzero terminal SNR $\alpha_{T}/\sigma_{T}$ in many diffusion schedules [40] can lead to cumulative errors with repeated backtracking.
+The threshold schedule $\delta_{t}$ is also set to ‘increase’ as in Eq.14 in our PointMaze experiments, enforcing tighter constraints for samples with lower noise.
+We use a simple constant $\delta_{t}=\delta$ schedule in our image experiments.
+In our experiments, we observed that when backtracking to $t_{\text{next}}=T$—thus fully restarting—the nonzero terminal SNR $\alpha_{T}/\sigma_{T}$ in many diffusion schedules (Lin et al., 2024) can lead to cumulative errors with repeated backtracking.
 Therefore, when backtracking to $t_{\text{next}}=T$, we initialize with fresh Gaussian noise.
@@
 Verifier $f$.
-  DFS input Budget for total number of backtracking $B=K$, backtracking depth $\Delta_{T}$ and threshold $\left\{\delta_{t}\right\}_{t=1}^{T}$. Set of evaluation time steps $\mathcal{S}$.
+DFS input: Budget for total number of backtracking $B=K$, backtracking depth $\Delta_{T}$ and threshold $\left\{\delta_{t}\right\}_{t=1}^{T}$.
 Init ${\bm{x}}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$, $t=T$. Init buffer with empty sets: $\texttt{buffer}(t)\leftarrow\left\{\right\}\,,t=1,2,\cdots,T$.
 while $t>0$ do
-     if $t\in\mathcal{S}$ then
-        Estimate the conditional mean and verifier score: ${\bm{x}}_{0|t}=\frac{{\bm{x}}_{t}-\sigma_{t}\epsilon_{\theta}({\bm{x}}_{t},t)}{\alpha_{t}}$, $f({\bm{x}}_{0|t})$.
+  Compute the conditional mean ${\bm{x}}_{0|t}=\frac{{\bm{x}}_{t}-\sigma_{t}\epsilon_{\theta}({\bm{x}}_{t},t)}{\alpha_{t}}$, and estimate the verifier score $f({\bm{x}}_{0|t})$.
+  Add the score-sample pair to the buffer: $\texttt{buffer}(t).\texttt{add}\left(f({\bm{x}}_{0|t}):{\bm{x}}_{t}\right)$
   if $f({\bm{x}}_{0|t})<\delta_{t}$ and budget $B>0$ then
@@
    Decrease the budget: $B\leftarrow B-1$
-           Add the score-value pair to the buffer: $\texttt{buffer}(t).\texttt{add}\left(f({\bm{x}}_{0|t}):{\bm{x}}_{t}\right)$
   else
-           if $f({\bm{x}}_{0|t})<\delta_{t}$ and budget $B=0$ then
-              Pop the best sample from buffer: ${\bm{x}}_{t}\leftarrow\texttt{buffer}(t).\texttt{max}$triangleright select the best sample from past explorations
+   if $B=0$ then
+    Pop the best sample from buffer: ${\bm{x}}_{t}\leftarrow\texttt{buffer}(t).\texttt{max}$   $\triangleright$ select the best sample from past explorations
    end if
-           Sample posterior: $t_{\text{next}}\leftarrow t-1$, ${\bm{x}}_{t_{\text{next}}}\sim\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$
-        end if
-     else
    Sample posterior: $t_{\text{next}}\leftarrow t-1$, ${\bm{x}}_{t_{\text{next}}}\sim\tilde{p}_{\theta}({\bm{x}}_{t-1}|{\bm{x}}_{t})$
@@
 In this section we provide the details of experimental setup and implementation for all our experiments. We run our experiments on clusters with Nvidia A100 GPUs, with over 1000 GPU hours used.
-E.1 Long Horizon Maze Planning
-Figure 7: Additional examples for PointMaze Ultra Fig.
-7(a) shows tha maze layout and a example trajectory rollout.
-Fig.
-7(b) shows the failure case of model generated plan.
-Fig.
-7(c) shows a successful generated plan after inference-time scaling.
+E.1 Ablation of BFS design space
+We directly adopt the official code base of FK-steering Singhal et al.
+(2025a) and use the samping methods provided in the code base of DAS (Kim et al., 2025).
+We use the ImageReward prompts as in (Singhal et al., 2025a) and report the average and standard deviation over 4 independent trials.
+For the temperature and resampling interval, we directly follow the implementation of FK-steering.
+For TreeG (Guo et al., 2025) we use a fixed branch out size of 2.
+The detailed hyper-parameters are below in Table.
+6:
+Table 6: BFS implementation parameters
+temperature 10
+resampling interval [20,40,80]
+sampling steps 100
+num seeds 4
+E.2 Text-to-Image Compositional Generation with DFS
+We use the SSD-1B model [^1] https://huggingface.co/segmind/SSD-1B which is distilled from SDXL, and we use the default sampling configuration with 50 steps of DDIM sampler.
+For DFS and BFS, we evaluate at time steps $\left\{25,35,45\right\}$ and set the backtrack depth $\Delta_{T}=25$.
+For BFS we additionally sweep the temperature in range $\left\{0.5,1,2,4,8\right\}$ and report the best performance.
+We provide example visualizations in Fig. 7.
+(a) w/o inference scaling
+(b) w/ inference scaling
+(c) Example from color dataset with prompt: a yellow book and a red vase
+(d) w/o inference scaling
+(e) w/ inference scaling
+(f) Example from texture dataset with prompt: a fabric hat and a glass mirror
+(g) w/o inference scaling
+(h) w/ inference scaling
+(i) Example from shape dataset with prompt: a round bagel and a square toaster
+Figure 7: Example of inference-time scaling with CompBench Above shows the example prompts from CompBench benchmark and images before and after inference scaling
+E.3 Long Horizon Maze Planning
 Maze environment
-For all our maze experiments we use the OGBench PointMaze [55].
-PointMaze Giant was directly adapted from their datasets, and we create PointMaze Ultra using the same protocol of Figure 5 in [49] [^1] https://github.com/mpetersen94/gcs/blob/main/models/maze.py, but with a smaller size of 20x20 cells.
-Dataset for Ultra is collected following the protocal in OGBench [55].
-A larger layout and example trajectory of PointMaze Ultra is shown in Fig.
-7.
-We evaluate the model on the default task 1 of OGBench [55], which is navigating from bottom left to top right.
+For all our maze experiments we use the OGBench PointMaze environment (Park et al., 2024).
+We created our maze layout using the same protocol of Figure 5 in Marcucci et al.
+(2023) [^2] https://github.com/mpetersen94/gcs/blob/main/models/maze.py, but with a smaller size of 20x20 cells.
+Dataset is collected following the protocal in OGBench (Park et al., 2024).
+We evaluate the model on the default task 1 of OGBench (Park et al., 2024), which is navigating from bottom left to top right.
 Empirically we discover that the diffusion model can perform well on short-horizon tasks without extra inference compute, but struggles heavily in the long horizon tasks.
 Model Training
-We train the model following diffuser [31], where we use a temporal U-Net to denoise the trajectory.
+We train the model following diffuser (Janner et al., 2022), where we use a temporal U-Net to denoise the trajectory
+$\bm{\tau}=\left[\begin{array}[]{cccc}\bm{s}_{1}&\bm{s}_{2}&\cdots&\bm{s}_{H}\\ \bm{a}_{1}&\bm{a}_{2}&\cdots&\bm{a}_{H}\end{array}\right]\,.$
 Since our objective start and goal is more distant than trajectories in dataset, we sample at longer horizons than training, which is enabled by the temporal U-Net architecture.
-Specifically, for PointMaze Giant, we train at horizon 400 and sample at 600; and for PointMaze Ultra, we train at horizon 2000 and sample at 2800.
-For the PointMaze Ultra model we extend the depth of the U-Net for 1 more layer, and keep all other parameters the same.
-We train the model for 1.2M steps using the same configuration as [31].
+We train the model for 1.2M steps using the same configuration as Janner et al.
+(2022).
+Model details are below in Table.
+7.
+Table 7: Hyper-parameters for maze model
+hidden dimension multipliers [1,4,8,16]
+training horizon 2000
+sampling horizon 2800
+training steps 1.2M
+batch size 64
+learning rate 1e-4
 Inference
@@
 For local search, we search the hyper-parameters $\bar{\rho}$ and $\bar{\mu}$ in Sec.
-C.5 with $\bar{\gamma}=0$, and we set the local search steps (recurrence steps) to $2$ for Giant and $6$ for Ultra.
-We set $N_{\text{iter}}=1$ as we observe they do not contribute to the performance gains, while the number of Langevin MCMC steps has the most impact on performance.
+C.5 with $\bar{\gamma}=0$.
 For global search with BFS we evaluate at steps $\left\{12,8,4\right\}$, and for DFS we evaluate at $\left\{12,11,\cdots,1\right\}$ with backtracking depth $\Delta_{T}=4$.
 We also observe that increasing backtracking depth to $12$ and evaluate at smaller time steps $\left\{4,3,2,1\right\}$ helps to scale up the performance with more compute.
-We use grid search to determine temperature and threshold, with a step size of $[1,5,10]$ for each magnitude.
-E.2 Offline RL
+The hyperparameter search results are below:
+N $\tau=0.2$ $\tau=0.005$ $\tau=0.1$
+2 $27.5\pm 4.3$ $32.5\pm 1.1$ $31.2\pm 4.2$
+4 $42.5\pm 5.2$ $48.1\pm 1.1$ $45.5\pm 2.3$
+8 $67.6\pm 1.1$ $71.2\pm 2.2$ $70.1\pm 1.1$
+Table 8: Hyperparameter search for temperature $\tau$ in PointMaze BFS
+Figure 8: Hyperparameter search for threshold $\delta$ in PointMaze DFS
+E.4 Offline RL
+Background
+Diffusion policy (Chi et al., 2023) is widely used for action generation in robot foundation models (Black et al., 2024; Liu et al., 2024).
+At inference time, policies can be guided by human trajectory constraints (Wang et al., 2022) or LLM-based value functions (Nakamoto et al., 2024).
+Exact sampling requires training a noise-dependent energy function (Lu et al., 2023), but this can degrade pretrained knowledge and demands additional data—often impractical in data-scarce robotic settings.
+In contrast, inference-scaling provides a more flexible approach, allowing seamless composition of pretrained diffusion policies with Q-functions without retraining.
 Setup
-We follow the setup in [44], and we directly use their pretrained diffusion model and Q-function, omitting the time-dependent energy function.
+We follow the setup in Lu et al.
+(2023), and we directly use their pretrained diffusion model and Q-function, omitting the time-dependent energy function.
 The diffusion model was trained to generate action $\bm{a}$ given state $\bm{s}$, and we sample with 15 steps of DDIM.
@@
 For strength parameters $\bar{\rho}$ and $\bar{\mu}$, we first search for the right magnitude.
-Then, we also follow [44] and search with step size [1,2,3,5,8,10] within the magnitude.
-Same as [44], we use 5 different seeds with 10 samples per seed for each task.
+Then, we also follow Lu et al.
+(2023) and search with step size [1,2,3,5,8,10] within the magnitude.
+Same as Lu et al.
+(2023), we use 5 different seeds with 10 samples per seed for each task.
 To avoid over fitting, we use different seeds for parameter search and evaluation.
@@
 Since the number of particles are small, we do not carry out BFS or DFS methods and simply use Best-of-N.
-We point out that the number of particles we use are much smaller than the 50 particles in [80] and the 32 particles in [8], highlighting the effectiveness of local search.
+We point out that the number of particles we use are much smaller than the 50 particles in Wang et al.
+(2022) and the 32 particles in Chen et al.
+(2023), highlighting the effectiveness of local search.
 Baseline
-We compare our method to a variety of baselines, including traditional state-of-the-art methods like CQL [35], BCQ [20] and IQL [34], diffusion-based policies such as diffuser [31], decision-diffuser (DD) [1], Diffusion-QL (D-QL) [80], SfBS [8] and QGPO [44].
-We directly take the numbers from [44].
-Among the baseline diffusion-based methods, both Diffuser [31] and QGPO [44] requires training a noise-dependent guidance function, and D-QL [80] requires updating the diffusion model during training using the Q-function iteratively, which needs to back-propagate through the diffusion sampling chain, introducing high computation and memory overheads.
-DD [1] uses classifier-free guidance [27] to generate high-return trajectories, which requires training a return-conditional model on labeled datasets, which can be expensive to obtain in robotics where only demonstration data is available [43].
+We compare our method to a variety of baselines, including traditional state-of-the-art methods IQL (Kostrikov et al., 2021) and diffusion-based policies such as diffuser (Janner et al., 2022), decision-diffuser (DD) (Ajay et al., 2022), Diffusion-QL (D-QL) (Wang et al., 2022), SfBC (Chen et al., 2023) and QGPO (Lu et al., 2023).
+We directly take the numbers from Lu et al.
+(2023).
+Among the baseline diffusion-based methods, both Diffuser (Janner et al., 2022) and QGPO (Lu et al., 2023) requires training a noise-dependent guidance function, and D-QL (Wang et al., 2022) requires updating the diffusion model during training using the Q-function iteratively, which needs to back-propagate through the diffusion sampling chain, introducing high computation and memory overheads.
+DD (Ajay et al., 2022) uses classifier-free guidance (Ho & Salimans, 2022) to generate high-return trajectories that requires training a return-conditional model on labeled datasets, which can be expensive to obtain in robotics where only demonstration data is available (Liu et al., 2024).
+For our reproduced baselines, TFG (Ye et al., 2024) is allowed up to 8 recurrence steps and DAS (Kim et al., 2025) up to 16 particles, resulting in a hyperparameter space and computational cost approximately twice that of our method.
+We sweep across all configurations for the baseline methods and report the best performance.
+For fair comparison we evaluate our method on different seeds used for hyperparameter search, with the results shown in Table.
+9.
 Dataset Environment particles $N_{\text{recur}}$ $N_{\text{iter}}$ $\bar{\rho}$ $\bar{\mu}$ Eval set Search set
@@
 Table 9: Hyper-parameters on D4RL locomotion tasks with test-time scaling. We report the performance on hyper-parameter search dataset and the evaluation dataset, highlighting the best number.
-E.3 Image Generation
-E.3.1 Text-to-Image Compositional Generation
+E.4.1 Offline Policy Distillation
 Setup
-We use the SSD-1B model [^2] https://huggingface.co/segmind/SSD-1B which is distilled from SDXL, and we use the default sampling configuration with 50 steps of DDIM sampler.
-We use the oracle verifier of each task following CompBench [29].
-Global search with DFS
-For DFS we evaluate at time steps $\left\{25,35,45\right\}$ and set the backtrack depth $\Delta_{T}=25$.
-We set the threshold for our experiments to $0.5$, $0.7$ and $0.9$, with the ‘constant’ schedule in Eq.
-14.
-We typically use a higher threshold for high computation budgets, representing using more compute for higher quality standard.
-We provide example visualizations in Fig. 8.
-(a) 1 Particle
-(b) 2 Particles
-(c) 4 Particles
-(d) 8 Particles
-Figure 8: Example of inference-time scaling with Numeracy dataset Above shows the results of Best-of-N with different particles, with prompt: “four dogs played with two toys”
-E.3.2 ImageNet Conditional Generation
-Setup
-We use the 256x256 DDPM ImageNet model from [12], and for local search we use the 224px ViT classifier [^3] https://huggingface.co/google/vit-base-patch16-224 and for global search we use the 384px version [^4] https://huggingface.co/google/vit-base-patch16-384 when using double-verifier.
-We evaluate on class label 222 and 444, with 256 samples per label.
-We sample with 100 DDPM steps.
-When evaluating the class accuracy we use the DeiT model [^5] https://huggingface.co/facebook/deit-small-patch16-224, and we evaluate the FID againt the images of the corresponding class in ImageNet.
-Global search with BFS
-Due to computation constraints we set the local search steps to $1$ in all experiments.
-For BFS we set the evaluation steps as $\left\{75,50,25\right\}$, and use $\bar{\tau}=1.0$ for label 222 and $\bar{\tau}=0.5$ for label 444 with the ‘increase’ schedule.
-We provide example visualizations in Fig. 9.
-(a) 2 Particles
-(b) 4 Particles
-(c) 8 Particles
-(d) 12 Particles
-Figure 9: Visualizations of ImageNet Conditional Generation Above shows the uncurated samples for class label 222 with BFS with different number of particles
+We adopt the Medium dataset from D4RL (Fu et al., 2020) and the corresponding pretrained models from Lu et al.
+(2023) as in the previous section.
+For the dataset, we replace the action of each state with the sample generated with test-time search (TTS), using the hyper-parameters in Table.
+9.
+We then finetune the models on the modified dataset with early-stopping.
+For evaluation of finetuned models, we use 5 different seeds different from the seed used during training, with 10 samples per seed.
+During evaluation, we sample from the finetuned diffusion model directly, without the presence of any verifier.
+Detailed training hyper-parameters are below in Table.
+10.
+epochs 5
+learning rate 1e-4
+batch size 16384
+eval every epoch 1
+sampling steps 15
+Table 10: Hyper-parameters for policy distillation
+For the Baseline DPPO (Ren et al., 2024), we directly adopt the numbers reported in their paper, which uses the same Medium dataset and models as ours.
+The training curves for policy finetuning are shown in Fig.
+9.
+In both Hopper and Walker2d environments, performance converges rapidly, within a single epoch over the dataset.
+Notably, our method is significantly faster than online finetuning approaches, as it does not require online data collection.
+Training one epoch on the full Medium dataset takes only about 4 minutes on a single Nvidia RTX 4090 GPU, including the time for data generation via inference-scaling sampling, whereas DPPO requires several hours to reach convergence.
+Figure 9: Training curves for policy distillation
+E.5 Mitigating reward hacking with double verifier
+Experimental setup
+We evaluate the proposed double-verifier on the challenging conditional ImageNet generation task, generating target-class samples from an unconditional model guided by a pretrained classifier.
+Specifically, we use two independent classifiers as verifiers [^3] https://huggingface.co/google/vit-base-patch16-224444https://huggingface.co/google/vit-base-patch16-384 for global and local search.
+We report the Fréchet Inception Distance (FID) computed on 256 generated samples against the corresponding ImageNet class, and measure class accuracy using a separate classifier [^5] https://huggingface.co/facebook/deit-small-patch16-224.
+Since we only apply the global verifier sparsely, double-verfier introduces negligible computational costs.
+For local search we search the hyper-parameters with step size [1,2,4,8,10], and for global search with BFS, we search the temperature in [1.0,2.0,4.0].
+For MSP score, we adopt a different imagenet classifier  [^6] https://huggingface.co/facebook/deit-base-patch16-224.
+For the MSP score (Hendrycks & Gimpel, 2016), it is defined as:
+$\text{MSP}({\bm{x}})=\max_{c}p(c|{\bm{x}})$
+Thus, higher MSP score indicates higher confidence that the image belongs to one of the classes from the dataset, which is less OOD.
+Visualizations
+Here we provide some visualizations of adversarial reward hacking and the effectiveness of our double-verifier approach.
+As shown in Fig.
+10, we can see the adversarial patches that exploits the weakness of a single verifier, making the verifier to predict the images as belonging to the wrong class.
+(a) Adversarial example with 0.96 class validity predicted by the local search verifier
+(b) Adversarial example with 0.95 class validity predicted by the local search verifier
+(c) Adversarial example with 0.98 class validity predicted by the local search verifier
+Figure 10: Examples of adversarial samples generated by gradient over-optimization Here we provide the adversarial samples generated by gradient guidance on class 222 (kuvasz).
+Although the samples have high probability of belonging to the target class as predicted by the verifier, they are caused by adversarial reward hacking.